### v1

# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

Abstract
Bridging the microscopic/individual and the macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control tasks. Here, we propose a combined manifold and machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations. The proposed framework builds upon our previous works on the development of next-generation Equation-free algorithms for dealing with the “curse of dimensionality” when learning surrogate models for high-dimensional and multiscale systems. The proposed approach is a four-stage one, explicitly conserving the mass of the reconstructed dynamics in the high-dimensional space. In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation (KDE). In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distribution. The third step involves learning reduced-order surrogate models (ROMs) in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs). Finally, we reconstruct the crowd dynamics in the high-dimensional space in terms of macroscopic density profiles. We demonstrate that the POD reconstruction of the density distribution via SVD conserves the mass. With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation (PDE) for the evolution of the density distribution. For our illustrations, we use the Social Force Model (SFM) to generate data in a corridor with an obstacle, imposing periodic boundary conditions. The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations. Notably, linear MVAR models surpass their nonlinear counterparts, such as LSTMs, in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
Keywords Crowd Dynamics, Multiscale modelling, Micro to Macro, Machine Learning, Numerical Analysis, Conservation laws, Reduced Order Models, Manifold Learning
1 Introduction
An important open challenge in the modelling of human crowd dynamics lies in bridging across the different modelling scales. Typically, models are formulated at three distinct levels: the microscopic (individual), mesoscopic (kinetic), and macroscopic (hydrodynamic) scales. Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [1] and its variants (see e.g. [2, 3]), cellular automata [4, 5, 6], etc. The mesoscopic scale serves as an intermediary between the microscopic and macroscopic approaches [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. At the macroscopic scale, modelling describes the emergent/collective behavior of crowds. For instance, in the seminal work of Hughes [17], the author models crowd dynamics by coupling a conservation law with an eikonal equation to represent how individuals choose paths based on a potential field. Moreover, in [18] the authors proposed nonlocal conservation laws for traffic and crowd dynamics, where the flux depends not only on local density but also on averaged information from the surrounding region. In [11] a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior. More recently, in [16] are highlighted new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
A major objective, when high-fidelity microsocpic/individual/agent-based simulators and/or detailed experimental data are available, is to derive mesoscopic or macroscopic models that can be then used for rigorous numerical analysis, optimization and control of the emergent/collective dynamics, for instance, in the optimal design of spaces, the reduction of evacuation times, and the management of crowd flows [11, 19, 20]. In this context, kinetic models serve as an intermediate bridge. They are typically derived analytically using statistical mechanics techniques, starting from microscopic interaction rules. These kinetic formulations, in turn, facilitate the derivation of macroscopic (continuum) models through asymptotic methods [10, 16].
The transition from high-dimensional agent-based representations to macroscopic or hydrodynamic descriptions is commonly achieved through closure techniques. These methods establish relationships between higher-order moments of the microscopic distributions and a limited, smaller set of lower-order moments that capture the observable emergent behavior [21, 22, 23]. However, such closures inherently rely on assumptions that introduce biases both for modelling and analyses tasks. Typically, these assumptions include infinitely large populations, homogeneous agents, and uniform and/or local interaction networks, but also unknown “behavioural” characteristics, which can significantly impact the accuracy and validity of collective-level analyses.
Recently, data-driven computational approaches based on physics-free machine learning (ML) have been employed for the modelling, numerical analysis, and control of pedestrian, crowd, and traffic dynamics. In this context, a wide range of ML architectures has been explored, including deep neural networks (DNNs) [24], Long Short-Term Memory (LSTM) networks [25, 26] for learning local movement decisions, Generative Adversarial Networks (GANs) [27, 28], Convolutional Neural Networks (CNNs) [29, 30, 31], and reinforcement learning methods [32, 33]. For a comprehensive review of these and related ML approaches, see [34]. These approaches have proven effective in forecasting individual and group trajectories within the framework of time-series modelling and forecasting, without relying on physics-based formulations. However, at the same time, such approaches lack interpretability and, importantly, do not incorporate information of the global “spatio-temporal” emergent dynamics, such as density fields, that are essential in modelling crowd dynamics.
An alternative data-driven paradigm that bypasses the need to explicitly learn surrogate models is the Equation-free multiscale framework [35]. By appropriately coupling on-demand microscopic simulations with coarse-graining (and fine-graining), this approach builds a coarse time-stepper for the evolution at the macroscopic scale, thus bypassing the need for analytically derived continuum equations [36, 37, 38]. While Equation-free framework facilitates powerful multiscale numerical analysis and control, it lacks an explicit, global macroscopic equation for the crowd/collective dynamics–that can be approximated, such as a mass-conserving PDE–limits interpretability, physical insight, and generalization for long-term simulations.
To address the aforementioned challenges, a few recent studies have utilized physics-informed neural networks (PINNs) [39] including physics-informed neural operators (NOs) [40, 41] for solving the inverse problem, i.e. that of constructing from data, hybrid models in the form of “grey-box” PDE models [42, 43, 21, 23, 41]. Such an approach, combines ML with underlying knowledge of physical principles, such as conservation laws [44] and closures [45], to create hybrid models that can both predict the spatio-temporal dynamics and ensure physical consistency, even under data scarcity. In the context of crowd dynamics, PINNs have been employed to construct both gray-box hydrodynamic PDEs [46] as well as for discovering the physics/rules governing interactions and motion at the microscopic level [47]. However, while PINNs provide an effective way to incorporate physical constraints and knowledge into machine learning models, they can also introduce biases due to their reliance on predefined forms of PDEs. Moreover, a key physical property in crowd modeling—mass conservation—is typically not enforced explicitly, but rather imposed as a soft constraint through the loss function [48, 44, 49]. This indirect enforcement can limit the generalizability and physical reliability of PINN-based models, particularly in heterogeneous or unforeseen scenarios where the true dynamics deviate from the assumed PDE forms.
For learning classes of PDEs that incorporate conservation laws—such as those found in crowd dynamics—the use of black-box deep neural networks (DNNs) and black-box neural operators (NOs) [50, 51, 49, 52, 53, 23] is inherently problematic. These models often lack physical consistency, particularly in preserving conservation laws, which are usually enforced only as soft constraints in the loss function. In addition, their training suffers from low approximation accuracy due to the curse of dimensionality, potentially leading to unphysical predictions and numerical instabilities during long-term simulations. Thus, to deal with that problem, RiemannONets were recently presented in [54], which are NOs based on DeepONets to solve Riemann problems encountered in compressible flows, thus enforcing the preservation of conservation laws. Towards solving the inverse problem, we have recently proposed GoRINNs [55]: Godunov-Riemann informed neural networks for learning hyperbolic PDEs that respect the conservation laws explicitly.
It is important to note, however, that despite the long-standing tradition of modeling crowd dynamics through (hydrodynamic) PDEs, such formulations inherently rely on an “infinite-size” assumption. In realistic settings involving crowds of moderate or finite size, this assumption often breaks down. Finite-size effects—such as spatial heterogeneity, individual interactions, and boundary influences—can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior.
Here, inspired by the Equation-free multiscale framework [35], and building on our previous recent works on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [56, 36, 57, 23, 58, 41], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics. The proposed framework bridges the microscopic and macroscopic modelling scales, while explicitly preserving the underlying conservation law of mass in the ambient space. The methodology is deployed in four main stages. First, we map discrete microscopic distributions to continuous macroscopic representations (i.e., density fields). Second, we project the macroscopic fields into a low-dimensional latent space using Proper Orthogonal Decomposition (POD). Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM networks. These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’ theorem [59] for reconstructing the phase-space structure (here in the latent space). The concept of the approach (i.e. that using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [60, 61] (see also results and discussion in [61, 56, 62, 63, 57]). In crowd dynamics, behavioural characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Taken’s theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics. Finally, we lift the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis. Crucially, this configuration ensures explicit mass conservation—a key physical constraint for any effective continuum description. For our numerical experiments, we employ the Social Force Model (SFM) [64] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions. The performance and generalization ability of the proposed framework are assessed on unseen initial conditions.
2 Methodology
2.1 Problem Statement
In this framework, we consider the problem of bridging the scales in crowd dynamics with particular focus on deriving ROMs from detailed, high-dimensional, agent-based simulations of pedestrian flows. This derivation involves reconstructing on demand the spatio-temporal dynamics under the strict constraint of mass conservation.
We utilize a detailed simulator that models the dynamics of a human crowd, consisting of $N$ pedestrians moving within a domain $\Omega\subset\mathbb{R}^{2}$. The overall state of the pedestrian $i\in\{1,\ldots,N\}$ is characterized by its position $\mathbf{x}_{i}=\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ and velocity $\mathbf{v}_{i}=\mathbf{v}_{i}(t)=(v_{ix}(t),v_{iy}(t))$, and also a set of “behavioral” variables $\bm{u}_{i}=\bm{u}_{i}(t)\in\mathbb{R}^{b}$, which may include phenomena such as decision-making strategies, herding behavior, but also panic and irrationality [65, 66, 67, 68, 69], where $b\in\mathbb{R}$ is the number of behavioral traits modeled for each pedestrian. Additionally, we introduce an auxiliary variable $\bm{z}_{i}\in\mathbb{R}^{b}$ to denote the rate of change of the behavioral variables $\bm{u}_{i}$. Then, the evolution of the position, velocity and behavior of each pedestrian is modelled by an individualistic/agent-based model, inspired by pseudo-Newtonian/molecular/Brownian dynamics formulations [70, 16], expressed in the general form:
$\frac{d\bm{u}_{i}}{dt}=\bm{z}_{i},\quad\frac{d\bm{z}_{i}}{dt}=\sum_{j\in\Omega_{i}}\,\bm{\psi}_{i}(\mathbf{x}_{i},\mathbf{v}_{i},\bm{u}_{i},\mathbf{x}_{j},\mathbf{v}_{j},\bm{u}_{j}),$ (1)
$\frac{d\mathbf{x}_{i}}{dt}=\mathbf{v}_{i},\quad\frac{d\mathbf{v}_{i}}{dt}=\sum_{j\in\Omega_{i}}\boldsymbol{\phi}_{i}(\mathbf{x}_{i},\mathbf{v}_{i},\bm{u}_{i},\mathbf{x}_{j},\mathbf{v}_{j},\bm{u}_{j}),$ (2)
where $j$ refers to all pedestrians within the interaction domain of the $i$-th pedestrian $\Omega_{i}$, and $\bm{\psi}_{i}(\cdot),\boldsymbol{\phi}_{i}(\cdot)$ denote pseudo-acceleration terms incorporating interaction rules. The majority of models in this category are based on the celebrated social force model (SFM) for pedestrian dynamics [1, 16], which was proposed in the mid-1990s.
In contrast, at the macroscopic scale, crowd dynamics are typically described using phenomenological continuum models, most often formulated as partial differential equations (PDEs). These models capture the collective behavior of large-scale pedestrian flows and are grounded in the theories of hydrodynamics and continuum mechanics. Well‑known examples include the so‑called Hughes model [17] and multi‑population nonlocal models [18]. However, such modelling approaches rely on several simplifying assumptions in order to derive algebraic closure relations that link individual-level motion with density‑scale dynamics. These assumptions often include, for example, that pedestrian speed depends solely on the local total density, that individuals aim to minimize travel time while avoiding congestion, or that their motion is guided by a potential field directing them toward their destination.
To address the limitations of phenomenological continuum models, several data-driven computational methods have been introduced to construct macroscopic hybrid models in the form of “black-box” or “grey-box” PDE models [50, 51, 49, 52, 53, 23, 42, 43, 21]. These methods are predominantly based on DNNs/NOs, thus suffering from the “curse of dimensionality” that is inherent in their training. To circumvent this challenge, learning Reduced Order Models (ROMs) in well-constructed latent spaces is crucial, since it significantly lowers computational cost while capturing the essential dynamics. These ROMs must also inherently preserve fundamental principles like mass conservation, ensuring physical accuracy in the reconstructed space. Towards this goal, we aim at bridging high-fidelity microscopic/agent-based simulations and/or experimental/real-world data with macroscopic scale modelling, using ROMs that preserve mass conservation, within the framework of learning discrete operators/coarse-timesteppers [35].
The proposed methodology, presented in detail in the following section, is based on our previous efforts to develop what we call next generation Equation-free framework [56, 36, 57, 23, 58]. This framework deals with the “curse of dimensionality” by learning surrogate ML models in appropriate latent spaces emerging from the complex spatio-temporal dynamics of high-dimensional systems. It originates from the celebrated Equation-free approach [35], which enables high-dimensional simulations to perform system-level numerical analysis tasks on an appropriate latent space. This is achieved via the construction of a lifting operator $\mathcal{L}$ and a restriction operator $\mathcal{R}$ bridging the gap between states in a high-dimensional space, say $\mathbf{u}(t)\in\mathbb{R}^{M}$, and states in a latent space, say $\mathbf{y}_{d}(t)\in\mathbb{R}^{d}$, with $d\ll M$, as:
$\mathbf{u}(t)=\mathcal{L}(\mathbf{y}_{d}(t))\quad\text{and}\quad\mathbf{y}_{d}(t)=\mathcal{R}(\mathbf{u}(t)).$ (3)
These operators are used to construct a coarse-timestepper, the discrete evolution operator at the latent space, $F_{\delta t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined as:
$\mathbf{y}_{d}(t+\delta t)=\mathcal{R}\Bigl{(}{\Phi_{\Delta t}}\big{(}\mathcal{L}(\mathbf{u}(t))\big{)}\Bigr{)}:=F_{\delta t}\big{(}\mathbf{y}_{d}(t)\big{)},$ (4)
where $\Phi_{\Delta t}:\mathbb{R}^{M}\to\mathbb{R}^{M}$ is the evolution operator at the high-dimensional space. In this work, we learn the discrete evolution operator $F_{\delta t}$ using autoregressive ROMs, thus effectively learning $\Phi_{\Delta t}$ via the restriction and lifting operators, as:
$\mathbf{u}(t+\Delta t)=\mathcal{L}\Bigl{(}F_{\delta t}\big{(}\mathcal{R}(\mathbf{u}(t)),\ldots,\mathcal{R}(\mathbf{u}(t-w\delta t))\big{)}\Bigl{)},$ (5)
where the learned ROMs account for $w$ delayed embeddings $\mathcal{R}(\mathbf{u}(t-j\delta t))$ for $j=1,\ldots,w$.
A key distinction between our proposed next generation EF and the traditional EF framework is that, in the latter, one constructs local (numerical) maps and computes the quantities needed for bifurcation analysis on demand via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured usually by a few moment distributions). In contrast, our proposed method learns a “global dynamical model“ in Eq. (5) via machine learning from long-term, high-fidelity simulations spanning the entire phase space, in low-dimensional latent spaces. This global dynamical-systems-based perspective enables the use of both linear (such as POD) and nonlinear manifold-learning techniques (such as Diffusion maps [71, 72]), and techniques for the solution of the pre-image problem, such as geometric harmonics [73, 74] or autoencoders [75, 76], to construct maps between the high-dimensional and the latent space. The methodology is described analytically in the following section.
The key advantage of our next generation EF framework is its ability to circumvent the need for explicit derivation or closure assumptions in PDEs—a process that inherently introduces biases in modelling crowd density dynamic. Instead, it directly learns the “actual” solution operator in Eq. (5)–-the map that advances observed macroscopic distributions (here the density) forward in time. This learned operator, embedded within computationally efficient linear or nonlinear autoregressive ROMs (MVAR or LSTM, respectively), effectively encodes the complex, often analytically intractable, dynamics of the crowd. By focusing on learning how densities evolve (the solution operator) rather than learning governing equations in the form of PDEs, the method circumvents the significant challenges of the “curse of dimensionality” associated with DNNs/NOs training of such PDEs. The resulting ROMs thus act as practical, data-driven surrogates that implicitly contain the action of the unknown PDE’s solution operator, enabling accurate prediction of future crowd states solely from observed density snapshots. Importantly, we demonstrate that for the proposed configuration, the POD method (used for identifying the latent space) conserves explicitly the mass of the reconstructed density field resulting from the ROMs. An explicit mass conservation of the continuum representations is an essential physical constraint for any credible surrogate model of crowd dynamics.
2.2 The proposed ML framework
The proposed methodology consists of four main steps. First, we map the discrete distributions of pedestrian positions into continuous density fields defined on a spatial grid using Kernel Density Estimation (KDE) [77]. Second, we construct a restriction operator to project the resulting density fields into an appropriate latent space. For this purpose, we employ the POD [78] method on normalized density profiles across the computational domain, ensuring that mass conservation is preserved in the reconstructed fields. Third, within this latent space, we train surrogate autoregressive ROMs using MVAR and LSTM time-series forecasting models to capture the complex dynamics present in the sequential data. Finally, in the fourth step, we lift the latent ROM dynamics back to the macroscopic density ambient space by solving the preimage problem. In the case of POD, this lifting step is computationally inexpensive, as it reduces to a simple linear projection onto the ambient space using the stored orthogonal basis. A schematic overview of the methodology is shown in Fig. 1.
Figure 1: Schematic of the proposed methodology. First, microscopic distributions of pedestrian positions are mapped to macroscopic density fields (Step 1) and embedded into a latent space using POD (Step 2). Second, surrogate reduced-order models (MVAR and LSTM) are trained in this latent space to capture and forecast the complex dynamics (Step 3). Finally, the learned dynamics are lifted back to the high-dimensional density fields via linear projection with the constructed POD basis (Step 4).
2.2.1 From discrete pedestrian positions to continuous density fields
The initial step in our framework involves constructing a continuous density field from the discrete positions of pedestrians. Given microscopic observations, obtained by agent-based simulators in the form of Eq. (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves. Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator in Appendix A. We next employ Multivariate Kernel Density Estimation (KDE) [79] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
$\rho(\mathbf{x},t)=\frac{1}{N}\sum_{i=1}^{N}K_{h}\left(\mathbf{x}-\mathbf{x}_{i}(t)\right),$ (6)
where $K_{h}$ is a smooth non-negative kernel function satisfying $\int_{\Omega}K(\mathbf{x})\,d\mathbf{x}=1$. Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(c-d)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq. (6) density fields are computed as $\rho(\mathbf{x},t)=[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $\rho((x_{i},y_{j}),t)$ corresponds to the density computed at the grid point $(x_{i},y_{j})$, for $i=1,\dots,n_{x}$, $j=1,\dots,n_{y}$. Then, the total mass in the domain is given by $S(t)=\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y$. While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct the density fields in Eq. (6). Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid-points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
2.2.2 The Restriction and Lifting Operators
Having obtained the macroscopic density fields, we now proceed to the second and fourth key elements of the approach, which, in analogy to the Equation-free framework, constitute the construction of an “embedding/restriction” and “lifting” operators to and from an appropriate low-dimensional latent space.
The restriction operator, say, $\mathcal{R}$, [35], is a map from density distributions $\rho(\mathbf{x},t)$ to a low-dimensional latent space coordinates, say $Y_{d}(t)$, i.e.,:
$\mathcal{R}:\rho(\mathbf{x},t)\in\mathbb{R}^{n_{c}}\mapsto Y_{d}(t)\in\mathbb{R}^{d},\quad d\ll n_{c},$ (7)
where $n_{c}=n_{x}\times n_{y}$. To construct the restriction operator, we apply Proper Orthogonal Decomposition (POD) [78] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq. (6), for different initial conditions; see Appendix B for details. Denoting the densities by $\rho^{(c)}(\mathbf{x},t_{0}+k\delta t)\equiv\rho^{(c)}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$, where $\delta t$ is a microscopic inner simulation step/observation sampling time of an experiment/real data, $t_{0}$ is the initial time, and the superscript $(c)$ denotes the case $c=1,\ldots,C$ of the different initial conditions considered, the training dataset generated by microscopic runs is represented by the matrix $X\in\mathbb{R}^{n_{c}\times(C\cdot K)}$:
$X=\Bigl{\{}\rho^{(c)}(\mathbf{x},t_{k})\,:\,k=0,\ldots,K,\,c=1,\ldots,C\Bigr{\}}.$ (8)
Compactly, we represent the above data set as:
$X=[x_{1},x_{2},\dots,x_{n_{t}}]\,\in\mathbb{R}^{n_{c}\times n_{t}},$ (9)
with columns, say, $x_{k}\in\mathbb{R}^{n_{c}}$ being the normalized density field along cases and timesteps $k=1,\ldots,n_{t}=C\cdot K$.
To complete the restriction operator at the second step, we employ POD on $X$, implemented via Singular Value Decomposition (SVD) on the centered matrix:
$\bar{X}=XH=U\Sigma V^{T},\quad H=I_{n_{t}}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top},$ (10)
where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular vectors, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim. unity matrix and $\mathbf{1}_{n_{t}}$ is the $n_{t}$-dim. unit column vector. The latent space is spanned by the first $d$ left-singular vectors forming $U_{d}$. Thus, projecting $X$ on the latent space we get, the low-dimensional embedding:
$Y_{d}=U_{d}^{T}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
For the lifting operator that maps $Y_{d}$ to the space of density profiles, we demonstrate that the POD reconstruction (aka the POD lifting operator) preserves the original density. This aligns with the general characteristic of POD in maintaining symmetries [80]. Here, in particular, we demonstrate the following proposition.
Proposition 1.
Let us assume a matrix $X\in\mathbb{R}^{n_{c}\times n_{t}}$, with columns being normalized density fields. If the density is preserved along timesteps, i.e., if
$\begin{bmatrix}\sum_{i=1}^{n_{c}}x_{i,1}&\sum_{i=1}^{n_{c}}x_{i,2}&\dots&\sum_{i=1}^{n_{c}}x_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}},$ (12)
then the reconstructed density field computed by the POD as:
$\tilde{X}=U_{d}U_{d}^{\top}\bar{X}+X(I_{n_{t}}-H),\,\,$ (13)
where $U_{d}=[u_{1},u_{2},\dots u_{d}]\in\mathbb{R}^{n_{c}\times d}$ is the orthonormal basis formed by the first $d$ left-singular vectors computed by the SVD, is also preserved, i.e.,:
$\begin{bmatrix}\sum_{i=1}^{n_{c}}\tilde{x}_{i,1}&\sum_{i=1}^{n_{c}}\tilde{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\tilde{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (14)
Proof.
As by hypothesis, the sum of each column of $X$ is equal to one (that is $\mathbf{1}_{n_{c}}^{\top}X=\mathbf{1}_{n_{t}}^{\top}$), we have that:
$\mathbf{1}_{n_{c}}^{\top}\bar{X}=\mathbf{1}_{n_{c}}^{\top}XH=\mathbf{1}_{n_{t}}^{\top}H=\mathbf{1}_{n_{t}}^{\top}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}^{\top}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top}=\mathbf{1}_{n_{t}}^{\top}-\mathbf{1}_{n_{t}}^{\top}=\boldsymbol{0}^{\top}_{n_{t}}.$ (15)
Therefore, the sum of each column of the centered matrix $\bar{X}$ is equal to zero.
This implies that $\mathbf{1}_{n_{c}}$ is orthogonal to the column space of $\bar{X}$, and therefore is orthogonal to the left singular vectors of the SVD decomposition of $\bar{X}$, which provide an orthogonal base for the column space, i.e.,:
$\mathbf{1}_{n_{c}}^{\top}U_{d}=\boldsymbol{0}^{\top}_{d},$ (16)
where $d=1,2,\dots r$, with $r=rank(X)$. Then, the multiplication of the reconstructed matrix in Eq. (13) by $\mathbf{1}_{n_{c}}^{\top}$ implies:
$\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\mathbf{1}_{n_{c}}^{\top}U_{d}U_{d}^{\top}\bar{X}+\mathbf{1}_{n_{c}}^{\top}X(I_{n_{t}}-H)=\mathbf{0}_{n_{t}}^{\top}+\frac{1}{n_{t}}\mathbf{1}_{n_{t}}^{\top}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top}=\mathbf{1}_{n_{t}}^{\top},$ (17)
thus getting:
$\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\begin{bmatrix}\sum_{i=1}^{n_{c}}\hat{x}_{i,1}&\sum_{i=1}^{n_{c}}\hat{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\hat{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (18)
Hence, the POD reconstruction preserves the original density. ∎
In summary, given a density distribution $\rho(\mathbf{x},t)$ obtained via Eq. (6), we obtain the normalized density along the $n_{c}$ grid points of $\Omega$ as:
$x(t)=\dfrac{1}{\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}}$ (19)
and define the restriction operator $\mathcal{R}$ by the projection on the POD basis of Eq. (11) as:
$\boldsymbol{y}(t)=\mathcal{R}(x(t))=U_{d}^{\top}\left(x(t)-\frac{1}{n_{t}}X\mathbf{1}_{n_{t}}\right)\in\mathbb{R}^{d},$ (20)
where $X$ is the dataset matrix in Eq. (8) and $U_{d}$ is the matrix containing the first $d$ left-singular vectors. On the other hand, the lifting operator $\mathcal{L}$ that maps latent coordinates, say $\boldsymbol{y}(t)$, back to the density distributions is defined via the POD reconstruction as:
$x(t)=\mathcal{L}(\boldsymbol{y}(t))=U_{d}\boldsymbol{y}(t)+\frac{1}{n_{t}}X\mathbf{1}_{n_{t}}\in\mathbb{R}^{n_{c}}.$ (21)
The above definitions of the restriction and lifting operators guarantee that the reconstructed, by the lifting operator, normalized density profiles preserve the total density within the domain $\Omega$, as per Proposition 1. We highlight here that while the operators are constructed based on the dataset $X$, their employment applies to any seen or unseen observation $x(t)\in\mathbb{R}^{n_{c}}$ obtained from Eq. (19) and $\boldsymbol{y}(t)\in\mathbb{R}^{d}$.
2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
With the restriction operator constructed, which maps density fields to latent coordinates, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data. We note that while the restriction and lifting operators $\mathcal{R}$ and $\mathcal{L}$ are, in principle, defined over continuous time inputs, we apply them at a discrete set of uniformly spaced time instances $t_{k}=t_{0}+k\delta t$ for $k=1,2,\ldots,K$, consistent with the temporal sampling of our data; see Eq. (8). To consider lags in the latent coordinate embeddings $\boldsymbol{y}(t_{k})$, we employ both linear and nonlinear autoregressive models, namely MVARs and LSTMs.
Let an MVAR model of order $w$ (i.e., with $w$ time lags) be expressed as:
$\boldsymbol{y}_{c}(t_{k})=\boldsymbol{A}_{0}+\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j})+\boldsymbol{\varepsilon}(t_{k}),$ (22)
where $\boldsymbol{y}_{c}(t_{k})$ is the latent variable vector, evaluated at discrete times $t_{k}$ for each of the $c=1,\ldots,C$ cases of different initial condition considered, $\boldsymbol{A}_{0}\in\mathbb{R}^{d}$ is the intercept vector and $\boldsymbol{A}_{j}\in\mathbb{R}^{d\times d}$ are the regression coefficient matrices for each lag $j$. The predictor vectors $\boldsymbol{y}_{c}(t_{k-j})\in\mathbb{R}^{d}$ include the $w$ past lag latent variables evaluated at discrete times $t_{k-1},\ldots,t_{k-w}$, while the error term $\boldsymbol{\varepsilon}(t_{k})\in\mathbb{R}^{d}$ is assumed to follow a multivariate white noise process:
$\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})]=\mathbf{0},\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{k})^{\top}]=\Sigma,\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{r})^{\top}]=\mathbf{0},\quad\text{for }t_{k}\neq t_{r},$ (23)
for every $k,r=1,\ldots,K$.
Given sequential (time series) data for $\{\boldsymbol{y}_{c}(t_{k})\}$, the unknown parameters of the MVAR in Eq. (22), i.e., the intercept vector $\boldsymbol{A}_{0}$ and the coefficient matrices $\boldsymbol{A}_{j}$, can be computed by the solution of the following least-squares problem:
$\operatorname*{argmin}_{\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}}\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right),$ (24)
where the MVAR loss function is:
$\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\sum_{c=1}^{C}\sum_{k=w+1}^{K}\Big{\lVert}\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big{\rVert}_{2}^{2},$ (25)
with the predictions of the MVAR model given by:
$\hat{\boldsymbol{y}}_{c}(t_{k})=\boldsymbol{A}_{0}+\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j}),$ (26)
for $c=1,\ldots,C$ and $k=1,\ldots,K$. The optimization problem in Eq. (24) can be solved using methods such as regularized least square methods; in particular, we used ridge regularization [81]. We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations. If the latent coordinate time series is non-stationary, pre-processing steps like differencing can be applied to achieve stationarity before solving the optimization problem. We highlight that MVAR model training further requires the selection of an optimal lag window size $w$, which determines the length of the time series history used to model the temporal dependencies. Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in Appendix D.
As a non-linear autoregressive model alternative to linear MVARs, we further consider LSTM networks, a type of Recurrent Neural Networks (RNNs) characterized by gating mechanisms, which allow the network to retain and propagate information over long sequences selectively, but also effectively handle vanishing gradients; a common challenge in traditional RNNs [82, 83, 84, 85]. Let an LSTM model with past $w$ latent states $\boldsymbol{y}_{c}(t_{k-j})$ for $j=1,\ldots,w$ be expressed as:
$(\mathbf{h}_{\tau},\mathbf{c}_{\tau})$ $=F_{\mathrm{LSTM}}\!\big{(}\boldsymbol{y}_{c}(\tau),\mathbf{h}_{\tau-1},\mathbf{c}_{\tau-1};\boldsymbol{\theta}\big{)},\quad\text{for}\ \tau=t_{k-w},\ldots,t_{k-1},$ (27)
$\hat{\boldsymbol{y}}_{c}(t_{k})$ $=\mathbf{W}_{y}\,\mathbf{h}_{t_{k-1}}+\mathbf{b}_{y},$ (28)
where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the, evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$. The gating mechanism of the LSTM is included in Eq. (27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for the conciseness of the presentation; a vanilla LSTM including 4 gates, contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
Similarly to the sequence-to-one learning of MVARs, given sequential data for $\boldsymbol{y}_{c}(t_{k})$ for a fixed lag window of width $w$, we train the LSTM model by solving the optimization problem:
$\operatorname*{argmin}_{\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}}\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right),$ (29)
where the LSTM loss function is the mean squared error (MSE) between the true value of the latent variable vector $\boldsymbol{y}_{c}(t_{k})$ and the predicted one $\hat{\boldsymbol{y}}_{c}(t_{k})$, as provided by Eqs. (27, 28), that is:
$\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big{\lVert}\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big{\rVert}^{2}.$ (30)
To solve the optimization problem in Eq. (30), one typically uses gradient-based methods; in particular, we employed the Adam optimizer [86], which is a stochastic variant of gradient descent, known for its adaptive learning-rate capabilities.
We highlight that while MVAR models are theoretically simple and efficient, their strict reliance on stationary data poses a significant limitation, requiring explicit pre‑processing that may distort dynamics or lose information [87]. In contrast, LSTMs as dynamical models can inherently handle non‑stationarity through their gated memory cells and adaptive hidden states. This allows LSTM to learn complex, long‑range temporal dependencies and nonlinear relationships directly from raw data. However, this flexibility comes at the cost of needing much larger datasets, greater computational resources, and often lacking interpretability compared to MVAR models.
Finally, the predicted latent coordinates $\hat{\boldsymbol{y}}_{c}(t_{k})$ obtained from the trained MVARs or LSTMs models are subsequently mapped back to the high-dimensional density space using the lifting operator defined in Eq. (21).
3 Case study and methodology implementation
3.1 Configuration
The proposed methodology is employed on a crowd dynamics configuration where pedestrians are moving from left to right in a corridor of dimensions $48\,\text{m}\times 12\,\text{m}$ ($x$: length, $y$: width) in the presence of an obstacle, as shown in Fig. 2.
Figure 2: The case study: The SFM simulating pedestrians moving in a corridor with an obstacle.
The square obstacle of area $12.96\,\text{m}^{2}$ is centered at $(25.8\,\text{m},1.8\,\text{m})$, with the pedestrians navigating around it towards reaching the corridor’s right end. Hence, the computational domain is $\Omega=\left([0,48]\times[0,12]\right)\backslash\left([24,27.6]\times[0,3.6]\right)$.
For modelling the pedestrians’ movement, we use the SFM [1] as a microscopic agent-based simulator. SFM falls within the general form of Eqs. (1, 2) and the idea behind it is that the motion of individuals is governed by “virtual forces”. These forces include acceleration toward the desired destination, as well as repulsive forces “exerted” by obstacles, borders, and nearby individuals to maintain a safe distance. Additionally, one may also consider an attractive force due to mimetic behavior, which can be observed in scenarios such as emergency evacuations or collective movement toward objects of common interest [1]. The derivation of SFM is based on the Langevin-type equation that contains a frictional term, a deterministic force term, and a random force term. In the SFM, these terms translate into a goal-directed force (playing the role of friction), social and repulsive (from obstacles) deterministic forces, and a random term, reading:
$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{i,j}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{i,B}(\bm{x}_{i})}_{\text{repulsive force}}+\mathbf{F}_{rand},$ (31)
where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time. Here, we avoid the inclusion of random forces, since we are interested in the macroscopic behavior that often emerges from the deterministic pedestrian interactions [1]. Details on the SFM forces included in Eq. (31) and the related parameters that were considered in this work are provided in Appendix A.
We simulate a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see in the Appendix B). Pedestrians are directed to move from left to right by setting their initial target at coordinates $(x,y)=(25\,\text{m},\,9.6\,\text{m})$, corresponding to the midpoint of the corridor in the horizontal direction and 80% of its vertical extent. Upon reaching this intermediate waypoint, their final destination is updated to $(x,y)=(48\,\text{m},\,4.2\,\text{m})$, which lies at the right end of the corridor and at 35% of its height. To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents that reach the right boundary at $x=48\,\text{m}$, at the left boundary at $x=0\,\text{m}$.
Pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator. Such configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.
3.2 Data generation of microscopic distributions
We generate synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy. Each simulation has a time span of $275\,\text{s}$. To reduce computational storage costs while retaining essential temporal features, we subsample the resulting pedestrian trajectories at intervals of $\delta t=0.25\,\text{s}$. As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios. In particular, we consider $C=10$ different initializations with zero velocities and position distributions including: (i) uniform distribution to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) piecewise linear distributions with higher densities near the wall boundaries to simulate characteristic wall-avoidance behavior, (iv) sinusoidal distributions to generate periodic macroscopic behavior, and (v) double Gaussian distributions to capture the formation of two distinct crowds. Details on the form and parameters of the above initializations are provided in Appendix B, distinguishing between those used for training and testing the proposed framework in Tables B.1 and B.2, respectively. The resulting datasets are comprised of 10 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1100}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1100$ timesteps are sampled from each trajectory. Thus, the total datasets overall cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(10\cdot 1100)}$ used for training and testing, respectively.
3.3 Extraction of continuous density fields
For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1. To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behaviors, while minimizing artificial grid effects. For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq. (6) tuned with bandwidths $\mathbf{H}=diag(h_{x},h_{y})=(3,2)$ for suppressing noise and preserving features of the microscopic data, while ensuring that local structures remain clear without introducing spurious artifacts. To account for the physical obstacle in $\Omega$, a binary mask is applied to explicitly zero out density estimation on the cell centroids of the obstacle geometry. To account for the periodic boundary conditions, we further perform domain augmentation for the KDE support. All the details regarding KDE implementation are provided in Appendix C.
Using the above procedure, we derive the density fields $\rho^{(c)}\big{(}x_{i},y_{j},t_{k}\big{)}$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing data sets $\mathcal{X}_{ts}$. For conforming to the assumptions of Proposition 1, we finally normalize each density field following Eq. (19), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq. (8), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=10\cdot 1100$.
3.4 POD for the restriction and lifting operators
For constructing the restriction and lifting operators in Eqs. (7) and (21) respectively, we employ POD on the training dataset $X_{tr}$ to determine the latent dimension $d$ and the POD basis $U_{d}$. Using the economy-size SVD (which reduces computational complexity in thin matrices while maintaining exact reconstruction capability), we determine $d$ by the best low-rank approximation of $X_{tr}$ retaining $99\%$ of the energy; thus, computing the minimum number $d$ satisfying the criterion:
$E_{d}=\dfrac{\sum_{i=1}^{d}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq 0.99$ (32)
where $r=\min(n_{c},n_{t})=1600$ is the rank of $X_{tr}$, and $\sigma_{i}>0$ for $i=1,\ldots,r$ are the singular values in descending order. Next, using the first $d$ left singular vectors of $X_{tr}$, we form $U_{d}$ for completing the construction of the restriction and lifting operators in Eqs. (7, 21). To assess the accuracy of restriction and lifting operators, we evaluate the POD reconstruction error on the training and testing data sets by the relative $L_{2}$ error:
$e^{2,rec}_{k}=\dfrac{\lVert x(t_{k})-\mathcal{L}(\mathcal{R}(x(t_{k})))\rVert_{2}}{\lVert x(t_{k})\rVert_{2}},$ (33)
where $x(t_{k})\in X_{tr},X_{ts}$ with $k=1,\ldots,K$ is the normalized density defined in Eq. (19).
3.5 Autoregressive ROMs and their training
Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct autoregressive ROMs for learning the dynamics in the latent space. For this purpose, we learn linear MVARs and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$. To conform to this sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq. (7) on the training data $X_{tr}$.
As a preliminary step, we first tune the lag window ww–within a bound of $w_{\text{max}}=100$–using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [88, 89, 90], the details of which are described in Appendix D. Both criteria are likelihood-based, with AIC tending to select larger windows than BIC. We note that BIC and AIC are well-suited for MVARs models since they assume stationarity of the data and white noise residuals. For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey-Fuller (ADF) test [91] with a threshold of significance level $p<0.01$. Then, we solve the least-squares optimization problem in Eq. (24) using ridge regression with a regularization parameter $\lambda=10^{-6}$ to compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq. (22).
For the LSTM model, we consider a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq. (28). To train the LSTM model, we solve the optimization problem in Eq. (29) using the Adam optimizer [86] with automatic differentiation for the gradients of the loss function, enabled through MATLAB’s Deep Learning Toolbox [92]. The model’s parameters are initialized with a Glorot uniform initialization [93], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints. The maximum number of epochs was set to $100$ as no further improvement was achieved using more epochs. To further avoid overfitting of the LSTM, we split the training data into training and validation sets using a 80-20% ratio, and shuffle each data set without breaking the internal temporal structure of the sequences. We compute the loss function value on the validation set for every epoch and stop the training after 5 consecutive validation checks with no improvement.
To assess the convergence of both MVARs and LSTMs training, we report the values of the loss functions in Eqs. (25) and (30), respectively, at the end of training. Since the MVARs loss function is based on $L_{2}$ error, we further report the MSE:
$MSE_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\rVert^{2},$ (34)
where $\hat{\boldsymbol{y}}_{c}(t_{k})$ is the prediction in Eq. (26), to enable direct comparison with the MSE-based loss function of LSTMs. Additionally, we report the computational times required for training both ROMs; we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the ridge regression employed for solving the related optimization problem.
3.6 Accuracy assessment of the proposed framework
Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $c=10$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details. In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$. Then, following the proposed restrict-evolve with ROM-lift framework in Eq. (5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities; i.e., computing:
$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(\hat{x}^{(c)}(t_{k-1})),\ldots,\mathcal{R}(\hat{x}^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (35)
where $\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators in Eqs. (7) and (21) and $\Phi_{\text{ROM}}$ is the autoregressive ROM with parameters $\mathbf{p}$ that can be either the trained MVARs or LSTMs model. Notice that while Eq. (35) could consider the restricted values of the normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, it instead considers the predicted normalized densities; thus, evaluating the proposed framework in a closed-loop (autoregressive) manner. This allows us to feed subsequent predictions recursively into Eq. (35) to advance the forecast over time.
To quantify the prediction accuracy, we compute the relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ per case $c$ and time step $k$ as:
$e^{1,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{1}}{\lVert x^{(c)}(t_{k})\rVert_{1}},\quad e^{2,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{2}}{\lVert x^{(c)}(t_{k})\rVert_{2}},\quad e^{\infty,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{\infty}}{\lVert x^{(c)}(t_{k})\rVert_{\infty}}.$ (36)
We report these errors overall cases and time steps, also computing the 10-90 percentiles, and track down the evolution of these errors along time steps.
All numerical computations were executed on a PC equipped with an 11th Gen Intel(R) Core(TM) i7-1165G7 2.80 GHz CPU and 16 GB of system memory.
4 Numerical results
Here, we present the results of the proposed framework on predicting pedestrian flow following the configuration of a rectangular corridor with an obstacle described in Section 3.1. As already discussed in Section 3.2, we first collected pedestrian trajectory data using the SFM microscopic agent-based simulator under a wide range of initial conditions. The resulting training dataset $\mathcal{X}_{tr}$ is used to train our framework, while the testing dataset $\mathcal{X}_{ts}$ is used to assess its efficiency and prediction accuracy.
At the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3. Next, we employed POD on $\mathcal{X}_{tr}$ to discover the latent space dimension; see Section 3.4. Following the criterion in Eq. (32), the minimum latent dimension required for retaining more than 99% of the energy in the dataset is $d=13$ modes, as shown in Fig. 3a).
(a) POD cumulative energy
(b) POD reconstruction error
Figure 3: Accuracy of the restriction and lifting operators in Eqs. (7) and (21). Panel a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set. Panel b) depicts the average (over the $c=10$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq. (33) with $d=13$ POD modes, in time over the training and testing datasets.
Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs. (7) and (21). To assess their accuracy, we computed the relative $L_{2}$ reconstruction error $e^{2,rec}_{k}$ in Eq. (33) across the $k=1,\ldots,K$ time steps of each trajectory. Figure 3b) displays the average relative error overall $c=10$ cases with different initial conditions included in either the training or testing datasets. While in the beginning, the reconstruction error is more than 10%, after the first $100$ time steps, it reduces to 5% for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$. We note that this level of reconstruction error serves as a baseline for the proposed framework, since the timeseries forecasting with ROMs in the latent space introduces additional approximation errors.
Having constructed the restriction and lifting operators (second and fourth step of the framework), we next train MVARs and LSTMs autoregressive ROMs to learn the latent dynamics; see Section 2.2.3. First, we obtained the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ from the density profiles of the training set, using the restriction operator. We then determined the lag window, using the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas the AIC favored (a wider, as expected) $w=9$. Both windows were used for training both MVARs and LSTMs models, denoted from now on as MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models. For performing sequence-to-one learning, we further divided the restricted training data set into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
For the MVARs training, we first verified that the latent dynamics datasets $Y_{f},Y_{t}$ constitute stationary time series. Using the ADF test, we confirmed that all $d$ variables are stationary processes with a significance level of $p=0.001$. We next trained all MVAR and LSTM models using algorithms, hyperparameters and stopping criteria discussed in detail in Section 3.5. The training results are shown in Table 1 for the four ROMs, including the loss function values attained at the end of training and the computational time (in seconds) required.
Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models. Loss function values at the end of training for MVARs $\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{c})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)$ in Eq. (25) and for LSTMs $\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{c})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)$ in Eq. (30), along with computational times (in seconds) required for training. For MVARs, the MSE in Eq. (34) is reported for comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s) $\downarrow$
MVAR(4) $1.66\times 10^{-3}$ $1.96\times 10^{-11}$ - 0.0055
MVAR(9) $1.01\times 10^{-3}$ $7.38\times 10^{-12}$ - 0.042
LSTM(4) - - $8.64\times 10^{-7}$ 180
LSTM(9) - - $3.51\times 10^{-7}$ 443
Since the MVARs loss function is $L_{2}$ based, we further report the MSE, computed in Eq. (34), for MVARs to allow comparison with the MSE-based loss of the LSTMs. Table 1 indicates that all the optimization problems converge, as evident from the low loss function values attained. MVARs are able to reach much smaller loss function values than LSTMs, providing a hint for their better accuracy. As expected, LSTMs require significantly higher computational time for training in comparison to MVARs. Table 1 further shows that the training of both types of ROM converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
Having constructed and trained all components of our framework, we now evaluate its end-to-end performance, using the forecasting closed-loop time-stepper in Eq. (35) for long-horizon forecasting. In particular, we quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ for the $c$, unseen in training, cases–of different initial conditions–included in the testing set $X_{ts}$ across $k$ time steps. Table 2 provides a summary of the relative errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ for all four ROMs, including the median and the 10-90 percentiles, overall cases and time steps of the testing set.
Table 2: End-to-end numerical accuracy of the closed-loop framework across the testing set, using the trained MVARs and LSTMs models for the latent dynamics. The relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors of the high-dimensional density space, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) are reported, for the four MVAR(4), MVAR(9), LSTM(4), LSTM(9) latent dynamics models; median and 10-90 percentiles (in parenthesis) are shown overall $C=10$ cases and $K=1091$ time steps.
Model $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
MVAR(4) 0.3599 (0.2408, 0.4535) 0.2536 (0.1609, 0.3098) 0.2251 (0.1361, 0.3435)
MVAR(9) 0.3783 (0.2865, 0.4655) 0.2765 (0.2040, 0.3367) 0.2569 (0.1935, 0.3330)
LSTM(4) 0.6323 (0.4313, 0.7941) 0.4160 (0.2889, 0.5269) 0.3372 (0.2532, 0.5037)
LSTM(9) 0.4646 (0.3172, 0.6820) 0.3185 (0.2151, 0.4626) 0.3010 (0.1993, 0.4393)
As shown in Table 2, the prediction accuracy of the closed-loop time stepper with the use of MVAR models outperforms the one with LSTMs, demonstrating robust generalization with tighter 10-90 percentile ranges. In addition, while MVARs with shorter lag demonstrate slightly higher approximation accuracy, LSTMs significantly improve with higher lags, albeit remaining inferior to MVARs. We hereby note that, considering the baseline 5-10% $L_{2}$ relative error of POD reconstruction (see Fig. 3), the $L_{2}$ errors are 3-10 times higher for the end-to-end closed-loop framework, underscoring the compounding effect of autoregressive prediction.
To highlight the robust performance of the proposed closed-loop time-stepping framework over time, we further depict the relative $L_{2}$ and $L_{\infty}$ errors, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in time, across the multiple cases of different, unseen initial conditions in Figs. 4 and 5 for MVARs and LSTMs, respectively. In particular, Figs. 4,5 show the average relative errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10-90 percentile ranges, over the $c=10$ cases considered, across time steps $k=1,\ldots,1100$.
(a) Relative Errors with MVAR(4)
(b) Relative Errors with MVAR(9)
Figure 4: End-to-end numerical accuracy of the closed-loop framework across the testing set over time, using the trained MVAR models. The relative $L_{2}$ (red) and $L_{\infty}$ (green) errors of the high-dimensional density space, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) are displayed for the MVAR(4) (panel a) and MVAR(9) (panel b) latent dynamics models; average (solid) and 10-90 percentiles (dashed) are shown overall $C=10$ cases per time step.
Figures 4a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both indicating that, in spite of the long-time prediction, the relative error remains bounded in values below 30%; 5-10% is the baseline error of POD reconstruction in Fig. 3. In addition, the tight 10-90 percentile ranges indicate good generalization. Although both models exhibit comparable error profiles, the MVAR(4) model demonstrates slightly improved accuracy while using fewer parameters, as also reported in Table 2.
Similar bounded results are reported in Fig. 5 (a,b), depicting the employment of LSTM(4) and LSTM(9) models for long-time predictions.
(a) Relative Errors with LSTM(4)
(b) Relative Errors with LSTM(9)
Figure 5: End-to-end numerical accuracy of the closed-loop framework across the testing set over time, using the trained LSTM models. The relative $L_{2}$ (red) and $L_{\infty}$ (green) errors of the high-dimensional density space, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (36) are displayed for the LSTM(4) (panel a) and LSTM(9) (panel b) latent dynamics models; average (solid) and 10-90 percentiles (dashed) are shown overall $C=10$ cases per time step.
As expected by the results in Table 2, the use of LSTMs provides less accurate predictions in comparison to MVARs. However, LSTMs provide tighter 10-90 percentile ranges towards the end of time integration, indicating a more robust behavior in comparison to MVARs. In addition, as also reported in Table 2, a wider lag window makes the scheme with LSTMs more accurate, while that of MVARs less accurate. Nevertheless, when comparing across model classes, the MVAR(4) model achieves slightly lower relative errors overall, while also requiring significantly fewer trainable parameters and hyperparameters. This makes MVAR(4) not only more accurate in this setting but also more computationally efficient.
For completeness, we provide a comparison of the “ground truth” normalized density fields $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ derived based on the closed-loop time-stepper in Eq. (35) for a particular case, in the testing data set, of an initially spread pedestrian group; Gaussian initialization corresponding to the 6th row of Table B.2. Figure 6 displays such a comparison for the MVAR(4) model in four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930) where the pedestrian crowd does not pass through the corridor boundary; similar comparisons are provided for the MVAR(9), LSTM(4) and LSTM(9) models in Figs. E.1, E.2 and E.3 of Appendix E, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the MVAR(4) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
As shown in Fig. 6, the predicted density fields preserve the global structure and propagation direction observed in the “ground truth” density, demonstrating the framework’s ability to capture dominant crowd dynamics and to deliver meaningful long-horizon predictions. The predicted density fields additionally succeed in correctly respecting the obstacle region. Comparison with the other (less accurate) models in Figs. E.1, E.2, E.3 reveals that while the above two crucial for crowd flow characteristics are preserved, the LSTM(4) model (the least accurate one) tends to underestimate peak density values, introduce artificial density in low density regions (for coping with conservation of mass), and exhibit greater dissipation in regions of high density; these erroneous behaviors relax as more accurate models are used, i.e., the LSTM(9) and the MVAR(9) ones.
As shown, the proposed framework provides a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics. In particular, for the benchmark case of $100$ pedestrians moving in the corridor for more than 275s275\ $\text${s}, the conventional numerical integration requires a computational time of 20s20\ $\text${s} for the microscopic SFM simulator simulation plus 30s30\ $\text${s} for the extraction of the density fields (in a coarser time discretization), summing up to a total of 50s50\ $\text${s}. In contrast, our framework requires just $0.0063$, 0.0161s0.0161\ $\text${s} for the MVARs models and $6.75$, 6.77s6.77\ $\text${s} for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively. Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under 0.1s0.1\ $\text${s} (as shown in Table 1). The MVARs variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts. These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches unfeasible.
5 Conclusion
In this work, we present a four-step data-driven framework that combines manifold learning and machine learning to construct a discrete evolution operator for emergent crowd dynamics. Our framework provides a highly computationally efficient and systematic approach for learning accurate ROMs for spatio-temporal emergent crowd dynamics. By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs—sidestepping the prohibitive computational cost of learning PDEs via for example convolutional neural networks (CNNs) or other high‐dimensional ML models such as DNNs. Working in latent space simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics.
Thus, our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution. The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction. As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics without sacrificing accuracy or stability, making them an ideal tool for large‐scale, real-world crowd‐dynamics analysis and control.
Finally, we note that linear MVARs exhibit a better performance in terms of the median relative $L_{1}$, $_{2}, LL_{\infty}$ errors when compared to LSTMs, thus offering a significantly lower complexity in their training and greater interpretability. This is in line with the results presented in other works [56, 94, 63, 57] which utilize PCA-based delay-coordinate embeddings to construct linear ROMs using in latent spaces to approximate complex nonlinear or chaotic dynamics.
Appendix A Social Force Model
Here, we provide the detailed expressions of the virtual forces included in the SFM Eq. (31) and the set of parameters selected for the microscopic simulations.
As a reminder, we consider the SFM including a goal-directed, a social, and a repulsive force term (the random force is neglected), reading:
$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{i,j}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{i,B}(\bm{x}_{i})}_{\text{repulsive force}},$ (A.1)
where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time. The social force term in Eq. (A.1) introduces a short-range repulsive interaction between pedestrians, to prevent collisions. The social force between pedestrians $i$ and $j$ is given by:
$\mathbf{F}_{ij}(\bm{x}_{i},\bm{x}_{j})=A_{i}\exp\left(\frac{r_{ij}-d_{ij}}{B_{i}}\right)\bm{n}_{ij}+k\,g(r_{ij}-d_{ij})\,\bm{n}_{ij}+\kappa\,g(r_{ij}-d_{ij})\,\Delta v_{ij}^{\mathrm{tang}}\,\bm{t}_{ij},$ (A.2)
where $A_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction, $d_{ij}=\|\bm{x}_{i}-\bm{x}_{j}\|$ is the distance between pedestrians, $r_{i,j}\in\mathbb{R}$ is the sum of their radii (minimal comfortable distance) [64], $B\in\mathbb{R}^{+}$ is the parameter controlling the spatial decay rate of the repulsive force, $\bm{n}_{ij}=\big{(}n^{1}_{i,j},n^{2}_{i,j}\big{)}=\frac{\bm{x}_{i}-\bm{x}_{j}}{d_{ij}}$ is the unit vector pointing from pedestrian $j$ to pedestrian $i$, $g(x)$ is a function which ensures that physical forces are only active when pedestrians overlap, i.e., when $(d_{i,j}<r_{i,j})$. Moreover, the constant $k>0$ governs the strength of the linear contact (body) force, which causes repulsion between pedestrians when they are in contact. The third term is a sliding friction term that involves a coefficient $\kappa>0$, which scales with the tangential component of the relative velocity $\Delta v_{ij}^{\mathrm{tang}}$; this component is projected onto the unit tangential vector $\mathbf{t}_{i,j}=\big{(}-n^{2}_{i,j},n^{1}_{i,j}\big{)}$. The repulsive force in Eq. (A.1), accounting for the repulsion that a pedestrian might experience due to static objects $B$, follows a similar, to Eq. (A.2), form, reading:
$\mathbf{F}_{iB}(\bm{x}_{i})=C_{i}\exp\left(-\frac{d_{iB}}{D_{i}}\right)\bm{n}_{iB}+k\,g(r_{i}-d_{iB})\,\bm{n}_{iB}.$ (A.3)
Here, $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle and $D_{i}$ is its corresponding decay parameter. Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exert dynamic resistance.
For our simulations, the model parameters were calibrated based on established values from previous studies [1, 64]. These works describe pedestrians with a standard average mass of 8080kg, for which a relaxation time $\tau=0.5\,\text{s}$ provides adequate behavior under normal conditions. The complete set of parameters governing repulsive potentials, including agent-agent and agent-obstacle interactions, along with the simulation environment specifications, are detailed in Table A.1. This parameter selection, grounded in empirical crowd studies, enables the simulation of diverse scenarios, including the present case study while maintaining consistency with validated bottleneck evacuations scenarios, and even though the parameters should vary individually, in the literature the parameters are often chosen equally for all pedestrians to maintain a reasonable number of parameters calibration [64].
Table A.1: Simulation parameters.
Simulation Parameters Value
Number of Pedestrians 100
Corridor length ($L_{x}$) $48\,\text{m}$
Corridor width ($L_{y}$) $12\,\text{m}$
Time Step $\delta_{t}$ $2.5\times 10^{-2}\,\text{s}$
Simulation time $275\,\text{s}$
Repulsion coefficient ($k$) $1.2\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
Sliding friction coefficient ($\kappa$) $2.4\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
Interaction strength ($A_{i}$) $2\times 10^{3}\,\text{N}$
Interaction range ($B_{i}$) $0.08\,\text{m}$
Interaction strength ($C_{i}$) $2\times 10^{3}\,\text{N}$
Interaction range ($D_{i}$) $0.08\,\text{m}$
Mass $80\,\text{kg}$
Relaxation time ($\tau$) $0.5\,\text{s}$
Pedestrian radius $0.2\,\text{m}$
Initial velocity in $x$-direction ($v_{x}$) $0\,\text{m/s}$
Initial velocity in $y$-direction ($v_{y}$) $0\,\text{m/s}$
Obstacle width $3.6\,\text{m}$
Obstacle height $3.6\,\text{m}$
$x$-coordinate of center $25.8\,\text{m}$
$y$-coordinate of center $1.8\,\text{m}$
Appendix B Initial Conditions of Microscopic Distributions
Here, we describe the initial conditions that we considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested. We considered zero initial velocities for all pedestrians, while we used the following distributions for initializing the position $\bm{x}_{i}=(x_{i},y_{i})\in\Omega$ of the $i$-th pedestrian.
  • Positions are sampled from a two dimensional normal distribution with mean $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
    $p(x_{i},y_{i})=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\left(-\frac{(x_{i}-\mu_{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y_{i}-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right),$ (B.1)
  • Uniform density distribution: Positions are sampled uniformly over the domain $\Omega$
    $p(x_{i},y_{i})=\mathcal{U}\left([x_{min},x_{max}]\times[y_{min},y_{max}]\right).$ (B.2)
  • Cosine distribution;
    $p(x_{i},y_{i})=\frac{1}{\pi\sigma_{x}\sigma_{y}}\cos\left(\frac{x_{i}-\mu_{x}}{\sigma_{x}}\right)\cos\left(\frac{y_{i}-\mu_{y}}{\sigma_{y}}\right).$ (B.3)
  • Piecewise linear density distribution:
    To initialize particle positions in a rectangular domain $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$, we assume a uniform distribution in $x$ and a non-uniform, piecewise linear distribution in $y$. This approach allows control over the particle concentration along the vertical axis.
    Particles are sampled uniformly in the $x$-direction over the interval $x_{\min},x_{\max}]$. In the $y$-direction, particle positions are sampled from a custom piecewise linear distribution designed to increase density near the top and bottom boundaries. Specifically, we use a triangular distribution defined over the interval $[0,12]$ by
    $p(y_{i})=\begin{cases}\frac{2}{12}\left(1-\frac{y_{i}}{6}\right),&0\leq y_{i}\leq 6,\\ \frac{2}{12}\left(\frac{y_{i}-6}{6}\right),&6<y_{i}\leq 12,\end{cases}$ (B.4)
    which has peaks at $y=0$ and $y=12$, and a minimum at $y=6$.
    To sample $y$-coordinates from this distribution, we use inverse transform sampling. First, a uniform random number $u\sim\mathcal{U}(0,1)$ is generated. If $u\leq 0.5$, the corresponding $y$ value is computed by solving $u=\frac{y}{12}(2-\frac{y}{6})$, yielding
    $y=6\left(1-\sqrt{1-2u}\right).$ (B.5)
    If $u>0.5$, we solve $u=\frac{1}{2}+\frac{(y-6)^{2}}{72}$, resulting in
    $y=6+\sqrt{72(u-0.5)}.$ (B.6)
    This method produces particle positions that are uniformly distributed along the horizontal axis and more concentrated near the top and bottom edges along the vertical axis.
For generating the training/testing data set, we used a wide range of initial conditions, based on the described distributions;in particular, for the gaussian sampling, the means and standard deviations were selected relative to the domain size to ensure spatial coverage while avoiding boundary artifacts. Specifically, the means were placed at the geometric center or fractional positions within the domain (e.g., $0.5$ or $0.25$ of the domain width), while standard deviations were set as fractions of the domain dimensions (e.g., one-fifth or one-half of the width or height). This ensures that the pedestrian positions lie within the domain, while allowing for spatial clustering or multimodal patterns. The parameters were chosen heuristically to reflect plausible crowd configurations and to introduce variety in the training and testing datasets. the ones used to construct the training are shown in Table B.1, while the ones for the testing data set are shown in Table B.2.
Table B.1: Microscopic distribution configurations and parameters for training the LSTM and MVAR models.
Distribution Type Parameters
Data set 1: Gaussian $\begin{aligned} &\mu_{x}=7.5,\;\mu_{y}=5.0,\\ &\sigma_{x}=2.2,\;\sigma_{y}=2.0\end{aligned}$
Data set 2: Uniform $\mathcal{U}\left([2,15]\times[3,6]\right)$
Data set 3: Gaussian $\begin{aligned} &\mu_{x}=8.4,\;\mu_{y}=4.0,\\ &\sigma_{x}=1.8,\;\sigma_{y}=1.3\end{aligned}$
Data set 4: Gaussian $\begin{aligned} &\mu_{x}=12.5,\;\mu_{y}=4.0,\\ &\sigma_{x}=7,\;\sigma_{y}=1.1\end{aligned}$
Data set 5: Double Gaussian $\begin{aligned} &\mu_{x1}=17.3,\;\mu_{x2}=13.6,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.7,\\ &\sigma_{y}=1.2;\;\mu_{y}=7.0\end{aligned}$
Data set 6: Double Gaussian $\begin{aligned} &\mu_{x1}=19.7,\;\mu_{x2}=13.7,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.62,\\ &\sigma_{y}=4.0;\;\mu_{y}=2.33\end{aligned}$
Data set 7: Double Gaussian $\begin{aligned} &\mu_{x1}=32.2,\;\mu_{x2}=27.75,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.35,\\ &\sigma_{y}=4.0;\;\mu_{y}=1.0\end{aligned}$
Data set 8: Piecewise Linear $x_{i}\sim\mathcal{U}\left(30,47\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
Data set 9: Piecewise Linear $x_{i}\sim\mathcal{U}\left(15,35\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
Data set 10: Cosine $\begin{aligned} &\mu_{x}=25,\;\mu_{y}=3.5,\\ &L_{x}=10,\;L_{y}=2.5\end{aligned}$
Table B.2: Microscopic distribution configurations and parameters for testing the LSTM and MVAR models.
Configuration Parameters
Dataset 1: Double Gaussian $\begin{aligned} &\mu_{x1}=41.2,\;\mu_{x2}=33.7,\\ &\sigma_{x1}=\sigma_{x2}=0.5,\\ &\sigma_{y}=2.0,\;\mu_{y}=9.0\end{aligned}$
Dataset 2: Double Gaussian $\begin{aligned} &\mu_{x1}=11.7,\;\mu_{x2}=5.2,\\ &\sigma_{x1}=\sigma_{x2}=0.4,\\ &\sigma_{y}=1.0,\;\mu_{y}=9.0\end{aligned}$
Dataset 3: Double Gaussian $\begin{aligned} &\mu_{x1}=42.5,\;\mu_{x2}=37.5,\\ &\sigma_{x1}=\sigma_{x2}=0.3,\\ &\sigma_{y}=2.0,\;\mu_{y}=5.0\end{aligned}$
Dataset 4: Gaussian $\begin{aligned} &\mu_{x}=36.0,\;\mu_{y}=4.0,\\ &\sigma_{x}=2.4,\;\sigma_{y}=1.5\end{aligned}$
Dataset 5: Gaussian $\begin{aligned} &\mu_{x}=23.0,\;\mu_{y}=8.5,\\ &\sigma_{x}=2.0,\;\sigma_{y}=2.5\end{aligned}$
Dataset 6:Gaussian $\begin{aligned} &\mu_{x}=7.85,\;\mu_{y}=8.0,\\ &\sigma_{x}=4.3,\;\sigma_{y}=2.0\end{aligned}$
Dataset 7: Cosine $\begin{aligned} &\mu_{x}=40,\;\mu_{y}=8.0,\\ &L_{x}=5,\;L_{y}=2.0\end{aligned}$
Dataset 8: Uniform $\mathcal{U}\left([10,20]\times[4,9]\right)$
Dataset 9: Piecewise Linear $x_{i}\sim\mathcal{U}\left(30,44\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
Dataset 10: Cosine $\begin{aligned} &\mu_{x}=39,\;\mu_{y}=6.5,\\ &L_{x}=7,\;L_{y}=2.5\end{aligned}$
Appendix C Kernel Density Estimation
Here, we provide additional details for the first step of the proposed framework, regarding the extraction of continuous density profiles from discrete pedestrians’ positions. As already discussed in Section 2.2.1, the Kernel Density Estimation (KDE) is used for this task, following the general form in Eq. (6). Here, we elaborate on the kernel selection, bandwidth tuning, and adjustments for boundaries and obstacles.
The choice of kernel function directly influences the smoothness and continuity of the density estimate. Common options include the uniform, triangular, Epanechnikov, and Gaussian kernels [95, 96]. The Epanechnikov kernel is theoretically optimal for minimizing mean integrated squared error (MISE) [97], but its bounded support can limit flexibility in practice. In contrast, the Gaussian kernel is widely preferred due to its smoothness, infinite support, and analytical tractability. Its multivariate form is given by:
$K_{h}(\mathbf{x}-\mathbf{x}_{i}(t))=\frac{1}{2\pi\det(\mathbf{H})^{1/2}}\exp\left(-\tfrac{1}{2}(\mathbf{x}-\mathbf{x}_{i}(t))^{\top}\mathbf{H}^{-1}(\mathbf{x}-\mathbf{x}_{i}(t))\right),$ (C.1)
where $\mathbf{H}$ is the symmetric positive-definite bandwidth matrix, $\mathbf{x}\in\mathbb{R}^{2}$ is the spatial domain and $\mathbf{x}_{i}(t)\in\mathbb{R}^{2}$ is the position of the $i$-th pedestrian.
The choice of the bandwidth matrix, which in this case takes the form $\mathbf{H}=\operatorname{diag}(h_{x},h_{y})$, plays a central role in the KDE estimate. Small values of $h_{x}/h_{y}$ result in undersmoothing and high variance, while large values can oversmooth the density and obscure relevant features. Silverman’s rule of thumb [95], originally formulated for univariate distributions, extends naturally to multivariate settings, indicating an optimal bandwidth:
$h_{i}=\left(\frac{4}{N(d+2)}\right)^{1/(d+4)}\sigma_{i}\quad(i=x,y).$ (C.2)
where $N$ is the number of data points, $d$ is the number of spatial dimensions, in this case $d=2$ and $\sigma_{i}$ is the standard deviations of the pedestrian positions. However, in our case, Silverman’s rule resulted in wide bandwidths, which did not allow us to consider sufficient detail in the density profiles. We thus tuned $h_{x}/h_{y}$ to the values reported in Section 2.2.1.
The standard symmetric kernels are known to introduce bias near domain boundaries, when handling cases where periodic boundary conditions are imposed, such as in the $x$-direction of pedestrian flow in our configuration. These artifacts arise because the kernel’s support extends beyond the domain, failing to reflect the periodic structure of the system. To mitigate these artifacts, we define an extended support for the KDE in the $x$-direction. Let the physical domain boundaries be $x=a$ and $x=b$, and introduce a small displacement $\Delta x=\frac{b-a}{n}$, where $n$ defines the extension of the domain along $x$-direction; here, we tuned $n=5$. We then define the extended domain as:
$\Omega^{\text{ext}}=[a-\Delta x,b+\Delta x]\times[c,d],$ (C.3)
while the support in the $y$-direction remains unbounded, i.e., $\Omega_{y}=(-\infty,\infty)$. To enforce continuity and reduce boundary-induced artifacts in the $x$-direction, we further augment the observed positions with “ghost” particles near the domain edges. Let $\epsilon=\Delta x$ define a threshold for proximity to the boundary. For each position $\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at
$\mathbf{x}_{i}^{\text{aug1}}(t)=(x_{i}(t)-(b-a),y_{i}(t)).$
Similarly, for every $\mathbf{x}_{i}(t)$ such that $x_{i}(t)\geq b-\epsilon$, we introduce a “mirrored” pedestrians at
$\mathbf{x}_{i}^{\text{aug2}}(t)=(x_{i}(t)+(b-a),y_{i}(t)).$
The resulting augmented set of pedestrian positions used for KDE is then given by
$\{\mathbf{x}_{i}^{\prime}(t)\}=\{\mathbf{x}_{i}(t)\}\cup\{\mathbf{x}_{i}^{\text{aug1}}(t)\}\cup\{\mathbf{x}_{i}^{\text{aug2}}(t)\}.$ (C.4)
A pseudo-code for the estimation of the density in $\Omega\subset\mathbb{R}^{2}$ is given in Algorithm C.1.
Algorithm C.1 Kernel Density Estimation (KDE) for extracting density fields from pedestrian positions.
1:Number of pedestrians NN, computational domain $\Omega=[a,b]\times[c,d]$, spatial discretization $n_{x}$$n_{y}$, obstacle subdomain $\Gamma=[a_{o},b_{o}]\times[c_{o},d_{o}]\subset\Omega$ and pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\in\Omega\backslash\Gamma$
2:Set bandwidth $\mathbf{H}=\operatorname{diag}(h_{x},h_{y})$.
3:Set augmentation on $x$-direction $\Delta x=(b-a)/n$
4:Define extended support $\Omega^{\text{ext}}$ to handle periodic boundary conditions $\triangleright$ Use Eq. (C.3)
5:Discretize extended domain $\Omega^{\text{ext}}$ in $(n_{x}+2n_{x}/n)\times n_{y}$ cells
6:Introduce “mirrored” pedestrians to form $\{\mathbf{x}_{i}^{\prime}(t)\}$ $\triangleright$ Use Eq. (C.4)
7:Initialize zero density field $\rho((x_{i},y_{j}),t)$ across the extended domain $\Omega^{\text{ext}}$
8:for each pedestrian $p=1,2,\ldots,N$ do $\triangleright$ Account for “mirrored” pedestrians
9:  Compute kernel contribution $K_{h}(\mathbf{x}-\mathbf{x}_{p}(t))$ over the extended support $\Omega^{\text{ext}}$ $\triangleright$ Use Eq. (C.1)
10:  Add contribution to density field $\rho((x_{i},y_{j}),t)$
11:end for
12:Normalize $\rho((x_{i},y_{j}),t)$ with total number of pedestrians, NN and “mirrored” ones
13:Retrieve density $\rho((x_{i},y_{j}),t)$ in computational domain $\Omega$
14:Employ binary mask function to ensure zero density in obstacle subdomain $\Gamma$, by setting $\triangleright$ See Eq. (C.5)
[…]
16:Compute total estimated density over the entire domain $S(t)=\sum_{i,j=1}^{n_{x},n_{y}}\rho((x_{i},y_{j}),t)\delta x\delta y$
17:Normalize density field $\rho((x_{i},y_{j}),t)\leftarrow\rho((x_{i},y_{j}),t)/S(t)$
18:return Normalized density $\rho((x_{i},y_{j}),t)$ across computational domain $\Omega$
In the presence of obstacles, formally described as a subdomain $\Gamma\subset\Omega$, the standard KDE must be adjusted to ensure that the density vanishes within and along the boundaries of these regions. This is typically achieved by introducing a binary mask function $\mathcal{M}(x,y)$:
$M(x,y)=\begin{cases}0,&\text{if }(x,y)\in\Gamma\subset\Omega,\\ 1,&\text{otherwise}.\end{cases}$ (C.5)
which is then applied to the initially estimated densities to enforce zero density in $\Gamma$, implying:
$\rho((x_{i},y_{j}),t)=\mathcal{M}(x_{i},y_{j})\,\rho((x_{i},y_{j}),t),\quad\forall(x_{i},y_{j})\in\Omega.$ (C.6)
Appendix D Lag window size selection via the Bayesian and the Akaike Information Criteria
As discussed in Section 3.5, to determine the optimal lag window size $w$ for the MVAR model, we used two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [88, 89, 90] . These criteria provide a balance between model fit and complexity, and are used to avoid underfitting or overfitting the temporal dynamics. Both tests work under the assumption of data stationarity. The BIC is defined as:
$\text{BIC}(w)=-2\ln(\mathcal{L}(w))+k\ln(n_{t}),$ (D.1)
where $\mathcal{L}$ is the maximized likelihood of the model. For an MVAR model with $d$ latent dimensions and $w$ order of the model, the total number of parameters is then $k=d^{2}\cdot w$. The AIC, which provides a different trade-off between model complexity and fit, is given by:
$\text{AIC}(w)=-2\ln(\mathcal{L}(w))+2k.$ (D.2)
While AIC tends to favor more complex models and is better suited for short-term predictive accuracy in smaller datasets, BIC imposes a stronger penalty on complexity, making it preferable for larger sample sizes. The log-likelihood $\ln(\mathcal{L}(w))$ is computed under the assumption of Gaussian noise, and is derived from the residual covariance matrix of the MVAR model.
Finally, to select the optimal lag window $w^{*}$, we evaluated the BIC and AIC for all candidate lags $w\in\{1,2,\dots,w_{\text{max}}\}$. The chosen lag $w^{*}$ corresponds to the value of $w$ that yielded the smallest BIC or AIC in this range, i.e.,:
$w^{*}_{\text{BIC}}=\underset{w\in\{1,\dots,w_{\text{max}}\}}{\arg\min}\,\text{BIC}(w),\quad w^{*}_{\text{AIC}}=\underset{w\in\{1,\dots,w_{\text{max}}\}}{\arg\min}\,\text{AIC}(w).$ (D.3)
Appendix E Comparison of the “ground truth” density fields with the predicted ones, using MVAR(9), LSTM(4) and LSTM(9) models
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
Figure E.1: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the MVAR(9) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the LSTM(4) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq. (35) using the LSTM(9) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.

### v2

@@
 Abstract
-Bridging the microscopic/individual and the macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control tasks.
+Bridging the microscopic/individual and macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control.
 Here, we propose a combined manifold and machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations.
@@
 The proposed approach is a four-stage one, explicitly conserving the mass of the reconstructed dynamics in the high-dimensional space.
-In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation (KDE).
-In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distribution.
-The third step involves learning reduced-order surrogate models (ROMs) in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs).
+In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation.
+In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distributions.
+The third step involves learning reduced-order surrogate models in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs).
 Finally, we reconstruct the crowd dynamics in the high-dimensional space in terms of macroscopic density profiles.
-We demonstrate that the POD reconstruction of the density distribution via SVD conserves the mass.
-With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation (PDE) for the evolution of the density distribution.
-For our illustrations, we use the Social Force Model (SFM) to generate data in a corridor with an obstacle, imposing periodic boundary conditions.
+We demonstrate that the POD reconstruction of the density distribution via Singular Value Decomposition conserves the mass.
+With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation for the evolution of the density distribution.
+For our illustrations, we use the Social Force Model to generate data in a corridor with an obstacle, imposing periodic boundary conditions.
 The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations.
-Notably, linear MVAR models surpass their nonlinear counterparts, such as LSTMs, in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
-Keywords Crowd Dynamics, Multiscale modelling, Micro to Macro, Machine Learning, Numerical Analysis, Conservation laws, Reduced Order Models, Manifold Learning
+Notably, linear MVAR models surpass nonlinear LSTMs in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
+Keywords Crowd Dynamics, Multiscale modelling, Machine Learning, Numerical Analysis, Conservation laws, Reduced Order Models
 1 Introduction
@@
 Moreover, in [18] the authors proposed nonlocal conservation laws for traffic and crowd dynamics, where the flux depends not only on local density but also on averaged information from the surrounding region.
-In [11] a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior.
-More recently, in [16] are highlighted new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
-A major objective, when high-fidelity microsocpic/individual/agent-based simulators and/or detailed experimental data are available, is to derive mesoscopic or macroscopic models that can be then used for rigorous numerical analysis, optimization and control of the emergent/collective dynamics, for instance, in the optimal design of spaces, the reduction of evacuation times, and the management of crowd flows [11, 19, 20].
+In [11], a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior.
+More recently, [16] highlights new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
+A major objective, when high-fidelity microsocpic/individual/agent-based simulators and/or detailed experimental data are available, is to derive mesoscopic or macroscopic models that can then be used for numerical analysis, optimization and control of the emergent/collective dynamics, for instance, in the optimal design of spaces, the reduction of evacuation times, and the control of the flows [11, 9, 19, 20, 21, 22, 23, 24].
 In this context, kinetic models serve as an intermediate bridge.
@@
 These methods establish relationships between higher-order moments of the microscopic distributions and a limited, smaller set of lower-order moments that capture the observable emergent behavior [25, 26, 27].
-However, such closures inherently rely on assumptions that introduce biases both for modelling and analyses tasks.
-Typically, these assumptions include infinitely large populations, homogeneous agents, and uniform and/or local interaction networks, but also unknown “behavioural” characteristics, which can significantly impact the accuracy and validity of collective-level analyses.
-Recently, data-driven computational approaches based on physics-free machine learning (ML) have been employed for the modelling, numerical analysis, and control of pedestrian, crowd, and traffic dynamics.
-In this context, a wide range of ML architectures has been explored, including deep neural networks (DNNs) [24], Long Short-Term Memory (LSTM) networks [25, 26] for learning local movement decisions, Generative Adversarial Networks (GANs) [27, 28], Convolutional Neural Networks (CNNs) [29, 30, 31], and reinforcement learning methods [32, 33].
+However, such closures inherently rely on assumptions that introduce biases both for modelling and analysis tasks.
+Typically, these assumptions include infinitely large populations, homogeneous agents, and uniform and/or local interaction networks, but also unknown “behavioral” characteristics, which can significantly impact the accuracy and validity of collective-level analyses.
+Recently, data-driven computational approaches based on physics-free machine learning (ML) have been employed for modelling, numerical analysis, and control of pedestrian, crowd, and traffic dynamics.
+In this context, a wide range of ML architectures has been explored, including deep neural networks (DNNs) [28], Long Short-Term Memory (LSTM) networks [29, 30], Generative Adversarial Networks (GANs) [31, 32], Convolutional Neural Networks (CNNs) [33, 34, 35], and reinforcement learning methods [36, 37].
 For a comprehensive review of these and related ML approaches, see [38].
 These approaches have proven effective in forecasting individual and group trajectories within the framework of time-series modelling and forecasting, without relying on physics-based formulations.
-However, at the same time, such approaches lack interpretability and, importantly, do not incorporate information of the global “spatio-temporal” emergent dynamics, such as density fields, that are essential in modelling crowd dynamics.
+However, at the same time, such approaches lack interpretability and, importantly, do not incorporate information of the global spatio-temporal emergent dynamics, such as density fields, that are essential in modelling crowd dynamics.
 An alternative data-driven paradigm that bypasses the need to explicitly learn surrogate models is the Equation-free multiscale framework [39].
-By appropriately coupling on-demand microscopic simulations with coarse-graining (and fine-graining), this approach builds a coarse time-stepper for the evolution at the macroscopic scale, thus bypassing the need for analytically derived continuum equations [36, 37, 38].
-While Equation-free framework facilitates powerful multiscale numerical analysis and control, it lacks an explicit, global macroscopic equation for the crowd/collective dynamics–that can be approximated, such as a mass-conserving PDE–limits interpretability, physical insight, and generalization for long-term simulations.
-To address the aforementioned challenges, a few recent studies have utilized physics-informed neural networks (PINNs) [39] including physics-informed neural operators (NOs) [40, 41] for solving the inverse problem, i.e.
-that of constructing from data, hybrid models in the form of “grey-box” PDE models [42, 43, 21, 23, 41].
+By appropriately coupling on-demand microscopic simulations with coarse-graining, or “restriction”, and subsequent fine-graining, or “lifting”, this approach builds a coarse time-stepper for the evolution at the macroscopic scale, thus bypassing the need for analytically derived continuum equations [40, 41, 42, 43, 44].
+Although the equation-free framework enables powerful multiscale numerical analysis and control, it is fundamentally based on the on-demand construction of local discrete maps rather than on deriving a global macroscopic model of the crowd or collective dynamics (for example, a mass-conserving partial differential equation (PDE)).
+That reliance on local maps reduces interpretability and physical insight, and it hinders reliable generalization in long-term simulations.
+To address the aforementioned challenges, a few recent studies have utilized physics-informed neural networks (PINNs) [45] including physics-informed neural operators (NOs) [46, 47] for solving the inverse problem, i.e., that of constructing hybrid models in the form of “grey-box” PDE models from data [48, 49, 25, 27, 47].
 Such an approach, combines ML with underlying knowledge of physical principles, such as conservation laws [50] and closures [51], to create hybrid models that can both predict the spatio-temporal dynamics and ensure physical consistency, even under data scarcity.
 In the context of crowd dynamics, PINNs have been employed to construct both gray-box hydrodynamic PDEs [52] as well as for discovering the physics/rules governing interactions and motion at the microscopic level [53].
-However, while PINNs provide an effective way to incorporate physical constraints and knowledge into machine learning models, they can also introduce biases due to their reliance on predefined forms of PDEs.
-Moreover, a key physical property in crowd modeling—mass conservation—is typically not enforced explicitly, but rather imposed as a soft constraint through the loss function [48, 44, 49].
+However, while PINNs provide an effective way to incorporate physical constraints and knowledge into ML models, they can also introduce biases due to their reliance on predefined forms of PDEs.
+Moreover, a key physical property in crowd modeling, mass conservation, is typically not enforced explicitly, but rather imposed as a soft constraint through the loss function [54, 50, 55].
 This indirect enforcement can limit the generalizability and physical reliability of PINN-based models, particularly in heterogeneous or unforeseen scenarios where the true dynamics deviate from the assumed PDE forms.
-For learning classes of PDEs that incorporate conservation laws—such as those found in crowd dynamics—the use of black-box deep neural networks (DNNs) and black-box neural operators (NOs) [50, 51, 49, 52, 53, 23] is inherently problematic.
+For learning classes of PDEs that incorporate conservation laws–such as those found in crowd dynamics–the use of black-box DNNs and black-box NOs [56, 57, 55, 58, 59, 27] is inherently problematic.
 These models often lack physical consistency, particularly in preserving conservation laws, which are usually enforced only as soft constraints in the loss function.
-In addition, their training suffers from low approximation accuracy due to the curse of dimensionality, potentially leading to unphysical predictions and numerical instabilities during long-term simulations.
+In addition, their training suffers from low approximation accuracy due to the “curse of dimensionality”, potentially leading to unphysical predictions and numerical instabilities during long-term simulations.
 Thus, to deal with that problem, RiemannONets were recently presented in [60], which are NOs based on DeepONets to solve Riemann problems encountered in compressible flows, thus enforcing the preservation of conservation laws.
-Towards solving the inverse problem, we have recently proposed GoRINNs [55]: Godunov-Riemann informed neural networks for learning hyperbolic PDEs that respect the conservation laws explicitly.
+Towards solving the inverse problem, we have recently proposed GoRINNs [61], neural networks coupled with Godunov numerical schemes and Riemann solvers for learning hyperbolic PDEs that explicitly respect the conservation laws by construction.
 It is important to note, however, that despite the long-standing tradition of modeling crowd dynamics through (hydrodynamic) PDEs, such formulations inherently rely on an “infinite-size” assumption.
-In realistic settings involving crowds of moderate or finite size, this assumption often breaks down.
-Finite-size effects—such as spatial heterogeneity, individual interactions, and boundary influences—can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior.
-Here, inspired by the Equation-free multiscale framework [35], and building on our previous recent works on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [56, 36, 57, 23, 58, 41], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics.
+In realistic settings involving crowds of moderate or finite-size, this assumption often breaks down.
+Finite-size effects, such as spatial heterogeneity, individual interactions, and boundary influences, can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior.
+Here, inspired by the Equation-free multiscale framework [39], and building on our previous recent works, on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [62, 42, 63, 27, 64, 47], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics.
 The proposed framework bridges the microscopic and macroscopic modelling scales, while explicitly preserving the underlying conservation law of mass in the ambient space.
@@
 Second, we project the macroscopic fields into a low-dimensional latent space using Proper Orthogonal Decomposition (POD).
-Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM networks.
+Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM models.
 These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’ theorem [65] for reconstructing the phase-space structure (here in the latent space).
-The concept of the approach (i.e.
-that using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [60, 61] (see also results and discussion in [61, 56, 62, 63, 57]).
-In crowd dynamics, behavioural characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Taken’s theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics.
+The concept of the approach (i.e., using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [66, 67] (see also results and discussion in [67, 62, 68, 69, 63]).
+In crowd dynamics, behavioral characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Takens’ theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics.
 Finally, we lift the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis.
@@
 We utilize a detailed simulator that models the dynamics of a human crowd, consisting of $N$ pedestrians moving within a domain $\Omega\subset\mathbb{R}^{2}$.
-The overall state of the pedestrian $i\in\{1,\ldots,N\}$ is characterized by its position $\mathbf{x}_{i}=\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ and velocity $\mathbf{v}_{i}=\mathbf{v}_{i}(t)=(v_{ix}(t),v_{iy}(t))$, and also a set of “behavioral” variables $\bm{u}_{i}=\bm{u}_{i}(t)\in\mathbb{R}^{b}$, which may include phenomena such as decision-making strategies, herding behavior, but also panic and irrationality [65, 66, 67, 68, 69], where $b\in\mathbb{R}$ is the number of behavioral traits modeled for each pedestrian.
+The state of each pedestrian $i\in\{1,\ldots,N\}$ is characterized by its position $\mathbf{x}_{i}=\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ and velocity $\mathbf{v}_{i}=\mathbf{v}_{i}(t)=(v_{ix}(t),v_{iy}(t))$, and also a set of “behavioral” variables $\bm{u}_{i}=\bm{u}_{i}(t)\in\mathbb{R}^{b}$, which may include phenomena such as decision-making strategies, herding behavior, but also panic and irrationality [71, 72, 73, 20, 74], where $b\in\mathbb{N}$ is the number of behavioral traits modeled for each pedestrian.
 Additionally, we introduce an auxiliary variable $\bm{z}_{i}\in\mathbb{R}^{b}$ to denote the rate of change of the behavioral variables $\bm{u}_{i}$.
@@
 The majority of models in this category are based on the celebrated social force model (SFM) for pedestrian dynamics [1, 16], which was proposed in the mid-1990s.
-In contrast, at the macroscopic scale, crowd dynamics are typically described using phenomenological continuum models, most often formulated as partial differential equations (PDEs).
+In contrast, at the macroscopic scale, crowd dynamics are typically described using phenomenological continuum models, most often formulated as PDEs.
 These models capture the collective behavior of large-scale pedestrian flows and are grounded in the theories of hydrodynamics and continuum mechanics.
@@
 These methods are predominantly based on DNNs/NOs, thus suffering from the “curse of dimensionality” that is inherent in their training.
-To circumvent this challenge, learning Reduced Order Models (ROMs) in well-constructed latent spaces is crucial, since it significantly lowers computational cost while capturing the essential dynamics.
+To circumvent this challenge, learning ROMs in well-constructed latent spaces is crucial, since it significantly lowers computational cost while capturing the essential dynamics.
 These ROMs must also inherently preserve fundamental principles like mass conservation, ensuring physical accuracy in the reconstructed space.
 Towards this goal, we aim at bridging high-fidelity microscopic/agent-based simulations and/or experimental/real-world data with macroscopic scale modelling, using ROMs that preserve mass conservation, within the framework of learning discrete operators/coarse-timesteppers [39].
-The proposed methodology, presented in detail in the following section, is based on our previous efforts to develop what we call next generation Equation-free framework [56, 36, 57, 23, 58].
+The proposed methodology, presented in detail in the following section, is based on our previous efforts to develop what we call the next generation Equation-free (EF) framework [62, 42, 63, 27, 64].
 This framework deals with the “curse of dimensionality” by learning surrogate ML models in appropriate latent spaces emerging from the complex spatio-temporal dynamics of high-dimensional systems.
-It originates from the celebrated Equation-free approach [35], which enables high-dimensional simulations to perform system-level numerical analysis tasks on an appropriate latent space.
+It originates from the Equation-free approach [39], which enables high-dimensional simulations to perform system-level numerical analysis tasks on an appropriate latent space.
 This is achieved via the construction of a lifting operator $\mathcal{L}$ and a restriction operator $\mathcal{R}$ bridging the gap between states in a high-dimensional space, say $\mathbf{u}(t)\in\mathbb{R}^{M}$, and states in a latent space, say $\mathbf{y}_{d}(t)\in\mathbb{R}^{d}$, with $d\ll M$, as:
@@
 A key distinction between our proposed next generation EF and the traditional EF framework is that, in the latter, one constructs local (numerical) maps and computes the quantities needed for bifurcation analysis on demand via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured usually by a few moment distributions).
-In contrast, our proposed method learns a “global dynamical model“ in Eq.
-(5) via machine learning from long-term, high-fidelity simulations spanning the entire phase space, in low-dimensional latent spaces.
+In contrast, our proposed method learns a “global dynamical model” in Eq.
+(5) via ML from long-term, high-fidelity simulations spanning the entire phase space, in low-dimensional latent spaces.
 This global dynamical-systems-based perspective enables the use of both linear (such as POD) and nonlinear manifold-learning techniques (such as Diffusion maps [76, 77]), and techniques for the solution of the pre-image problem, such as geometric harmonics [78, 79] or autoencoders [80, 81], to construct maps between the high-dimensional and the latent space.
-The methodology is described analytically in the following section.
-The key advantage of our next generation EF framework is its ability to circumvent the need for explicit derivation or closure assumptions in PDEs—a process that inherently introduces biases in modelling crowd density dynamic.
-Instead, it directly learns the “actual” solution operator in Eq.
+The methodology is described in detail in the following section.
+The key advantage of our next generation EF framework is its ability to circumvent the need for explicit derivation or closure assumptions in PDEs—a process that inherently introduces biases in modelling crowd density dynamics.
+Instead, it directly learns the “actual” evolution operator in Eq.
 (5)–-the map that advances observed macroscopic distributions (here the density) forward in time.
-This learned operator, embedded within computationally efficient linear or nonlinear autoregressive ROMs (MVAR or LSTM, respectively), effectively encodes the complex, often analytically intractable, dynamics of the crowd.
-By focusing on learning how densities evolve (the solution operator) rather than learning governing equations in the form of PDEs, the method circumvents the significant challenges of the “curse of dimensionality” associated with DNNs/NOs training of such PDEs.
+This learned operator, embedded within computationally efficient linear or nonlinear autoregressive ROMs (MVARs or LSTMs, respectively), effectively encodes the complex, often analytically intractable, dynamics of the crowd.
+By focusing on learning how densities evolve (the evolution operator) rather than learning governing equations in the form of PDEs, the method circumvents the significant challenges of the “curse of dimensionality” associated with DNNs/NOs training of such PDEs.
 The resulting ROMs thus act as practical, data-driven surrogates that implicitly contain the action of the unknown PDE’s solution operator, enabling accurate prediction of future crowd states solely from observed density snapshots.
-Importantly, we demonstrate that for the proposed configuration, the POD method (used for identifying the latent space) conserves explicitly the mass of the reconstructed density field resulting from the ROMs.
+Importantly, we demonstrate that for the proposed configuration, the POD approach (used for identifying the latent space) conserves explicitly the mass of the reconstructed density field resulting from the ROMs.
 An explicit mass conservation of the continuum representations is an essential physical constraint for any credible surrogate model of crowd dynamics.
@@
 The proposed methodology consists of four main steps.
-First, we map the discrete distributions of pedestrian positions into continuous density fields defined on a spatial grid using Kernel Density Estimation (KDE) [77].
+First, we map the discrete distributions of pedestrian positions into continuous density fields defined on a spatial grid using KDE [82].
 Second, we construct a restriction operator to project the resulting density fields into an appropriate latent space.
 For this purpose, we employ the POD [83] method on normalized density profiles across the computational domain, ensuring that mass conservation is preserved in the reconstructed fields.
-Third, within this latent space, we train surrogate autoregressive ROMs using MVAR and LSTM time-series forecasting models to capture the complex dynamics present in the sequential data.
-Finally, in the fourth step, we lift the latent ROM dynamics back to the macroscopic density ambient space by solving the preimage problem.
+Third, within the POD latent space, we train surrogate autoregressive ROMs using MVAR and LSTM time-series forecasting models to capture the complex dynamics present in the data.
+Finally, in the fourth step, we lift the latent ROM dynamics back to the macroscopic density ambient space by solving the pre-image problem.
 In the case of POD, this lifting step is computationally inexpensive, as it reduces to a simple linear projection onto the ambient space using the stored orthogonal basis.
@@
 First, microscopic distributions of pedestrian positions are mapped to macroscopic density fields (Step 1) and embedded into a latent space using POD (Step 2).
-Second, surrogate reduced-order models (MVAR and LSTM) are trained in this latent space to capture and forecast the complex dynamics (Step 3).
+Second, surrogate reduced-order models (MVARs and LSTMs) are trained in this latent space to capture and forecast the complex dynamics (Step 3).
 Finally, the learned dynamics are lifted back to the high-dimensional density fields via linear projection with the constructed POD basis (Step 4).
@@
 (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves.
-Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator in Appendix A.
-We next employ Multivariate Kernel Density Estimation (KDE) [79] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
+Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator, detailed in Appendix A.
+We next employ KDE [84] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
 $\rho(\mathbf{x},t)=\frac{1}{N}\sum_{i=1}^{N}K_{h}\left(\mathbf{x}-\mathbf{x}_{i}(t)\right),$ (6)
@@
 (6).
-Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid-points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
+Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
 2.2.2 The Restriction and Lifting Operators
-Having obtained the macroscopic density fields, we now proceed to the second and fourth key elements of the approach, which, in analogy to the Equation-free framework, constitute the construction of an “embedding/restriction” and “lifting” operators to and from an appropriate low-dimensional latent space.
-The restriction operator, say, $\mathcal{R}$, [35], is a map from density distributions $\rho(\mathbf{x},t)$ to a low-dimensional latent space coordinates, say $Y_{d}(t)$, i.e.,:
+Having obtained the macroscopic density fields, we now proceed to the second and fourth key elements of the approach, which, in analogy to the EF framework, constitute the construction of an “embedding/restriction” and “lifting” operators to and from an appropriate low-dimensional latent space.
+The restriction operator, say, $\mathcal{R}$, [39], is a map from density distributions $\rho(\mathbf{x},t)$ to low-dimensional latent space coordinates, say $Y_{d}(t)$, i.e.,:
 $\mathcal{R}:\rho(\mathbf{x},t)\in\mathbb{R}^{n_{c}}\mapsto Y_{d}(t)\in\mathbb{R}^{d},\quad d\ll n_{c},$ (7)
 where $n_{c}=n_{x}\times n_{y}$.
-To construct the restriction operator, we apply Proper Orthogonal Decomposition (POD) [78] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq.
+To construct the restriction operator, we apply POD [83] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq.
 (6), for different initial conditions; see Appendix B for details.
@@
 $X=[x_{1},x_{2},\dots,x_{n_{t}}]\,\in\mathbb{R}^{n_{c}\times n_{t}},$ (9)
-with columns, say, $x_{k}\in\mathbb{R}^{n_{c}}$ being the normalized density field along cases and timesteps $k=1,\ldots,n_{t}=C\cdot K$.
+with columns, say $x_{k}=\rho^{(c)}(\mathbf{x},t_{k})/S(t_{k})\in\mathbb{R}^{n_{c}}$, being the normalized density fields along cases and time steps $k=1,\ldots,n_{t}=C\cdot K$.
 To complete the restriction operator at the second step, we employ POD on $X$, implemented via Singular Value Decomposition (SVD) on the centered matrix:
-$\bar{X}=XH=U\Sigma V^{T},\quad H=I_{n_{t}}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top},$ (10)
+$\bar{X}=XH=U\Sigma V^{\top},\quad H=I_{n_{t}}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top},$ (10)
 where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular vectors, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim.
@@
 The latent space is spanned by the first $d$ left-singular vectors forming $U_{d}$.
-Thus, projecting $X$ on the latent space we get, the low-dimensional embedding:
-$Y_{d}=U_{d}^{T}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
-For the lifting operator that maps $Y_{d}$ to the space of density profiles, we demonstrate that the POD reconstruction (aka the POD lifting operator) preserves the original density.
+Thus, the projection of $X$ on the latent space provides the low-dimensional embedding:
+$Y_{d}=U_{d}^{\top}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
+For the lifting operator that maps $Y_{d}$ to the space of density profiles, we demonstrate that the POD reconstruction (aka the POD lifting operator) preserves the original total density.
 This aligns with the general characteristic of POD in maintaining symmetries [85].
@@
 Proposition 1.
-Let us assume a matrix $X\in\mathbb{R}^{n_{c}\times n_{t}}$, with columns being normalized density fields. If the density is preserved along timesteps, i.e., if
+Let us assume a matrix $X\in\mathbb{R}^{n_{c}\times n_{t}}$, with columns being normalized density fields. If the density is preserved along time steps, i.e., if:
 $\begin{bmatrix}\sum_{i=1}^{n_{c}}x_{i,1}&\sum_{i=1}^{n_{c}}x_{i,2}&\dots&\sum_{i=1}^{n_{c}}x_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}},$ (12)
-then the reconstructed density field computed by the POD as:
+then, the reconstructed density field, computed by the POD as:
 $\tilde{X}=U_{d}U_{d}^{\top}\bar{X}+X(I_{n_{t}}-H),\,\,$ (13)
-where $U_{d}=[u_{1},u_{2},\dots u_{d}]\in\mathbb{R}^{n_{c}\times d}$ is the orthonormal basis formed by the first $d$ left-singular vectors computed by the SVD, is also preserved, i.e.,:
+where $U_{d}=[u_{1},u_{2},\dots u_{d}]\in\mathbb{R}^{n_{c}\times d}$, is the orthonormal basis formed by the first $d$ left-singular vectors computed by the SVD, is also preserved, i.e.,:
 $\begin{bmatrix}\sum_{i=1}^{n_{c}}\tilde{x}_{i,1}&\sum_{i=1}^{n_{c}}\tilde{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\tilde{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (14)
 Proof.
-As by hypothesis, the sum of each column of $X$ is equal to one (that is $\mathbf{1}_{n_{c}}^{\top}X=\mathbf{1}_{n_{t}}^{\top}$), we have that:
+By construction, since the sum of each column of $X$ is equal to one (that is $\mathbf{1}_{n_{c}}^{\top}X=\mathbf{1}_{n_{t}}^{\top}$), we have that:
 $\mathbf{1}_{n_{c}}^{\top}\bar{X}=\mathbf{1}_{n_{c}}^{\top}XH=\mathbf{1}_{n_{t}}^{\top}H=\mathbf{1}_{n_{t}}^{\top}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}^{\top}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top}=\mathbf{1}_{n_{t}}^{\top}-\mathbf{1}_{n_{t}}^{\top}=\boldsymbol{0}^{\top}_{n_{t}}.$ (15)
 Therefore, the sum of each column of the centered matrix $\bar{X}$ is equal to zero.
-This implies that $\mathbf{1}_{n_{c}}$ is orthogonal to the column space of $\bar{X}$, and therefore is orthogonal to the left singular vectors of the SVD decomposition of $\bar{X}$, which provide an orthogonal base for the column space, i.e.,:
+This implies that $\mathbf{1}_{n_{c}}$ is orthogonal to the column space of $\bar{X}$, and therefore is orthogonal to the left singular vectors of $\bar{X}$, which provide an orthogonal basis for the column space, i.e.,:
 $\mathbf{1}_{n_{c}}^{\top}U_{d}=\boldsymbol{0}^{\top}_{d},$ (16)
-where $d=1,2,\dots r$, with $r=rank(X)$. Then, the multiplication of the reconstructed matrix in Eq. (13) by $\mathbf{1}_{n_{c}}^{\top}$ implies:
+where $d=1,2,\dots r$ with $r=rank(X)$. Then, the multiplication of the reconstructed matrix in Eq. (13) by $\mathbf{1}_{n_{c}}^{\top}$ implies:
 $\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\mathbf{1}_{n_{c}}^{\top}U_{d}U_{d}^{\top}\bar{X}+\mathbf{1}_{n_{c}}^{\top}X(I_{n_{t}}-H)=\mathbf{0}_{n_{t}}^{\top}+\frac{1}{n_{t}}\mathbf{1}_{n_{t}}^{\top}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top}=\mathbf{1}_{n_{t}}^{\top},$ (17)
 thus getting:
-$\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\begin{bmatrix}\sum_{i=1}^{n_{c}}\hat{x}_{i,1}&\sum_{i=1}^{n_{c}}\hat{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\hat{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (18)
-Hence, the POD reconstruction preserves the original density. ∎
+$\mathbf{1}_{n_{c}}^{\top}\tilde{X}=\begin{bmatrix}\sum_{i=1}^{n_{c}}\tilde{x}_{i,1}&\sum_{i=1}^{n_{c}}\tilde{x}_{i,2}&\dots&\sum_{i=1}^{n_{c}}\tilde{x}_{i,n_{t}}\end{bmatrix}=\mathbf{1}^{\top}_{n_{t}}.$ (18)
+Hence, the POD reconstruction preserves the total original density. ∎
 In summary, given a density distribution $\rho(\mathbf{x},t)$ obtained via Eq. (6), we obtain the normalized density along the $n_{c}$ grid points of $\Omega$ as:
-$x(t)=\dfrac{1}{\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}}$ (19)
-and define the restriction operator $\mathcal{R}$ by the projection on the POD basis of Eq. (11) as:
+$x(t)=\dfrac{1}{S(t)}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}},$ (19)
+where $S(t)$ is the total mass on time $t$, and define the restriction operator $\mathcal{R}$ by the projection on the POD basis of Eq. (11) as:
 $\boldsymbol{y}(t)=\mathcal{R}(x(t))=U_{d}^{\top}\left(x(t)-\frac{1}{n_{t}}X\mathbf{1}_{n_{t}}\right)\in\mathbb{R}^{d},$ (20)
@@
 2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
-With the restriction operator constructed, which maps density fields to latent coordinates, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data.
+With the restriction operator, which maps density fields to latent coordinates, constructed, the third step of the proposed approach involves training ROMs to learn the latent dynamics from sequential data.
 We note that while the restriction and lifting operators $\mathcal{R}$ and $\mathcal{L}$ are, in principle, defined over continuous time inputs, we apply them at a discrete set of uniformly spaced time instances $t_{k}=t_{0}+k\delta t$ for $k=1,2,\ldots,K$, consistent with the temporal sampling of our data; see Eq.
@@
 The predictor vectors $\boldsymbol{y}_{c}(t_{k-j})\in\mathbb{R}^{d}$ include the $w$ past lag latent variables evaluated at discrete times $t_{k-1},\ldots,t_{k-w}$, while the error term $\boldsymbol{\varepsilon}(t_{k})\in\mathbb{R}^{d}$ is assumed to follow a multivariate white noise process:
-$\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})]=\mathbf{0},\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{k})^{\top}]=\Sigma,\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{r})^{\top}]=\mathbf{0},\quad\text{for }t_{k}\neq t_{r},$ (23)
+$\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})]=\mathbf{0},\qquad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{k})^{\top}]=\Sigma\quad\text{and}\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{r})^{\top}]=\mathbf{0}\quad\forall k\neq r,$ (23)
 for every $k,r=1,\ldots,K$.
@@
 The optimization problem in Eq.
-(24) can be solved using methods such as regularized least square methods; in particular, we used ridge regularization [81].
+(24) can be solved using methods such as regularized least squares; in particular, we used ridge regularization [86].
 We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations.
@@
 $\hat{\boldsymbol{y}}_{c}(t_{k})$ $=\mathbf{W}_{y}\,\mathbf{h}_{t_{k-1}}+\mathbf{b}_{y},$ (28)
-where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the, evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$.
+where $\mathbf{h}_{t_{k}}\in\mathbb{R}^{N_{h}}$ and $\mathbf{c}_{t_{k}}\in\mathbb{R}^{N_{h}}$ are the evaluated at discrete times $t_{k}$, hidden and memory cell states, respectively, $N_{h}$ corresponds to the number of hidden units and $\mathbf{W}_{y}\in\mathbb{R}^{d\times N_{h}},\mathbf{b}_{y}\in\mathbb{R}^{d}$ define an output layer mapping the last hidden state $\mathbf{h}_{t_{k-1}}$ to the predicted latent variable vector $\hat{\boldsymbol{y}}_{c}(t_{k})\in\mathbb{R}^{d}$.
 The gating mechanism of the LSTM is included in Eq.
-(27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for the conciseness of the presentation; a vanilla LSTM including 4 gates, contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
-Similarly to the sequence-to-one learning of MVARs, given sequential data for $\boldsymbol{y}_{c}(t_{k})$ for a fixed lag window of width $w$, we train the LSTM model by solving the optimization problem:
+(27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for conciseness of the presentation; a vanilla LSTM including 4 gates, and contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
+Similarly to the sequence-to-one learning of MVARs, given sequential data for $\{\boldsymbol{y}_{c}(t_{k})\}$ for a fixed lag window of width $w$, we train the LSTM model by solving the optimization problem:
 $\operatorname*{argmin}_{\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}}\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right),$ (29)
@@
 2.
-Figure 2: The case study: The SFM simulating pedestrians moving in a corridor with an obstacle.
+Figure 2: The case study: The SFM simulating pedestrians moving in a corridor from the left to the right with an obstacle.
 The square obstacle of area $12.96\,\text{m}^{2}$ is centered at $(25.8\,\text{m},1.8\,\text{m})$, with the pedestrians navigating around it towards reaching the corridor’s right end.
@@
 In the SFM, these terms translate into a goal-directed force (playing the role of friction), social and repulsive (from obstacles) deterministic forces, and a random term, reading:
-$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{i,j}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{i,B}(\bm{x}_{i})}_{\text{repulsive force}}+\mathbf{F}_{rand},$ (31)
+$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{ij}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{iB}(\bm{x}_{i})}_{\text{repulsive force}}+\mathbf{F}_{rand},$ (31)
 where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time.
@@
 Upon reaching this intermediate waypoint, their final destination is updated to $(x,y)=(48\,\text{m},\,4.2\,\text{m})$, which lies at the right end of the corridor and at 35% of its height.
-To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents that reach the right boundary at $x=48\,\text{m}$, at the left boundary at $x=0\,\text{m}$.
+To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents at the left boundary at $x=0\,\text{m}$, when agents reach the right boundary at $x=48\,\text{m}$.
 Pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator.
-Such configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.
+Such a configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.
 3.2 Data generation of microscopic distributions
 We generate synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy.
-Each simulation has a time span of $275\,\text{s}$.
+Each simulation has a time span of $250\,\text{s}$.
 To reduce computational storage costs while retaining essential temporal features, we subsample the resulting pedestrian trajectories at intervals of $\delta t=0.25\,\text{s}$.
 As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios.
-In particular, we consider $C=10$ different initializations with zero velocities and position distributions including: (i) uniform distribution to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) piecewise linear distributions with higher densities near the wall boundaries to simulate characteristic wall-avoidance behavior, (iv) sinusoidal distributions to generate periodic macroscopic behavior, and (v) double Gaussian distributions to capture the formation of two distinct crowds.
+In particular, we consider $C=20$ different initializations with zero velocities and positions distributed via: (i) uniform distributions to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) “double bell-shaped” Gaussian distributions to capture the formation of two distinct crowd groups and (iv) sinusoidal distributions to generate periodic macroscopic behavior.
 Details on the form and parameters of the above initializations are provided in Appendix B, distinguishing between those used for training and testing the proposed framework in Tables B.1 and B.2, respectively.
-The resulting datasets are comprised of 10 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1100}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1100$ timesteps are sampled from each trajectory.
-Thus, the total datasets overall cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(10\cdot 1100)}$ used for training and testing, respectively.
+The resulting datasets are comprised of 20 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1000}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1000$ time steps correspond to the samples recorded from each trajectory.
+Thus, the total datasets over all cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(20\cdot 1000)}$ used for training and testing, respectively.
 3.3 Extraction of continuous density fields
 For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1.
-To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behaviors, while minimizing artificial grid effects.
+To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behavior, while minimizing artificial grid effects.
 For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq.
@@
 All the details regarding KDE implementation are provided in Appendix C.
-Using the above procedure, we derive the density fields $\rho^{(c)}\big{(}x_{i},y_{j},t_{k}\big{)}$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing data sets $\mathcal{X}_{ts}$.
+Using the above procedure, we derive the density fields $\rho^{(c)}\big(x_{i},y_{j},t_{k}\big)$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing $\mathcal{X}_{ts}$ data sets.
 For conforming to the assumptions of Proposition 1, we finally normalize each density field following Eq.
 (19), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq.
-(8), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=10\cdot 1100$.
+(9), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=20\cdot 1000$.
 3.4 POD for the restriction and lifting operators
@@
 Next, using the first $d$ left singular vectors of $X_{tr}$, we form $U_{d}$ for completing the construction of the restriction and lifting operators in Eqs.
-(7, 21).
+(20, 21).
 To assess the accuracy of restriction and lifting operators, we evaluate the POD reconstruction error on the training and testing data sets by the relative $L_{2}$ error:
@@
 where $x(t_{k})\in X_{tr},X_{ts}$ with $k=1,\ldots,K$ is the normalized density defined in Eq. (19).
-3.5 Autoregressive ROMs and their training
-Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct autoregressive ROMs for learning the dynamics in the latent space.
-For this purpose, we learn linear MVARs and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$.
+3.5 ROMs for latent space dynamics and their training
+Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct ROMs for learning the dynamics in the latent space.
+For this purpose, we learn linear MVAR and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$.
 To conform to this sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq.
 (20) on the training data $X_{tr}$.
-As a preliminary step, we first tune the lag window ww–within a bound of $w_{\text{max}}=100$–using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [88, 89, 90], the details of which are described in Appendix D.
+As a preliminary step, we first tune the lag window $w$, within a bound of $w_{\text{max}}=100$, using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [93, 94, 95]; details are provided in Appendix D.
 Both criteria are likelihood-based, with AIC tending to select larger windows than BIC.
@@
 For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
-For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey-Fuller (ADF) test [91] with a threshold of significance level $p<0.01$.
-Then, we solve the least-squares optimization problem in Eq.
-(24) using ridge regression with a regularization parameter $\lambda=10^{-6}$ to compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq.
+For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey–Fuller (ADF) test [96] with a threshold of significance level $p<0.01$.
+Then, we compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq.
+(22) using ordinary least-squares (OLS) to solve the optimization problem in Eq.
 (24).
@@
 The model’s parameters are initialized with a Glorot uniform initialization [98], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints.
-The maximum number of epochs was set to $100$ as no further improvement was achieved using more epochs.
+The maximum number of epochs was set to $200$ as no further improvement was achieved using more epochs.
 To further avoid overfitting of the LSTM, we split the training data into training and validation sets using a 80-20% ratio, and shuffle each data set without breaking the internal temporal structure of the sequences.
@@
 (25) and (30), respectively, at the end of training.
-Since the MVARs loss function is based on $L_{2}$ error, we further report the MSE:
+To further quantify the uncertainty associated with the stochastic training of the LSTM models, we repeated their training procedure 50 times, each with different randomly initialized learnable parameters.
+We report the mean error of the final losses and the 10–90% error percentiles across the 50 runs.
+Since the MVARs loss function is based on $L_{2}$ error, we further report the MSE at the latent space:
 $MSE_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\rVert^{2},$ (34)
@@
 (26), to enable direct comparison with the MSE-based loss function of LSTMs.
-Additionally, we report the computational times required for training both ROMs; we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the ridge regression employed for solving the related optimization problem.
+Additionally, we report the computational times required for training both ROMs (for LSTMs, we report the mean error and 10–90% error percentiles statistics across the 50 runs); we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the OLS method employed for solving the related optimization problem.
 3.6 Accuracy assessment of the proposed framework
-Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $c=10$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details.
-In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$.
+Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $C=20$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details.
+In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$, using Eq.
+(19).
 Then, following the proposed restrict-evolve with ROM-lift framework in Eq.
-(5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities; i.e., computing:
+(5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities in both an open-loop and closed-loop setting.
+For the open-loop setting evaluation, we compute the prediction:
+$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(x^{(c)}(t_{k-1})),\ldots,\mathcal{R}(x^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (35)
+while for the closed-loop setting evaluation, we compute the prediction:
 $\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(\hat{x}^{(c)}(t_{k-1})),\ldots,\mathcal{R}(\hat{x}^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (36)
 where $\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators in Eqs.
-(7) and (21) and $\Phi_{\text{ROM}}$ is the autoregressive ROM with parameters $\mathbf{p}$ that can be either the trained MVARs or LSTMs model.
-Notice that while Eq.
-(35) could consider the restricted values of the normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, it instead considers the predicted normalized densities; thus, evaluating the proposed framework in a closed-loop (autoregressive) manner.
-This allows us to feed subsequent predictions recursively into Eq.
-(35) to advance the forecast over time.
-To quantify the prediction accuracy, we compute the relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ per case $c$ and time step $k$ as:
+(20) and (21) and $\Phi_{\text{ROM}}$ is the latent dynamics ROM with parameters $\mathbf{p}$ that can be either the trained MVAR or LSTM model.
+In particular, for the LSTM model, we consider the parameter set resulting in the best training performance among the 50 runs.
+Notice that the open-loop prediction in Eq.
+(35) considers the restricted values of the observed normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, while the closed-loop prediction in Eq.
+(36) instead considers the predicted–from previous steps–normalized densities $\mathcal{R}(\hat{x}^{(c)}(t_{k-j}))$ in a recursive fashion.
+Since the open-loop setting requires the “ground truth” densities, we use it here to evaluate the one-step predictions of our framework.
+For the desired, in practice, longer predictions, we utilize the closed-loop setting, which enables us to recursively feed subsequent predictions into Eq.
+(36), resulting in an autoregressive time-stepping framework also considering the effect of one-step forecast error accumulation.
+To quantify the end-to-end prediction accuracy, we compute the relative, reconstructed in the ambient-space, $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ (either for the open-loop or the closed-loop setting) per case $c$ in the testing set and time step $k$ as:
 $e^{1,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{1}}{\lVert x^{(c)}(t_{k})\rVert_{1}},\quad e^{2,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{2}}{\lVert x^{(c)}(t_{k})\rVert_{2}},\quad e^{\infty,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{\infty}}{\lVert x^{(c)}(t_{k})\rVert_{\infty}}.$ (37)
-We report these errors overall cases and time steps, also computing the 10-90 percentiles, and track down the evolution of these errors along time steps.
+For a summary of statistics, we report the mean and the 10–90% percentiles of the above errors over all cases and time steps.
+To track down the prediction accuracy in time, we further depict the mean and 10–90% percentiles of the error evolution in time over all cases.
 All numerical computations were executed on a PC equipped with an 11th Gen Intel(R) Core(TM) i7-1165G7 2.80 GHz CPU and 16 GB of system memory.
@@
 The resulting training dataset $\mathcal{X}_{tr}$ is used to train our framework, while the testing dataset $\mathcal{X}_{ts}$ is used to assess its efficiency and prediction accuracy.
-At the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3.
+In the first step, we extracted the normalized density fields data $X_{tr},X_{ts}$ from the discrete pedestrian positions in $\mathcal{X}_{tr},\mathcal{X}_{ts}$ using KDE, as described in Section 3.3.
 Next, we employed POD on $\mathcal{X}_{tr}$ to discover the latent space dimension; see Section 3.4.
 Following the criterion in Eq.
-(32), the minimum latent dimension required for retaining more than 99% of the energy in the dataset is $d=13$ modes, as shown in Fig.
-3a).
+(32), the minimum latent dimension required for retaining more than 99% of the energy in the training dataset is $d=13$ modes, as shown in Fig.
+3(a).
 (a) POD cumulative energy
@@
 (20) and (21).
-Panel a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set.
-Panel b) depicts the average (over the $c=10$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq.
-(33) with $d=13$ POD modes, in time over the training and testing datasets.
+Panel (a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set.
+Panel (b) depicts the average (over the $C=20$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq.
+(33) with $d=13$ POD modes, in time for the training and testing datasets.
 Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs.
@@
 (33) across the $k=1,\ldots,K$ time steps of each trajectory.
-Figure 3b) displays the average relative error overall $c=10$ cases with different initial conditions included in either the training or testing datasets.
-While in the beginning, the reconstruction error is more than 10%, after the first $100$ time steps, it reduces to 5% for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$.
-We note that this level of reconstruction error serves as a baseline for the proposed framework, since the timeseries forecasting with ROMs in the latent space introduces additional approximation errors.
-Having constructed the restriction and lifting operators (second and fourth step of the framework), we next train MVARs and LSTMs autoregressive ROMs to learn the latent dynamics; see Section 2.2.3.
-First, we obtained the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ from the density profiles of the training set, using the restriction operator.
-We then determined the lag window, using the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas the AIC favored (a wider, as expected) $w=9$.
-Both windows were used for training both MVARs and LSTMs models, denoted from now on as MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models.
-For performing sequence-to-one learning, we further divided the restricted training data set into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
-For the MVARs training, we first verified that the latent dynamics datasets $Y_{f},Y_{t}$ constitute stationary time series.
-Using the ADF test, we confirmed that all $d$ variables are stationary processes with a significance level of $p=0.001$.
+Fig.
+3(b) displays the average relative error over all the $C=20$ cases with different initial conditions included in either the training or testing datasets.
+While initially the reconstruction error is $\sim$15%, it decreases to 10% at the first 30s30\ $\text${s} and further reduces to $\sim$7% after 50s50\ $\text${s} for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$.
+We note that this level of reconstruction error serves as an infimum approximation error baseline, i.e., the best possible error for the proposed framework at the ambient–density profiles–space.
+ROM forecasts in the long term are therefore expected to surpass this baseline error due to the additional error contributions from model-form error and training/generalization errors.
+We next train MVAR and LSTM models to learn ROMs for the latent dynamics (see Section 2.2.3), where the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ are obtained from the density profiles of the training set via the restriction operator.
+To determine the lag window, we employed the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas AIC favored a wider (as expected) $w=9$.
+Both windows were used for training both MVAR and LSTM models, denoted from now on as MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models.
+The training data sets were divided into features $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$ and targets $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}$.
+For the MVARs training, we first verified that the latent dynamics datasets $Y_{f},Y_{t}$ constitute stationary time series; the ADF test confirmed stationarity with a significance level of $p<0.001$ for all $d$ variables.
 We next trained all MVAR and LSTM models using algorithms, hyperparameters and stopping criteria discussed in detail in Section 3.5.
 The training results are shown in Table 1 for the four ROMs, including the loss function values attained at the end of training and the computational time (in seconds) required.
-Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models.
-Loss function values at the end of training for MVARs $\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{c})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)$ in Eq.
-(25) and for LSTMs $\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{c})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)$ in Eq.
-(30), along with computational times (in seconds) required for training.
-For MVARs, the MSE in Eq.
+For the LSTMs, as each model was trained over 50 runs with different parameter initializations, we report in Table 1 the mean and the 10–90% error percentiles across the 50 runs.
+Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models at the latent space.
+Loss functions $\mathcal{L}_{\text{MVAR}}\left(\cdot\right)$ in Eq.
+(25) for MVARs and $\mathcal{L}_{\text{LSTM}}\left(\cdot\right)$ in Eq.
+(30) for LSTMs at the end of training, along with computational times (in seconds) required for training.
+For the LSTMs, the mean and 10–90% error percentiles (in parentheses) are reported across the 50 training processes with different parameter initializations.
+Additionally, the MSE loss function of MVARs in Eq.
 (34) is reported for comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
-Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s) $\downarrow$
-MVAR(4) $1.66\times 10^{-3}$ $1.96\times 10^{-11}$ - 0.0055
-MVAR(9) $1.01\times 10^{-3}$ $7.38\times 10^{-12}$ - 0.042
-LSTM(4) - - $8.64\times 10^{-7}$ 180
-LSTM(9) - - $3.51\times 10^{-7}$ 443
-Since the MVARs loss function is $L_{2}$ based, we further report the MSE, computed in Eq.
-(34), for MVARs to allow comparison with the MSE-based loss of the LSTMs.
-Table 1 indicates that all the optimization problems converge, as evident from the low loss function values attained.
+Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s)
+MVAR(4) $1.91\times 10^{-4}$ $1.41\times 10^{-13}$ - $0.0021$
+MVAR(9) $1.34\times 10^{-4}$ $7.05\times 10^{-14}$ - $0.0045$
+LSTM(4) - - $2.35\ (2.22,2.50)\times 10^{-5}$ 252(236,302)252\ (236,302)
+LSTM(9) - - $1.81\ (1.68,1.93)\times 10^{-5}$ 310(302,326)310\ (302,326)
+Since the loss function of MVARs is $L_{2}$-based, we additionally compute the MSE in Eq.
+(34) to enable direct comparison with the MSE-based LSTM loss.
 MVARs are able to reach much smaller loss function values than LSTMs, providing a hint for their better accuracy.
 As expected, LSTMs require significantly higher computational time for training in comparison to MVARs.
-Table 1 further shows that the training of both types of ROM converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
-Having constructed and trained all components of our framework, we now evaluate its end-to-end performance, using the forecasting closed-loop time-stepper in Eq.
-(35) for long-horizon forecasting.
-In particular, we quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
-(36) between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ for the $c$, unseen in training, cases–of different initial conditions–included in the testing set $X_{ts}$ across $k$ time steps.
-Table 2 provides a summary of the relative errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ for all four ROMs, including the median and the 10-90 percentiles, overall cases and time steps of the testing set.
-Table 2: End-to-end numerical accuracy of the closed-loop framework across the testing set, using the trained MVARs and LSTMs models for the latent dynamics.
-The relative $L_{1}$, $L_{2}$ and $L_{\infty}$ errors of the high-dimensional density space, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
-(36) are reported, for the four MVAR(4), MVAR(9), LSTM(4), LSTM(9) latent dynamics models; median and 10-90 percentiles (in parenthesis) are shown overall $C=10$ cases and $K=1091$ time steps.
+Table 1 further shows that the training of both types of ROMs converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
+One-Step/Open-loop Predictions
+Having constructed and trained all components of our framework, we now evaluate its end-to-end performance in one-step predictions in the ambient–density profile–space using the open-loop time-stepper in Eq.
+(35); note that for the LSTMs, we consider the model with the best training performance among the 50 model runs.
+We quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$, and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
+(37), computed between the “ground-truth” density profiles $x^{(c)}(t_{k})$ and the predicted/reconstructed ones $\hat{x}^{(c)}(t_{k})$ for the $C$, unseen in training, cases (of different initial conditions) included in the testing set $X_{ts}$ across time steps $k$.
+Table 2 provides a summary of the relative reconstructed errors in the ambient space, for one-step predictions, for all four ROMs, including the mean error and the 10–90% error percentiles, aggregated over all cases and time steps in the testing set.
+Table 2: Open-loop (one-step) prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models.
+The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
+(37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models.
+Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
 Model $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
-MVAR(4) 0.3599 (0.2408, 0.4535) 0.2536 (0.1609, 0.3098) 0.2251 (0.1361, 0.3435)
-MVAR(9) 0.3783 (0.2865, 0.4655) 0.2765 (0.2040, 0.3367) 0.2569 (0.1935, 0.3330)
-LSTM(4) 0.6323 (0.4313, 0.7941) 0.4160 (0.2889, 0.5269) 0.3372 (0.2532, 0.5037)
-LSTM(9) 0.4646 (0.3172, 0.6820) 0.3185 (0.2151, 0.4626) 0.3010 (0.1993, 0.4393)
-As shown in Table 2, the prediction accuracy of the closed-loop time stepper with the use of MVAR models outperforms the one with LSTMs, demonstrating robust generalization with tighter 10-90 percentile ranges.
-In addition, while MVARs with shorter lag demonstrate slightly higher approximation accuracy, LSTMs significantly improve with higher lags, albeit remaining inferior to MVARs.
-We hereby note that, considering the baseline 5-10% $L_{2}$ relative error of POD reconstruction (see Fig.
-3), the $L_{2}$ errors are 3-10 times higher for the end-to-end closed-loop framework, underscoring the compounding effect of autoregressive prediction.
-To highlight the robust performance of the proposed closed-loop time-stepping framework over time, we further depict the relative $L_{2}$ and $L_{\infty}$ errors, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in time, across the multiple cases of different, unseen initial conditions in Figs.
-4 and 5 for MVARs and LSTMs, respectively.
-In particular, Figs.
-4,5 show the average relative errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10-90 percentile ranges, over the $c=10$ cases considered, across time steps $k=1,\ldots,1100$.
+MVAR(4) 0.127(0.072,0.201)0.127\ (0.072,0.201) 0.078(0.045,0.119)0.078\ (0.045,0.119) 0.073(0.039,0.117)0.073\ (0.039,0.117)
+MVAR(9) 0.127(0.072,0.200)0.127\ (0.072,0.200) 0.078(0.045,0.118)0.078\ (0.045,0.118) 0.072(0.039,0.116)0.072\ (0.039,0.116)
+LSTM(4) 0.280(0.265,0.325)0.280\ (0.265,0.325) 0.167(0.158,0.190)0.167\ (0.158,0.190) 0.138(0.121,0.167)0.138\ (0.121,0.167)
+LSTM(9) 0.309(0.297,0.351)0.309\ (0.297,0.351) 0.185(0.174,0.207)0.185\ (0.174,0.207) 0.135(0.115,0.167)0.135\ (0.115,0.167)
+As shown in Table 2, the open-loop (one-step) prediction accuracy with the use of MVAR models outperforms that of LSTMs; all LSTM relative errors are approximately twice as high as those of the MVAR models.
+However, LSTMs demonstrate narrower 10–90% error percentiles in comparison to MVARs, indicating less uncertainty in one-step predictions.
+Table 2 additionally shows that the accuracy of MVARs does not improve with a longer lag, while that of LSTMs slightly degrades.
 (a) Relative Errors with MVAR(4)
 (b) Relative Errors with MVAR(9)
-Figure 4: End-to-end numerical accuracy of the closed-loop framework across the testing set over time, using the trained MVAR models.
-The relative $L_{2}$ (red) and $L_{\infty}$ (green) errors of the high-dimensional density space, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
-(36) are displayed for the MVAR(4) (panel a) and MVAR(9) (panel b) latent dynamics models; average (solid) and 10-90 percentiles (dashed) are shown overall $C=10$ cases per time step.
-Figures 4a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both indicating that, in spite of the long-time prediction, the relative error remains bounded in values below 30%; 5-10% is the baseline error of POD reconstruction in Fig.
-3.
-In addition, the tight 10-90 percentile ranges indicate good generalization.
-Although both models exhibit comparable error profiles, the MVAR(4) model demonstrates slightly improved accuracy while using fewer parameters, as also reported in Table 2.
-Similar bounded results are reported in Fig. 5 (a,b), depicting the employment of LSTM(4) and LSTM(9) models for long-time predictions.
-(a) Relative Errors with LSTM(4)
-(b) Relative Errors with LSTM(9)
-Figure 5: End-to-end numerical accuracy of the closed-loop framework across the testing set over time, using the trained LSTM models.
-The relative $L_{2}$ (red) and $L_{\infty}$ (green) errors of the high-dimensional density space, $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
-(36) are displayed for the LSTM(4) (panel a) and LSTM(9) (panel b) latent dynamics models; average (solid) and 10-90 percentiles (dashed) are shown overall $C=10$ cases per time step.
-As expected by the results in Table 2, the use of LSTMs provides less accurate predictions in comparison to MVARs.
-However, LSTMs provide tighter 10-90 percentile ranges towards the end of time integration, indicating a more robust behavior in comparison to MVARs.
-In addition, as also reported in Table 2, a wider lag window makes the scheme with LSTMs more accurate, while that of MVARs less accurate.
-Nevertheless, when comparing across model classes, the MVAR(4) model achieves slightly lower relative errors overall, while also requiring significantly fewer trainable parameters and hyperparameters.
-This makes MVAR(4) not only more accurate in this setting but also more computationally efficient.
-For completeness, we provide a comparison of the “ground truth” normalized density fields $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ derived based on the closed-loop time-stepper in Eq.
-(35) for a particular case, in the testing data set, of an initially spread pedestrian group; Gaussian initialization corresponding to the 6th row of Table B.2.
-Figure 6 displays such a comparison for the MVAR(4) model in four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930) where the pedestrian crowd does not pass through the corridor boundary; similar comparisons are provided for the MVAR(9), LSTM(4) and LSTM(9) models in Figs.
-E.1, E.2 and E.3 of Appendix E, respectively.
+(c) Relative Errors with LSTM(4)
+(d) Relative Errors with LSTM(9)
+Figure 4: One-step (open-loop) prediction errors, in the ambient–density profile–space, for the testing set over time, using trained MVAR and LSTM models.
+The relative reconstructed $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq.
+(37) (using the one-step prediction in Eq.
+(35)) are shown for the MVAR(4) panel (a), MVAR(9) panel (b), LSTM(4) panel (c), and LSTM(9) panel (d) latent dynamics models.
+Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step.
+For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
+Fig.
+4 displays the average relative reconstructed errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10–90% percentile ranges, for one-step predictions, over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$.
+Fig.
+4(a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both demonstrating a remarkable approximation accuracy.
+The mean $L_{2}$ reconstructed relative errors are initially $\sim$15% and reduce to $\sim$7% after the first 50s50\ $\text${s} of the simulation.
+This level of errors matches the infimum approximation error baseline of POD reconstruction, shown in Fig.
+3(b).
+This very high approximation accuracy is not met by the LSTM(4) and LSTM(9) models, shown in Fig.
+4(c,d).
+In particular, the mean relative reconstructed $L_{2}$ errors of the LSTM(4) and LSTM(9) models are $\sim$17% and $\sim$18%, respectively, towards the end of the simulations.
+Fig.
+4 further confirms that the lag window does not affect the one-step prediction accuracy of MVARs, while it slightly reduces that of LSTMs.
+Recursive/Closed-loop Predictions
+For obtaining long-term (closed-loop) predictions with the proposed framework, we next adopt the recursive closed-loop time-stepper in Eq.
+(36), which is only supplied with the first $w$ observations and predicts the next time steps in an autoregressive fashion.
+To quantify the prediction accuracy, we again use the relative reconstructed $L_{1}$, $L_{2}$, and $L_{\infty}$ errors in the ambient space, computed for the cases included in the testing set $X_{ts}$ across time steps.
+Table 3 provides a summary of these reconstructed errors, including the mean and the 10–90% error percentiles, aggregated over all cases and time steps.
+Table 3: Closed-loop/recursive simulation errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models.
+The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
+(37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models.
+Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
+Model $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
+MVAR(4) 0.231(0.139,0.376)0.231\ (0.139,0.376) 0.153(0.087,0.267)0.153\ (0.087,0.267) 0.141(0.074,0.247)0.141\ (0.074,0.247)
+MVAR(9) 0.184(0.114,0.292)0.184\ (0.114,0.292) 0.124(0.073,0.209)0.124\ (0.073,0.209) 0.119(0.062,0.210)0.119\ (0.062,0.210)
+LSTM(4) 0.419(0.221,0.648)0.419\ (0.221,0.648) 0.284(0.139,0.512)0.284\ (0.139,0.512) 0.281(0.115,0.501)0.281\ (0.115,0.501)
+LSTM(9) 0.433(0.230,0.590)0.433\ (0.230,0.590) 0.310(0.145,0.460)0.310\ (0.145,0.460) 0.285(0.121,0.441)0.285\ (0.121,0.441)
+As shown in Table 3, the long-term (closed-loop) prediction accuracy of the proposed framework with MVAR models outperforms significantly that of LSTM models.
+In addition, MVARs demonstrate narrower 10–90% error percentiles than LSTMs, indicating more reliable long-term predictions.
+Table 3 further shows that the consideration of a longer lag makes MVARs more accurate, while slightly degrading the accuracy of LSTMs.
+We next assessed the evolution/accumulation of the relative reconstructed errors $L_{2}$ and $L_{\infty}$ over time for the closed-loop setting.
+Fig.
+5 depicts the average reconstructed relative errors in the ambient–density profile–space, and their 10–90% error percentile ranges across the multiple cases of different, unseen initial conditions in the testing set for all four ROMs.
+(a) Relative Errors with MVAR(4)
+(b) Relative Errors with MVAR(9)
+(c) Relative Errors with LSTM(4)
+(d) Relative Errors with LSTM(9)
+(e) Relative Errors with SFM, $1\%$ perturbation
+(f) Relative Errors with SFM, $1\%$ perturbation
+Figure 5: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the trained MVAR and LSTM models.
+Panels (a)–(d) display the relative $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq.
+(37) for MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models, respectively.
+Panels (e) and (f) provide a system’s “baseline” relative error estimation, obtained from the SFM simulations using $\pm 1\%$ perturbations in the initial conditions for the density profiles (see in Section 3.1).
+Mean relative error (solid) and 10–90% percentiles (dashed) are shown over $C=20$ cases per time step.
+For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
+Indicatively, for our illustrations, to provide a comparative “baseline” of the reconstructed error for tiny perturbations of the density profiles, we quantified the sensitivity of the SFM simulations per se with respect to small perturbations of the initial conditions (see in Section 3.1).
+Specifically, we introduced $\pm 1\%$ perturbations in the initial conditions of the density profiles of the reference test cases and computed the relative reconstructed errors between these perturbed initial conditions and their unperturbed counterparts.
+The MVAR models Fig.
+5(a,b) demonstrate a remarkable approximation accuracy, with the mean relative $L_{2}$ error remaining bounded below $\sim$14%(17%) for the MVAR(9)(MVAR(4)) models, despite the long-term horizon.
+The respective 10–90% error percentiles are bounded in (10, 30)% relative errors for the MVAR(4) model, while for the MVAR(9) model, the 10–90% error percentiles reduce further to (7, 21)% relative errors.
+The mean values of errors, and especially of the MVAR(9) model, is well within the 10–90% error percentiles of the system’s deviation/“baseline” 10–90% $L_{2}$ error percentiles (2, 18)% relative error, as quantified by the SFM simulations of the testing set using the $\pm$ 1% perturbed initial conditions for the density profiles, shown in Fig.
+5(e,f).
+In contrast, the LSTM models Fig.
+5(c,d) provide less accurate predictions than the MVAR models, with mean relative $L_{2}$ error plateauing at $\sim$35% for both LSTM(4) and LSTM(9) models, towards the end of the simulation, well above the expected “baseline” error bounds.
+For completeness, we directly compare the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted fields $\hat{x}^{(c)}(t_{k})$ generated by the closed-loop/recursive simulations.
+This analysis uses a case from the test dataset featuring a Gaussian initialization corresponding to the 6th row of Table B.2.
+Figure 6 displays such comparison for the MVAR(9) model at four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930), during which the pedestrian crowd does not pass through the corridor boundary.
+Panels (a,c,e,g) show the ground truth at these time steps, while panels (b,d,f,h) present the corresponding MVAR(9) model predictions; analogous results for the MVAR(4), LSTM(4) and LSTM(9) models are provided in Appendix E, Figs.
+E.1, E.2 and E.3, respectively.
 (a) Ground truth density at $t_{250}=62.5\,\text{s}$.
@@
 (h) Predicted density at $t_{930}=232.5\,\text{s}$.
-Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq.
-(35) using the MVAR(4) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2.
-Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
-As shown in Fig.
-6, the predicted density fields preserve the global structure and propagation direction observed in the “ground truth” density, demonstrating the framework’s ability to capture dominant crowd dynamics and to deliver meaningful long-horizon predictions.
-The predicted density fields additionally succeed in correctly respecting the obstacle region.
-Comparison with the other (less accurate) models in Figs.
-E.1, E.2, E.3 reveals that while the above two crucial for crowd flow characteristics are preserved, the LSTM(4) model (the least accurate one) tends to underestimate peak density values, introduce artificial density in low density regions (for coping with conservation of mass), and exhibit greater dissipation in regions of high density; these erroneous behaviors relax as more accurate models are used, i.e., the LSTM(9) and the MVAR(9) ones.
-As shown, the proposed framework provides a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics.
-In particular, for the benchmark case of $100$ pedestrians moving in the corridor for more than 275s275\ $\text${s}, the conventional numerical integration requires a computational time of 20s20\ $\text${s} for the microscopic SFM simulator simulation plus 30s30\ $\text${s} for the extraction of the density fields (in a coarser time discretization), summing up to a total of 50s50\ $\text${s}.
-In contrast, our framework requires just $0.0063$, 0.0161s0.0161\ $\text${s} for the MVARs models and $6.75$, 6.77s6.77\ $\text${s} for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively.
-Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under 0.1s0.1\ $\text${s} (as shown in Table 1).
+Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(36) using the MVAR(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
+The predictions from all four surrogate models successfully preserve the global structure and propagation direction of the ground truth density, demonstrating the framework’s capacity to capture dominant crowd dynamics and produce meaningful long-horizon predictions.
+The models also correctly respect the obstacle regions.
+The primary observed discrepancy is a minor underestimation of the peak density values coupled with the introduction of artificial density in low-density regions, a likely consequence of enforcing mass conservation via the POD restriction operator.
+This artifact is faintest in the MVAR(9) model, and remains only slightly more pronounced in the MVAR(4).
+By contrast, it becomes considerably stronger in the less accurate LSTM(4) and LSTM(9) models.
+MVAR vs. LSTM performance
+The comparison between model classes reveals that the use of MVARs provides a more accurate and reliable framework for long-term predictions than that of LSTMs.
+This is clearly demonstrated by their significantly lower errors and narrower error percentiles in the closed-loop setting (Table 3, Fig.
+5).
+Crucially, this performance gap widens substantially when compared to the open-loop results (Table 2, Fig.
+4), where LSTMs performed more competitively.
+For instance, the MVAR(9) model’s mean $L_{2}$ error across time increased from $\sim$7% (open-loop) to $\sim$14% (closed-loop) within the “baseline” reconstruction error from POD and perturbed initial conditions, respectively, while the LSTM(9) model’s error changing respectively from $\sim$19% to $\sim$35% is well above these “baselines”.
+This superior performance of MVARs is attributed to the challenge of error accumulation in closed-loop forecasting; the linear structure of the MVAR model proves to be much less susceptible to this instability than complex, nonlinear LSTMs.
+While LSTMs can theoretically capture richer nonlinear dynamics, their performance severely degrades due to the “curse of dimensionality” in their training, i.e., finding a global optimum, accompanied by a distribution shift [99, 100] (i.e., the error accumulation of the recursive scheme during inference, that shifts the distribution upon which the models were trained), despite their ability to learn the one-step dynamics effectively.
+In addition, there are several challenges regarding the “correct” initialization of LSTMs [101], especially in dynamical systems and multiscale modeling tasks.
+In contrast, the training of the linear MVAR models is based only on OLS, thus giving a global optimum.
+For the particular applications, the results suggest that MVARs are a much better choice for long-term horizon predictions in the latent space.
+This is a result in line also with other studies (see e.g., [62]).
+Beyond accuracy, the proposed framework offers a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics.
+In particular, for the benchmark case of $100$ pedestrians moving in the corridor for $250\text{s}$, the conventional numerical integration requires a computational time of $20\text{s}$ for the microscopic SFM simulator simulation plus $30\text{s}$ for the extraction of the density fields (in a coarser time discretization), summing up to a total of $50\text{s}$.
+In contrast, our framework requires just $0063\text{s}, 0.0$$0161\text{s}$ for the MVARs models and $6.75\text{s}$, $6.77\text{s}$ for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively.
+Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under $0.005\text{s}$ (as shown in Table 1).
 The MVARs variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts.
-These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches unfeasible.
-5 Conclusion
-In this work, we present a four-step data-driven framework that combines manifold learning and machine learning to construct a discrete evolution operator for emergent crowd dynamics.
-Our framework provides a highly computationally efficient and systematic approach for learning accurate ROMs for spatio-temporal emergent crowd dynamics.
-By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs—sidestepping the prohibitive computational cost of learning PDEs via for example convolutional neural networks (CNNs) or other high‐dimensional ML models such as DNNs.
-Working in latent space simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics.
+These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches infeasible.
+5 Conclusions
+In this work, we present a four-step data-driven framework that combines manifold and machine learning to construct a discrete evolution operator for emergent crowd dynamics.
+Our framework provides a computationally efficient and systematic approach for learning ROMs for spatio-temporal emergent crowd dynamics.
+By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs, bypassing the prohibitive computational cost of learning PDEs via, for example, CNNs or other high‐dimensional ML models such as DNNs.
+Working in latent space with delayed POD coordinates simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics.
+In fact, under the usual generic assumptions of Takens’ theorem, applying time-delay embeddings to POD modes produces a delay-coordinate latent model that is diffeomorphic to the original system dynamics (see also results and discussion in [67, 62, 102, 68, 69]).
 Thus, our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution.
 The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction.
-As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics without sacrificing accuracy or stability, making them an ideal tool for large‐scale, real-world crowd‐dynamics analysis and control.
-Finally, we note that linear MVARs exhibit a better performance in terms of the median relative $L_{1}$, $_{2}, LL_{\infty}$ errors when compared to LSTMs, thus offering a significantly lower complexity in their training and greater interpretability.
-This is in line with the results presented in other works [56, 94, 63, 57] which utilize PCA-based delay-coordinate embeddings to construct linear ROMs using in latent spaces to approximate complex nonlinear or chaotic dynamics.
+As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics.
+Critically, we demonstrate that linear MVAR models within our framework not only outperform more complex LSTMs in long-term predictive accuracy and reliability, but also achieve a computational speed-up of over three orders of magnitude compared to conventional simulation.
+This combination of interpretability, speed, and accuracy makes our approach, particularly with MVARs, an efficient approach for large‐scale, real-world crowd‐dynamics analysis and distributed control [103, 104, 19, 21, 24, 22, 105].
+Finally, our finding that simple linear MVAR models outperform nonlinear LSTM networks in the latent space in terms of complexity, training time, and interpretability is a significant insight.
+It aligns with a growing body of literature [62, 102, 69, 63] that demonstrates the power of linear ROMs constructed in judiciously chosen latent spaces (e.g., POD-based delay-coordinate embeddings, Koopman operator) for approximating complex nonlinear or chaotic dynamics, offering a promising path toward interpretable and efficient data-driven modeling.
+Acknowledgments
+C.S.
+acknowledges partial support by the PNRR MUR projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).
+I.K.
+acknowledges partially support by the US National Science Foundation.
 Appendix A Social Force Model
@@
 As a reminder, we consider the SFM including a goal-directed, a social, and a repulsive force term (the random force is neglected), reading:
-$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{i,j}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{i,B}(\bm{x}_{i})}_{\text{repulsive force}},$ (A.1)
-where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is their maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time.
+$m_{i}\frac{d{\bm{\dot{x}}}_{i}}{dt}=\underbrace{-\frac{1}{\tau_{i}}({v_{0}}_{i}\bm{e}_{i}-\bm{v}_{i})}_{\text{goal-directed force}}+\underbrace{\sum_{j\neq i}\mathbf{F}_{ij}(\bm{x}_{i},\bm{x}_{j})}_{\text{social force}}+\underbrace{\sum_{\text{obstacles}}\mathbf{F}_{iB}(\bm{x}_{i})}_{\text{repulsive force}},$ (A.1)
+where $m_{i}$ is the mass of the $i$-th pedestrian, ${v_{0}}_{i}$ is the maximum walking speed, $\bm{e}_{i}$ represents the desired direction (serving as a simplistic auxiliary/behavioral variable), and $\tau_{i}$ is a characteristic relaxation time.
 The social force term in Eq.
-(A.1) introduces a short-range repulsive interaction between pedestrians, to prevent collisions.
-The social force between pedestrians $i$ and $j$ is given by:
+(A.1) introduces a short-range repulsive interaction between pedestrians to maintain social distances and avoid physical contact [1].
+Its expression between two pedestrians $i$ and $j$ is given by:
 $\mathbf{F}_{ij}(\bm{x}_{i},\bm{x}_{j})=A_{i}\exp\left(\frac{r_{ij}-d_{ij}}{B_{i}}\right)\bm{n}_{ij}+k\,g(r_{ij}-d_{ij})\,\bm{n}_{ij}+\kappa\,g(r_{ij}-d_{ij})\,\Delta v_{ij}^{\mathrm{tang}}\,\bm{t}_{ij},$ (A.2)
-where $A_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction, $d_{ij}=\|\bm{x}_{i}-\bm{x}_{j}\|$ is the distance between pedestrians, $r_{i,j}\in\mathbb{R}$ is the sum of their radii (minimal comfortable distance) [64], $B\in\mathbb{R}^{+}$ is the parameter controlling the spatial decay rate of the repulsive force, $\bm{n}_{ij}=\big{(}n^{1}_{i,j},n^{2}_{i,j}\big{)}=\frac{\bm{x}_{i}-\bm{x}_{j}}{d_{ij}}$ is the unit vector pointing from pedestrian $j$ to pedestrian $i$, $g(x)$ is a function which ensures that physical forces are only active when pedestrians overlap, i.e., when $(d_{i,j}<r_{i,j})$.
-Moreover, the constant $k>0$ governs the strength of the linear contact (body) force, which causes repulsion between pedestrians when they are in contact.
-The third term is a sliding friction term that involves a coefficient $\kappa>0$, which scales with the tangential component of the relative velocity $\Delta v_{ij}^{\mathrm{tang}}$; this component is projected onto the unit tangential vector $\mathbf{t}_{i,j}=\big{(}-n^{2}_{i,j},n^{1}_{i,j}\big{)}$.
-The repulsive force in Eq.
-(A.1), accounting for the repulsion that a pedestrian might experience due to static objects $B$, follows a similar, to Eq.
-(A.2), form, reading:
-$\mathbf{F}_{iB}(\bm{x}_{i})=C_{i}\exp\left(-\frac{d_{iB}}{D_{i}}\right)\bm{n}_{iB}+k\,g(r_{i}-d_{iB})\,\bm{n}_{iB}.$ (A.3)
-Here, $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle and $D_{i}$ is its corresponding decay parameter.
-Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exert dynamic resistance.
+where the first term represents a repulsive force modelling the physiological desire for personal space, the second term represents a physical contact (body) force that is only active when pedestrians are in contact, and the third term represents a sliding friction force that occurs during physical contact.
+The repulsive force is modelled in Eq.
+(A.2) by an exponential term, where $A_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction, $d_{ij}=\|\bm{x}_{i}-\bm{x}_{j}\|_{2}$ is the euclidean distance between pedestrians, $r_{ij}=r_{i}+r_{j}\in\mathbb{R}$ is the sum of their radii (minimal comfortable distance [70]), $B_{i}\in\mathbb{R}^{+}$ is the parameter controlling the spatial decay rate of the repulsive force, and $\bm{n}_{ij}=\big(n^{1}_{ij},n^{2}_{ij}\big)=(\bm{x}_{i}-\bm{x}_{j})/d_{ij}$ is the unit vector pointing from pedestrian $j$ to pedestrian $i$.
+The physical contact force is modelled in Eq.
+(A.2) by a linear term, where $k>0$ controls the strength of the contact (body) force, and $g(x)=\text{max}(0,x)$ is a ramp function ensuring that physical forces are only active when pedestrians overlap, i.e., when $(d_{ij}<r_{ij})$.
+The friction force term in Eq.
+(A.2) is modelled by the same ramp function $g(x)$, scaled by a friction coefficient $\kappa>0$ and difference in the tangential velocities $\Delta v_{ij}^{\mathrm{tang}}=\|(\bm{v}_{i}-\bm{v}_{j})-\left((\bm{v}_{i}-\bm{v}_{j})\cdot\bm{n}_{ij}\right)\bm{n}_{ij}\|_{2}$; this component is projected onto the unit tangential vector $\mathbf{t}_{ij}=\big(-n^{2}_{ij},n^{1}_{ij}\big)$.
+The repulsive force in Eq. (A.1), accounting for the repulsion that a pedestrian might experience due to static objects $B$, follows a similar form to Eq. (A.2), reading:
+$\mathbf{F}_{iB}(\bm{x}_{i})=C_{i}\exp\left(-\frac{d_{iB}}{D_{i}}\right)\bm{n}_{iB}+k\,g(r_{i}-d_{iB})\,\bm{n}_{iB},$ (A.3)
+where $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle, $D_{i}$ is its corresponding decay parameter, and $d_{iB}$ is the euclidean distance between the pedestrian and the closest point of the object $B$; $\bm{n}_{iB}$ is the respective the unit vector.
+Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exerft dynamic resistance.
+Table A.1: Social Force Model simulation parameters.
+Simulation Parameters Value Simulation Parameters Value
+Number of Pedestrians ($N$) 100 Repulsion coefficient ($k$) $1.2\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
+Corridor length ($L_{x}$) $48\,\text{m}$ Sliding friction coefficient ($\kappa$) $2.4\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
+Corridor width ($L_{y}$) $12\,\text{m}$ Interaction strength ($A_{i}$) $2\times 10^{3}\,\text{N}$
+Obstacle height $3.6\,\text{m}$ Interaction range ($B_{i}$) $0.08\,\text{m}$
+Obstacle width $3.6\,\text{m}$ Interaction strength ($C_{i}$) $2\times 10^{3}\,\text{N}$
+Obstacle $x$-coord. center $25.8\,\text{m}$ Interaction range ($D_{i}$) $0.08\,\text{m}$
+Obstacle $y$-coord. center $1.8\,\text{m}$ Pedestrian mass ($m_{i}$) $80\,\text{kg}$
+Initial velocity in $x$-direction ($v_{x}$) $0\,\text{m/s}$ Relaxation time ($\tau$) $0.5\,\text{s}$
+Initial velocity in $y$-direction ($v_{y}$) $0\,\text{m/s}$ Pedestrian radius ($r_{i}$) $0.2\,\text{m}$
+Simulation time $250\,\text{s}$ Time Step ($\delta_{t}$) $2.5\times 10^{-2}\,\text{s}$
 For our simulations, the model parameters were calibrated based on established values from previous studies [1, 70].
-These works describe pedestrians with a standard average mass of 8080kg, for which a relaxation time $\tau=0.5\,\text{s}$ provides adequate behavior under normal conditions.
-The complete set of parameters governing repulsive potentials, including agent-agent and agent-obstacle interactions, along with the simulation environment specifications, are detailed in Table A.1.
-This parameter selection, grounded in empirical crowd studies, enables the simulation of diverse scenarios, including the present case study while maintaining consistency with validated bottleneck evacuations scenarios, and even though the parameters should vary individually, in the literature the parameters are often chosen equally for all pedestrians to maintain a reasonable number of parameters calibration [64].
-Table A.1: Simulation parameters.
-Simulation Parameters Value
-Number of Pedestrians 100
-Corridor length ($L_{x}$) $48\,\text{m}$
-Corridor width ($L_{y}$) $12\,\text{m}$
-Time Step $\delta_{t}$ $2.5\times 10^{-2}\,\text{s}$
-Simulation time $275\,\text{s}$
-Repulsion coefficient ($k$) $1.2\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
-Sliding friction coefficient ($\kappa$) $2.4\times 10^{5}\,\frac{\text{kg}\cdot\text{m}}{\text{s}^{2}}$
-Interaction strength ($A_{i}$) $2\times 10^{3}\,\text{N}$
-Interaction range ($B_{i}$) $0.08\,\text{m}$
-Interaction strength ($C_{i}$) $2\times 10^{3}\,\text{N}$
-Interaction range ($D_{i}$) $0.08\,\text{m}$
-Mass $80\,\text{kg}$
-Relaxation time ($\tau$) $0.5\,\text{s}$
-Pedestrian radius $0.2\,\text{m}$
-Initial velocity in $x$-direction ($v_{x}$) $0\,\text{m/s}$
-Initial velocity in $y$-direction ($v_{y}$) $0\,\text{m/s}$
-Obstacle width $3.6\,\text{m}$
-Obstacle height $3.6\,\text{m}$
-$x$-coordinate of center $25.8\,\text{m}$
-$y$-coordinate of center $1.8\,\text{m}$
+These works describe pedestrians with a standard average mass of 80kg80\ kg, for which a relaxation time $\tau=0.5\,\text{s}$ provides adequate behavior under normal conditions.
+The complete set of parameters governing repulsive potentials, including pedestrian-pedestrian and pedestrian-obstacle interactions, along with the simulation environment specifications, are detailed in Table A.1.
+This parameter selection, which is uniform for all pedestrians, ensures consistency with validated bottleneck evacuation scenarios.
+Although individual variations exist in reality, using a common set of parameters is a standard approach in the literature [70] to maintain a manageable calibration process while enabling the reliable simulation of diverse crowd dynamics, including the present case study.
 Appendix B Initial Conditions of Microscopic Distributions
-Here, we describe the initial conditions that we considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested.
+Here, we describe the initial conditions considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested.
 We considered zero initial velocities for all pedestrians, while we used the following distributions for initializing the position $\bm{x}_{i}=(x_{i},y_{i})\in\Omega$ of the $i$-th pedestrian.
-  • Positions are sampled from a two dimensional normal distribution with mean $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
+Positions are sampled from:
+  • Uniform distributions over the domain $x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\subset\Omega$:
+    $p(x_{i},y_{i})=\mathcal{U}\left([x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\right).$ (B.1)
+  • Two dimensional normal distributions with means $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
     $p(x_{i},y_{i})=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\left(-\frac{(x_{i}-\mu_{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y_{i}-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right),$ (B.2)
-  • Uniform density distribution: Positions are sampled uniformly over the domain $\Omega$
-    $p(x_{i},y_{i})=\mathcal{U}\left([x_{min},x_{max}]\times[y_{min},y_{max}]\right).$ (B.2)
-  • Cosine distribution;
-    $p(x_{i},y_{i})=\frac{1}{\pi\sigma_{x}\sigma_{y}}\cos\left(\frac{x_{i}-\mu_{x}}{\sigma_{x}}\right)\cos\left(\frac{y_{i}-\mu_{y}}{\sigma_{y}}\right).$ (B.3)
-  • Piecewise linear density distribution:
-    To initialize particle positions in a rectangular domain $x_{\min},x_{\max}]\times[y_{\min},y_{\max}]$, we assume a uniform distribution in $x$ and a non-uniform, piecewise linear distribution in $y$.
-This approach allows control over the particle concentration along the vertical axis.
-    Particles are sampled uniformly in the $x$-direction over the interval $x_{\min},x_{\max}]$.
-In the $y$-direction, particle positions are sampled from a custom piecewise linear distribution designed to increase density near the top and bottom boundaries.
-Specifically, we use a triangular distribution defined over the interval $[0,12]$ by
-    $p(y_{i})=\begin{cases}\frac{2}{12}\left(1-\frac{y_{i}}{6}\right),&0\leq y_{i}\leq 6,\\ \frac{2}{12}\left(\frac{y_{i}-6}{6}\right),&6<y_{i}\leq 12,\end{cases}$ (B.4)
-    which has peaks at $y=0$ and $y=12$, and a minimum at $y=6$.
-    To sample $y$-coordinates from this distribution, we use inverse transform sampling.
-First, a uniform random number $u\sim\mathcal{U}(0,1)$ is generated.
-If $u\leq 0.5$, the corresponding $y$ value is computed by solving $u=\frac{y}{12}(2-\frac{y}{6})$, yielding
-    $y=6\left(1-\sqrt{1-2u}\right).$ (B.5)
-    If $u>0.5$, we solve $u=\frac{1}{2}+\frac{(y-6)^{2}}{72}$, resulting in
-    $y=6+\sqrt{72(u-0.5)}.$ (B.6)
-    This method produces particle positions that are uniformly distributed along the horizontal axis and more concentrated near the top and bottom edges along the vertical axis.
-For generating the training/testing data set, we used a wide range of initial conditions, based on the described distributions;in particular, for the gaussian sampling, the means and standard deviations were selected relative to the domain size to ensure spatial coverage while avoiding boundary artifacts.
-Specifically, the means were placed at the geometric center or fractional positions within the domain (e.g., $0.5$ or $0.25$ of the domain width), while standard deviations were set as fractions of the domain dimensions (e.g., one-fifth or one-half of the width or height).
-This ensures that the pedestrian positions lie within the domain, while allowing for spatial clustering or multimodal patterns.
-The parameters were chosen heuristically to reflect plausible crowd configurations and to introduce variety in the training and testing datasets.
-the ones used to construct the training are shown in Table B.1, while the ones for the testing data set are shown in Table B.2.
-Table B.1: Microscopic distribution configurations and parameters for training the LSTM and MVAR models.
-Distribution Type Parameters
-Data set 1: Gaussian $\begin{aligned} &\mu_{x}=7.5,\;\mu_{y}=5.0,\\ &\sigma_{x}=2.2,\;\sigma_{y}=2.0\end{aligned}$
-Data set 2: Uniform $\mathcal{U}\left([2,15]\times[3,6]\right)$
-Data set 3: Gaussian $\begin{aligned} &\mu_{x}=8.4,\;\mu_{y}=4.0,\\ &\sigma_{x}=1.8,\;\sigma_{y}=1.3\end{aligned}$
-Data set 4: Gaussian $\begin{aligned} &\mu_{x}=12.5,\;\mu_{y}=4.0,\\ &\sigma_{x}=7,\;\sigma_{y}=1.1\end{aligned}$
-Data set 5: Double Gaussian $\begin{aligned} &\mu_{x1}=17.3,\;\mu_{x2}=13.6,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.7,\\ &\sigma_{y}=1.2;\;\mu_{y}=7.0\end{aligned}$
-Data set 6: Double Gaussian $\begin{aligned} &\mu_{x1}=19.7,\;\mu_{x2}=13.7,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.62,\\ &\sigma_{y}=4.0;\;\mu_{y}=2.33\end{aligned}$
-Data set 7: Double Gaussian $\begin{aligned} &\mu_{x1}=32.2,\;\mu_{x2}=27.75,\\ &\sigma_{x_{1}}=\sigma_{x_{2}}=0.35,\\ &\sigma_{y}=4.0;\;\mu_{y}=1.0\end{aligned}$
-Data set 8: Piecewise Linear $x_{i}\sim\mathcal{U}\left(30,47\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
-Data set 9: Piecewise Linear $x_{i}\sim\mathcal{U}\left(15,35\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
-Data set 10: Cosine $\begin{aligned} &\mu_{x}=25,\;\mu_{y}=3.5,\\ &L_{x}=10,\;L_{y}=2.5\end{aligned}$
-Table B.2: Microscopic distribution configurations and parameters for testing the LSTM and MVAR models.
-Configuration Parameters
-Dataset 1: Double Gaussian $\begin{aligned} &\mu_{x1}=41.2,\;\mu_{x2}=33.7,\\ &\sigma_{x1}=\sigma_{x2}=0.5,\\ &\sigma_{y}=2.0,\;\mu_{y}=9.0\end{aligned}$
-Dataset 2: Double Gaussian $\begin{aligned} &\mu_{x1}=11.7,\;\mu_{x2}=5.2,\\ &\sigma_{x1}=\sigma_{x2}=0.4,\\ &\sigma_{y}=1.0,\;\mu_{y}=9.0\end{aligned}$
-Dataset 3: Double Gaussian $\begin{aligned} &\mu_{x1}=42.5,\;\mu_{x2}=37.5,\\ &\sigma_{x1}=\sigma_{x2}=0.3,\\ &\sigma_{y}=2.0,\;\mu_{y}=5.0\end{aligned}$
-Dataset 4: Gaussian $\begin{aligned} &\mu_{x}=36.0,\;\mu_{y}=4.0,\\ &\sigma_{x}=2.4,\;\sigma_{y}=1.5\end{aligned}$
-Dataset 5: Gaussian $\begin{aligned} &\mu_{x}=23.0,\;\mu_{y}=8.5,\\ &\sigma_{x}=2.0,\;\sigma_{y}=2.5\end{aligned}$
-Dataset 6:Gaussian $\begin{aligned} &\mu_{x}=7.85,\;\mu_{y}=8.0,\\ &\sigma_{x}=4.3,\;\sigma_{y}=2.0\end{aligned}$
-Dataset 7: Cosine $\begin{aligned} &\mu_{x}=40,\;\mu_{y}=8.0,\\ &L_{x}=5,\;L_{y}=2.0\end{aligned}$
-Dataset 8: Uniform $\mathcal{U}\left([10,20]\times[4,9]\right)$
-Dataset 9: Piecewise Linear $x_{i}\sim\mathcal{U}\left(30,44\right)$, $y_{i}\sim p(y_{i})$ in Eq. (B.4)
-Dataset 10: Cosine $\begin{aligned} &\mu_{x}=39,\;\mu_{y}=6.5,\\ &L_{x}=7,\;L_{y}=2.5\end{aligned}$
-Appendix C Kernel Density Estimation
+  • “Double bell-shaped” independent joint distributions, which are comprised by normal distributions along the yy–axis with mean $\mu_{y}$ and standard deviation $\sigma_{y}$, and a mixture of normal distributions along the xx–axis with means $(\mu_{x_{1}},\mu_{x_{2}})$ and standard deviations $(\sigma_{x_{1}},\sigma_{x_{2}})$:
+    $p(x_{i},y_{i})$ $=\frac{1}{\sqrt{2\pi}\,\sigma_{y}}\,\exp\!\left(-\frac{(y-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right)$ (B.3)
+    $\quad\times\frac{1}{2}\left[\frac{1}{\sqrt{2\pi}\,\sigma_{x_{1}}}\exp\!\left(-\frac{(x-\mu_{x_{1}})^{2}}{2\sigma_{x_{1}}^{2}}\right)+\frac{1}{\sqrt{2\pi}\,\sigma_{x_{2}}}\exp\!\left(-\frac{(x-\mu_{x_{2}})^{2}}{2\sigma_{x_{2}}^{2}}\right)\right].$
+  • Cosine distributions centered at $(c_{x},c_{y})$ with scale parameters $(s_{x},s_{y})$:
+    $p(x_{i},y_{i})=\frac{1}{\pi s_{x}s_{y}}\cos\left(\frac{x_{i}-c_{x}}{s_{x}}\right)\cos\left(\frac{y_{i}-c_{y}}{s_{y}}\right).$ (B.4)
+The training and testing data sets were generated from $C=20$ initial conditions, with 5 samples drawn from each of the four distributions described above.
+To ensure plausible and varied crowd configurations, the parameters for these distributions were selected heuristically to provide a broad spatial coverage while mitigating boundary artifacts.
+Furthermore, during the sampling process, pedestrian positions were resampled if they were within a critical distance equal to $2r_{i}$ of an existing pedestrian or a wall/obstacle to prevent initial overlaps and collisions.
+The specific parameter configurations used for the training and testing data sets are detailed Table B.1 and Table B.2, respectively.
+Table B.1: Microscopic initial distribution configurations and parameters for constructing the training dataset.
+Case c Distribution Type Parameters
+1 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=15,y_{\text{min}}=3,y_{\text{max}}=9$
+2 Uniform, Eq. (B.1) $x_{\text{min}}=33,x_{\text{max}}=46,y_{\text{min}}=3,y_{\text{max}}=9$
+3 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=20,y_{\text{min}}=4,y_{\text{max}}=8$
+4 Uniform, Eq. (B.1) $x_{\text{min}}=28,x_{\text{max}}=46,y_{\text{max}}=4,y_{\text{max}}=8$
+5 Uniform, Eq. (B.1) $x_{\text{min}}=3,x_{\text{max}}=17,y_{\text{max}}=4,y_{\text{max}}=8$
+6 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+7 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
+8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
+9 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=5,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=8,\;\mu_{y}=7,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=0.8$
+12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
+13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=28,\;\mu_{y}=6,\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=32,\;\sigma_{x_{2}}=1.1$
+14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=36,\;\mu_{y}=6,\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=40,\;\sigma_{x_{2}}=1.2$
+15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=9,\;\mu_{y}=6,\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=1$
+16 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
+17 Cosine, Eq. (B.4) $c_{x}=20,\;c_{y}=8,s_{x}=10,\;s_{y}=3$
+18 Cosine, Eq. (B.4) $c_{x}=30,\;c_{y}=7,s_{x}=10,\;s_{y}=3$
+19 Cosine, Eq. (B.4) $c_{x}=40,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
+20 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=9,\;s_{y}=3$
+Table B.2: Microscopic initial distribution configurations and parameters for constructing the testing dataset.
+Case c Distribution Type Parameters
+1 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=16,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
+2 Uniform, Eq. (B.1) $x_{\text{min}}=32,\;x_{\text{max}}=46,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
+3 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=19,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
+4 Uniform, Eq. (B.1) $x_{\text{min}}=29,\;x_{\text{max}}=46,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
+5 Uniform, Eq. (B.1) $x_{\text{min}}=4,\;x_{\text{max}}=17,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
+6 Gaussian, Eq. (B.2) $\mu_{x}=12,\;\mu_{y}=7,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
+7 Gaussian, Eq. (B.2) $\mu_{x}=11,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=2$
+8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
+9 Gaussian, Eq. (B.2) $\mu_{x}=39,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=7,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=7,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.1,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=0.9$
+12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
+13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=29,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=33,\;\sigma_{x_{2}}=1.1$
+14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=37,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=41,\;\sigma_{x_{2}}=1.2$
+15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\;\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=1$
+16 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=4,\;s_{x}=8,\;s_{y}=3$
+17 Cosine, Eq. (B.4) $c_{x}=25,\;c_{y}=6,\;s_{x}=10,\;s_{y}=3$
+18 Cosine, Eq. (B.4) $c_{x}=35,\;c_{y}=5,\;s_{x}=10,\;s_{y}=3$
+19 Cosine, Eq. (B.4) $c_{x}=45,\;c_{y}=7,\;s_{x}=5,\;s_{y}=3$
+20 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=6,\;s_{x}=7,\;s_{y}=3$
+Appendix C Density fields extraction via Kernel Density Estimation
 Here, we provide additional details for the first step of the proposed framework, regarding the extraction of continuous density profiles from discrete pedestrians’ positions.
-As already discussed in Section 2.2.1, the Kernel Density Estimation (KDE) is used for this task, following the general form in Eq.
-(6).
-Here, we elaborate on the kernel selection, bandwidth tuning, and adjustments for boundaries and obstacles.
+As already discussed in Section 2.2.1, we used KDE for this task, following the general form in Eq.
+(6); here, we elaborate on the kernel selection, bandwidth tuning, and adjustments for boundaries and obstacles.
 The choice of kernel function directly influences the smoothness and continuity of the density estimate.
 Common options include the uniform, triangular, Epanechnikov, and Gaussian kernels [106, 107].
-The Epanechnikov kernel is theoretically optimal for minimizing mean integrated squared error (MISE) [97], but its bounded support can limit flexibility in practice.
+The Epanechnikov kernel is theoretically optimal for minimizing mean integrated squared error [108], but its bounded support can limit flexibility in practice.
 In contrast, the Gaussian kernel is widely preferred due to its smoothness, infinite support, and analytical tractability.
@@
 Silverman’s rule of thumb [106], originally formulated for univariate distributions, extends naturally to multivariate settings, indicating an optimal bandwidth:
-$h_{i}=\left(\frac{4}{N(d+2)}\right)^{1/(d+4)}\sigma_{i}\quad(i=x,y).$ (C.2)
-where $N$ is the number of data points, $d$ is the number of spatial dimensions, in this case $d=2$ and $\sigma_{i}$ is the standard deviations of the pedestrian positions.
+$h_{i}=\sigma_{i}\left(\frac{4}{N(d+2)}\right)^{1/(d+4)},$ (C.2)
+where $N$ is the number of data points, $d$ is the number of spatial dimensions, in this case $d=2$, and $\sigma_{i}$ is the standard deviation of the pedestrian positions along $i=x,y$ spatial directions.
 However, in our case, Silverman’s rule resulted in wide bandwidths, which did not allow us to consider sufficient detail in the density profiles.
@@
 $\Omega^{\text{ext}}=[a-\Delta x,b+\Delta x]\times[c,d],$ (C.3)
-while the support in the $y$-direction remains unbounded, i.e., $\Omega_{y}=(-\infty,\infty)$.
+while the support in the $y$-direction remains unbounded.
 To enforce continuity and reduce boundary-induced artifacts in the $x$-direction, we further augment the observed positions with “ghost” particles near the domain edges.
@@
 (h) Predicted density at $t_{930}=232.5\,\text{s}$.
-Figure E.1: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq.
-(35) using the MVAR(9) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2.
-Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
+Figure E.1: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(36) using the MVAR(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
 (a) Ground truth density at $t_{250}=62.5\,\text{s}$.
@@
 (h) Predicted density at $t_{930}=232.5\,\text{s}$.
-Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq.
-(35) using the LSTM(4) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2.
-Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
+Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(36) using the LSTM(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
 (a) Ground truth density at $t_{250}=62.5\,\text{s}$.
@@
 (h) Predicted density at $t_{930}=232.5\,\text{s}$.
-Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ via the closed-loop time-stepper in Eq.
-(35) using the LSTM(9) model, normalized densities for an unseen case of the testing set; initialization at the 6th row of Table B.2.
-Panels a,c,e, and g show the “ground truth” density distribution in $\Omega$, while panels b,d,f, and h, show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
+Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(36) using the LSTM(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.

### v3

@@
-Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning
+arXiv:2508.03926v3[cs.LG] 28 Feb 2026
 Abstract
 Bridging the microscopic/individual and macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control.
-Here, we propose a combined manifold and machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations.
+Here, we propose a manifold-informed machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations.
 The proposed framework builds upon our previous works on the development of next-generation Equation-free algorithms for dealing with the “curse of dimensionality” when learning surrogate models for high-dimensional and multiscale systems.
@@
 In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation.
-In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distributions.
-The third step involves learning reduced-order surrogate models in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs).
-Finally, we reconstruct the crowd dynamics in the high-dimensional space in terms of macroscopic density profiles.
-We demonstrate that the POD reconstruction of the density distribution via Singular Value Decomposition conserves the mass.
-With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation for the evolution of the density distribution.
-For our illustrations, we use the Social Force Model to generate data in a corridor with an obstacle, imposing periodic boundary conditions.
+In the second step, based on manifold learning, we construct a map from the density-field space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distributions.
+The third step involves learning reduced-order surrogate models in the latent space using machine learning techniques, particularly Long Short-Term Memory networks and Multivariate Autoregressive models.
+Finally, we reconstruct the crowd dynamics in the high-dimensional space with POD, demonstrating that the POD reconstruction conserves the mass.
+Thus, with this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation for the evolution of the density distribution.
+For our illustrations, we used the Social Force Model to generate data in a corridor with an obstacle, imposing periodic boundary conditions in two scenarios: (i) an unidirectional flow, and (ii) a counterflow.
 The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations.
-Notably, linear MVAR models surpass nonlinear LSTMs in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
 Keywords Crowd Dynamics, Multiscale modelling, Machine Learning, Numerical Analysis, Conservation laws, Reduced Order Models
@@
 Typically, models are formulated at three distinct levels: the microscopic (individual), mesoscopic (kinetic), and macroscopic (hydrodynamic) scales.
-Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [1] and its variants (see e.g.
+Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [62, 61] and its variants (see e.g.
 [32, 111]), cellular automata [87, 21, 41], etc.
@@
 Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM models.
-These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’ theorem [65] for reconstructing the phase-space structure (here in the latent space).
+These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’/Whitney’s embedding theorems [106, 99] for reconstructing the phase-space structure (here in the latent space).
 The concept of the approach (i.e., using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [75, 102] (see also results and discussion in [91, 49, 102, 71, 9]).
-In crowd dynamics, behavioral characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Takens’ theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics.
-Finally, we lift the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis.
-Crucially, this configuration ensures explicit mass conservation—a key physical constraint for any effective continuum description.
-For our numerical experiments, we employ the Social Force Model (SFM) [70] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions.
+In crowd dynamics, behavioral characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Takens’/Whitney’s embedding theorems [106, 99], to reconstruct the effective phase space containing these (unobservable) latent dynamics.
+Finally, we lift on-demand the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis.
+Crucially, this configuration ensures explicit mass conservation-—a key physical constraint for any effective continuum description.
+For our numerical experiments, we employ the SFM [62, 61] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions.
+We particularly consider a unidirectional pedestrian flow and a more challenging, dense counterflow of two pedestrian groups.
 The performance and generalization ability of the proposed framework are assessed on unseen initial conditions.
@@
 where $\Phi_{{\delta}t}:\mathbb{R}^{M}\to\mathbb{R}^{M}$ is the evolution operator at the high-dimensional space.
-In this work, we learn the discrete evolution operator $F_{\delta t}$ using autoregressive ROMs, thus effectively learning $\Phi_{\Delta t}$ via the restriction and lifting operators, as:
-$\mathbf{u}(t+\Delta t)=\mathcal{L}\Bigl(F_{\delta t}\big(\mathcal{R}(\mathbf{u}(t)),\ldots,\mathcal{R}(\mathbf{u}(t-w\delta t))\big)\Bigl),$ (5)
-where the learned ROMs account for $w$ delayed embeddings $\mathcal{R}(\mathbf{u}(t-j\delta t))$ for $j=1,\ldots,w$.
+In this work, we learn the discrete evolution operator $F_{\delta t}$ using autoregressive ROMs, thus effectively learning $\Phi_{{\delta}t}$ as:
+$\mathbf{u}(t+\delta t)=\mathcal{L}\left(\mathbf{y}_{d}(t+\delta t)\right),\quad\mathbf{y}_{d}(t+\delta t)=F_{\delta t}\big(\mathbf{y}_{d}(t),\ldots,\mathbf{y}_{d}(t-w\delta t)\big),$ (5)
+where the learned ROMs account for $w$ delayed latent points $\mathbf{y}_{d}(t-j\delta t)$ for $j=1,\ldots,w$.
+The latent points are obtained from previous predictions of $F_{\delta t}$ resulting in a closed-loop/recursive forecasting operator requiring only the initial $j=1,\ldots,w$ embeddings $\mathcal{R}(\mathbf{u}(j\delta t))$.
+This formulation enables efficient on-demand reconstruction of the high-dimensional space, as $F_{\delta t}$ evolves in latent space and lifting is performed only when a high-dimensional reconstruction is demanded.
 A key distinction between our proposed next generation EF and the traditional EF framework is that, in the latter, one constructs local (numerical) maps and computes the quantities needed for bifurcation analysis on demand via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured usually by a few moment distributions).
@@
 (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves.
-Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator, detailed in Appendix A.
+Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator, detailed in A.
 We next employ KDE [38] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
@@
 where $K_{h}$ is a smooth non-negative kernel function satisfying $\int_{\Omega}K(\mathbf{x})\,d\mathbf{x}=1$.
-Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(c-d)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq.
+Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(d-c)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq.
 (6) density fields are computed as $\rho(\mathbf{x},t)=[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $\rho((x_{i},y_{j}),t)$ corresponds to the density computed at the grid point $(x_{i},y_{j})$, for $i=1,\dots,n_{x}$, $j=1,\dots,n_{y}$.
 Then, the total mass in the domain is given by $S(t)=\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y$.
-While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct the density fields in Eq.
-(6).
-Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
+The above procedure can be applied in multiple groups of pedestrians to derive (normalized) group macroscopic density fields.
+While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct density fields.
+Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid points of the obstacle), and the treatment of periodic boundary conditions are provided in B.
 2.2.2 The Restriction and Lifting Operators
@@
 To construct the restriction operator, we apply POD [104] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq.
-(6), for different initial conditions; see Appendix B for details.
+(6), for different initial conditions; see C for details.
 Denoting the densities by $\rho^{(c)}(\mathbf{x},t_{0}+k\delta t)\equiv\rho^{(c)}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$, where $\delta t$ is a microscopic inner simulation step/observation sampling time of an experiment/real data, $t_{0}$ is the initial time, and the superscript $(c)$ denotes the case $c=1,\ldots,C$ of the different initial conditions considered, the training dataset generated by microscopic runs is represented by the matrix $X\in\mathbb{R}^{n_{c}\times(C\cdot K)}$:
-$X=\Bigl\{\rho^{(c)}(\mathbf{x},t_{k})\,:\,k=0,\ldots,K,\,c=1,\ldots,C\Bigr\}.$ (8)
+$X=\Bigl\{\rho^{(c)}(\mathbf{x},t_{k})\,:\,k=1,\ldots,K,\,c=1,\ldots,C\Bigr\}.$ (8)
 Compactly, we represent the above data set as:
@@
 $\bar{X}=XH=U\Sigma V^{\top},\quad H=I_{n_{t}}-\frac{1}{n_{t}}\mathbf{1}_{n_{t}}\mathbf{1}_{n_{t}}^{\top},$ (10)
-where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular vectors, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim.
+where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular values, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim.
 unity matrix and $\mathbf{1}_{n_{t}}$ is the $n_{t}$-dim.
@@
 Thus, the projection of $X$ on the latent space provides the low-dimensional embedding:
-$Y_{d}=U_{d}^{\top}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
+$Y_{d}=U_{d}^{\top}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}.$ (11)
 For the lifting operator that maps $Y_{d}$ to the space of density profiles, we demonstrate that the POD reconstruction (aka the POD lifting operator) preserves the original total density.
@@
 Hence, the POD reconstruction preserves the total original density. ∎
-In summary, given a density distribution $\rho(\mathbf{x},t)$ obtained via Eq. (6), we obtain the normalized density along the $n_{c}$ grid points of $\Omega$ as:
-$x(t)=\dfrac{1}{S(t)}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}},$ (19)
+We next consider a more challenging case of two interacting pedestrian groups in the domain $\Omega$, a typical scenario in dense counterflow settings.
+As in the one-group case, we collect the density profiles $\rho^{(c)}_{l}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$ of each group $l=1,2$ via Eq.
+(6), and represent them in the snapshot matrices:
+$X^{(1)}=\Big[x_{1}^{(1)},\ldots,x_{n_{t}}^{(1)}\Big]\in\mathbb{R}^{n_{c}\times n_{t}},\quad X^{(2)}=\Big[x_{1}^{(2)},\ldots,x_{n_{t}}^{(2)}\Big]\in\mathbb{R}^{n_{c}\times n_{t}},$ (19)
+with columns $x^{(l)}_{k}=\rho^{(c)}_{l}(\mathbf{x},t_{k})/S_{l}(t_{k})\in\mathbb{R}^{n_{c}}$ for $k=1,\ldots,n_{t}=C\cdot K$.
+Applying the single‑group POD (Proposition 1) separately to $X^{(1)}$ and $X^{(2)}$ yields reconstructions that preserve each group’s density, but ignore any statistical coupling between the groups.
+To retain mass‑preservation while also incorporating dependencies among groups, we construct enriched projection bases that combine the individual POD modes with cross‑covariance modes, as shown in the following proposition.
+Proposition 2.
+Let us assume the snapshot matrices $X^{(1)},\,X^{(2)}\in\mathbb{R}^{n_{c}\times n_{t}}$ with columns being normalized density fields of two interacting population groups.
+Furthermore, let $W\in\mathbb{R}^{n_{c}\times m}$ and $T\in\mathbb{R}^{n_{c}\times m}$ be the left and right singular vectors of the cross-covariance matrix $C^{(12)}=\frac{1}{n_{t}}\bar{X}^{(1)}\left(\bar{X}^{(2)}\right)^{\top}\in\mathbb{R}^{n_{c}\times n_{c}}$, and consider the augmented orthonormal projection bases:
+$\tilde{U}^{(1)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{1}}^{(1)}&\hat{W}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{1}+m+1)},\quad\tilde{U}^{(2)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{2}}^{(2)}&\hat{T}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{2}+m+1)},$ (20)
+where $U_{d_{1}}^{(1)}\in\mathbb{R}^{n_{c}\times d_{1}}$ and $U_{d_{2}}^{(2)}\in\mathbb{R}^{n_{c}\times d_{2}}$ are the first $d_{1}$ and $d_{2}$ left singular vectors from the SVD of $\bar{X}^{(1)}$ and $\bar{X}^{(2)}$, respectively, and the terms $\hat{W}=W_{\perp}(W_{\perp}^{\top}W_{\perp})^{-1/2}$ and $\hat{T}=T_{\perp}(T_{\perp}^{\top}T_{\perp})^{-1/2}$ are computed by the orthogonal projections $W_{\perp}=\left(I_{n_{c}}-\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}/n_{c}-U_{d_{1}}^{(1)}(U_{d_{1}}^{(1)})^{\top}\right)W\in\mathbb{R}^{n_{c}\times m}$ and $T_{\perp}=\left(I_{n_{c}}-\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}/n_{c}-U_{d_{2}}^{(2)}(U_{d_{2}}^{(2)})^{\top}\right)T\in\mathbb{R}^{n_{c}\times m}$.
+If the original densities are mass-preserving along time steps, i.e., if $\mathbf{1}^{\top}_{n_{c}}X^{(1)}=\mathbf{1}^{\top}_{n_{t}}$ and $\mathbf{1}^{\top}_{n_{c}}X^{(2)}=\mathbf{1}^{\top}_{n_{t}}$, then the reconstructed densities computed by projecting $X^{(1)}$ and $X^{(2)}$ onto the augmented bases as:
+$\tilde{X}^{(1)}=\tilde{U}^{(1)}\left(\tilde{U}^{(1)}\right)^{\top}\bar{X}^{(1)}+X^{(1)}(I_{n_{t}}-H),\quad\tilde{X}^{(2)}=\tilde{U}^{(2)}\left(\tilde{U}^{(2)}\right)^{\top}\bar{X}^{(2)}+X^{(2)}(I_{n_{t}}-H),$ (21)
+are also mass-preserving, i.e., $\mathbf{1}^{\top}_{n_{c}}\tilde{X}^{(1)}=\mathbf{1}^{\top}_{n_{t}}$ and $\mathbf{1}^{\top}_{n_{c}}\tilde{X}^{(2)}=\mathbf{1}^{\top}_{n_{t}}$.
+Proof.
+Assume the original densities are mass-preserving for each group. From the properties of the single-population POD (see Eq. (15) and Eq. (16)), it follows that:
+$\mathbf{1}_{n_{c}}^{\top}\bar{X}^{(l)}=\boldsymbol{0}^{\top}_{n_{t}},\quad\mathbf{1}_{n_{c}}^{\top}U^{(l)}_{d_{l}}=\mathbf{0}_{d_{l}}^{\top},$ (22)
+where $\bar{X}^{(l)}=X^{(l)}H$ is the centered data matrix and $U_{d_{l}}^{(l)}\in\mathbb{R}^{n_{c}\times d_{l}}$ are the first $d_{l}$ left singular vectors from the SVD of $\bar{X}^{(l)}$ for the $l=1,2$ groups.
+Let’s consider now the SVD of the cross-covariance matrix:
+$C^{(12)}=\frac{1}{n_{t}}\bar{X}^{(1)}\left(\bar{X}^{(2)}\right)^{\top}=W\Sigma_{C^{(12)}}T^{\top},$ (23)
+with $W\in\mathbb{R}^{n_{c}\times m}$, $\Sigma_{C^{(12)}}\in\mathbb{R}^{m\times m}$ and $T\in\mathbb{R}^{n_{c}\times m}$, where $m$ denotes the number of retained cross-covariance modes.
+Each column of $C^{(12)}$ is a linear combination of the columns of $\bar{X}^{(1)}$, and each row a linear combination of the rows of $\bar{X}^{(2)}$.
+Consequently:
+$\mathcal{R}(W)\equiv\mathcal{R}\big(C^{(12)}\big)\subseteq\mathcal{R}(\bar{X}^{(1)}),\quad\mathcal{R}(T)\equiv\mathcal{R}\big((C^{(12)})^{\top}\big)\subseteq\mathcal{R}\big(\bar{X}^{(2)}\big).$ (24)
+To obtain interaction modes orthogonal to both the unity vectors $\mathbf{1}_{n_{c}}$ and the individual POD bases $U^{(l)}_{d_{l}}$ for $l=1,2$, we project the $W$ and $T$ onto the complementary subspace:
+$W_{\perp}=\left(I_{n_{c}}-\frac{1}{n_{c}}\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}-U_{d_{1}}^{(1)}(U_{d_{1}}^{(1)})^{\top}\right)W\in\mathbb{R}^{n_{c}\times m},\quad T_{\perp}=\left(I_{n_{c}}-\frac{1}{n_{c}}\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}-U_{d_{2}}^{(2)}(U_{d_{2}}^{(2)})^{\top}\right)T\in\mathbb{R}^{n_{c}\times m},$ (25)
+and define their orthonormalized versions as:
+$\hat{W}=W_{\perp}(W_{\perp}^{\top}W_{\perp})^{-1/2},\quad\hat{T}=T_{\perp}(T_{\perp}^{\top}T_{\perp})^{-1/2}.$ (26)
+By construction of $\hat{W}$ and $\hat{T}$, we obtain:
+$\mathbf{1}_{n_{c}}^{\top}\hat{W}=\mathbf{0}^{\top}_{m},\quad(U_{d_{1}}^{(1)})^{\top}\hat{W}=0\in\mathbb{R}^{d_{1}\times n_{c}},\quad\mathbf{1}_{n_{c}}^{\top}\hat{T}=\mathbf{0}^{\top}_{k},\quad(U_{d_{2}}^{(2)})^{\top}\hat{T}=0\in\mathbb{R}^{d_{2}\times n_{c}}.$ (27)
+The augmented orthonormal bases (Eq. (20)) are therefore:
+$\tilde{U}^{(1)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{1}}^{(1)}&\hat{W}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{1}+m+1)},\quad\tilde{U}^{(2)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{2}}^{(2)}&\hat{T}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{2}+m+1)},$ (28)
+according to which, the reduced coordinates of the snapshots can be obtained as:
+$Y^{(1)}=(\tilde{U}^{(1)})^{\top}\bar{X}^{(1)}\in\mathbb{R}^{(d_{1}+m+1)\times n_{t}},\qquad Y^{(2)}=(\tilde{U}^{(2)})^{\top}\bar{X}^{(2)}\in\mathbb{R}^{(d_{2}+m+1)\times n_{t}}.$ (29)
+Multiplying the augmented basis in Eq.
+(28) by $\mathbf{1}_{n_{c}}^{\top}$ from the left, and using Eq.
+(27) together with the orthogonality $\mathbf{1}_{n_{c}}^{\top}U^{(l)}_{d_{l}}=\mathbf{0}_{d_{l}}^{\top}$ in Eq.
+(22) gives:
+$\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(1)}=\sqrt{n_{c}}\mathbf{e}_{1}^{\top},\quad\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(2)}=\sqrt{n_{c}}\mathbf{e}_{1}^{\top},$ (30)
+where $\mathbf{e}_{1}=[1,0,\ldots,0]^{\top}$. Hence, for the projected centered data, we get:
+$\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(l)}(\tilde{U}^{(l)})^{\top}\bar{X}^{(l)}=\sqrt{n_{c}}\mathbf{e}_{1}^{\top}(\tilde{U}^{(l)})^{\top}\bar{X}^{(l)}=\mathbf{1}_{n_{c}}^{\top}\bar{X}^{(l)}=\mathbf{0}_{n_{t}}^{\top},\quad l=1,2.$ (31)
+The last equality uses the mass-preservation property for the original densities $\mathbf{1}_{n_{c}}^{\top}\bar{X}^{(l)}=\boldsymbol{0}^{\top}_{n_{t}}$ in Eq. (22).
+Finally, using the reconstruction formula in Eq. (21), and the identities derived in Eqs. (22) and (31), we obtain:
+$\mathbf{1}^{\top}_{n_{c}}\tilde{X}^{(l)}=\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(l)}(\tilde{U}^{(l)})^{\top}\bar{X}^{(l)}+\mathbf{1}^{\top}_{n_{c}}X^{(l)}(I_{n_{t}}-H)=\mathbf{1}^{\top}_{n_{c}}X^{(l)}(I_{n_{t}}-H)=\mathbf{1}_{n_{t}},\quad l=1,2.$ (32)
+Thus, the reconstructions $\tilde{X}^{(1)}$ and $\tilde{X}^{(2)}$, obtained via the projection of $X^{(1)}$ and $X^{(2)}$ onto the augmented basis in Eq.
+(28), retain the mass-preservation property of the original data.
+∎
+In summary, given a density distribution $\rho(\mathbf{x},t)$ obtained via Eq. (6) for a single population, we obtain the normalized density along the $n_{c}$ grid points of $\Omega$ as:
+$x(t)=\dfrac{1}{S(t)}\big[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)\big]^{\top}\in\mathbb{R}^{n_{c}},$ (33)
 where $S(t)$ is the total mass on time $t$, and define the restriction operator $\mathcal{R}$ by the projection on the POD basis of Eq. (11) as:
@@
 $x(t)=\mathcal{L}(\boldsymbol{y}(t))=U_{d}\boldsymbol{y}(t)+\frac{1}{n_{t}}X\mathbf{1}_{n_{t}}\in\mathbb{R}^{n_{c}}.$ (35)
-The above definitions of the restriction and lifting operators guarantee that the reconstructed, by the lifting operator, normalized density profiles preserve the total density within the domain $\Omega$, as per Proposition 1.
-We highlight here that while the operators are constructed based on the dataset $X$, their employment applies to any seen or unseen observation $x(t)\in\mathbb{R}^{n_{c}}$ obtained from Eq.
-(19) and $\boldsymbol{y}(t)\in\mathbb{R}^{d}$.
+In the case of two interacting populations with normalized densities $x^{(l)}(t)$ for $l=1,2$, we consider the augmented bases $\tilde{U}^{(l)}$ in Proposition 2 and adjust the restriction operator $\mathcal{R}$ to:
+$\boldsymbol{y}(t)=[\boldsymbol{y}^{(1)}(t),\boldsymbol{y}^{(2)}(t)]^{\top}\in\mathbb{R}^{d_{1}+d_{2}+2m+2},\quad\boldsymbol{y}^{(l)}(t)=\mathcal{R}^{(l)}(x^{(l)}(t))=(\tilde{U}^{(l)})^{\top}\left(x^{(l)}(t)-\frac{1}{n_{t}}X^{(l)}\mathbf{1}_{n_{t}}\right)\in\mathbb{R}^{d_{l}+m+1},$ (36)
+and the lifting operator $\mathcal{L}$ to:
+$x(t)=[x^{(1)}(t),x^{(2)}(t)]^{\top}\in\mathbb{R}^{2n_{c}},\quad x^{(l)}(t)=\mathcal{L}(\boldsymbol{y}^{(l)}(t))=\tilde{U}^{(l)}\boldsymbol{y}^{(l)}(t)+\frac{1}{n_{t}}X^{(l)}\mathbf{1}_{n_{t}}\in\mathbb{R}^{n_{c}}.$ (37)
+In both cases, the above definitions of the restriction and lifting operators guarantee that the reconstructed normalized density profiles preserve the total per-population density within the domain $\Omega$, as per Propositions 1 and 2.
+We highlight here that while the operators are constructed based on the dataset $X$ (or $X^{(l)}$ for $l=1,2$), their employment applies to any seen or unseen observation $x(t)$ obtained from Eq.
+(33) and $\boldsymbol{y}(t)$.
+Hereafter, for conciseness of the presentation, we will carry the single-population dimensions of $x(t)$ and $\boldsymbol{y}(t)$, unless otherwise specified.
+Remark 2.1.
+Mass preservation and element-wise positivity in the reconstructed field are distinct properties.
+Standard POD/SVD truncation does not guarantee non-negativity in reconstruction: negative pointwise values may appear even when the original snapshots are positive.
+Since the reconstructed field is obtained by adding back the original mean, small negative entries—when they occur—are expected to appear in the low-density tails of the distribution.
+In such cases, we post-process the corresponding snapshots only for observation purposes, by setting these values to zero, and renormalizing the reconstruction by the total mass of the nonnegative part.
+This ensures positivity while preserving total mass, with normalization factors remaining near 1.
+Remark 2.2.
+The reconstruction step is only observational in our pipeline, as density fields are reconstructed via $\mathcal{L}$ in Eq.
+(35) or (37), while the ROM evolves entirely in the latent space (see Eq.
+(5)).
+Since learning is performed with the original snapshots (which are nonnegative by construction), any post-process non-negativity enforcement does not influence the latent representation learning.
+Constrained reconstruction approaches, such as nearest-neighbor based methods commonly used in Diffusion Maps latent spaces [93, 26], can guarantee non-negativity by construction but require higher computational cost; integrating such techniques into our pipeline is straightforward, although a detailed comparative study is out of the scope in the current work.
 2.2.3 Learning the dynamics in latent spaces using MVAR and LSTM
@@
 (8).
-To consider lags in the latent coordinate embeddings $\boldsymbol{y}(t_{k})$, we employ both linear and nonlinear autoregressive models, namely MVARs and LSTMs.
-Let an MVAR model of order $w$ (i.e., with $w$ time lags) be expressed as:
-$\boldsymbol{y}_{c}(t_{k})=\boldsymbol{A}_{0}+\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j})+\boldsymbol{\varepsilon}(t_{k}),$ (22)
-where $\boldsymbol{y}_{c}(t_{k})$ is the latent variable vector, evaluated at discrete times $t_{k}$ for each of the $c=1,\ldots,C$ cases of different initial condition considered, $\boldsymbol{A}_{0}\in\mathbb{R}^{d}$ is the intercept vector and $\boldsymbol{A}_{j}\in\mathbb{R}^{d\times d}$ are the regression coefficient matrices for each lag $j$.
-The predictor vectors $\boldsymbol{y}_{c}(t_{k-j})\in\mathbb{R}^{d}$ include the $w$ past lag latent variables evaluated at discrete times $t_{k-1},\ldots,t_{k-w}$, while the error term $\boldsymbol{\varepsilon}(t_{k})\in\mathbb{R}^{d}$ is assumed to follow a multivariate white noise process:
+To capture temporal dependencies via delayed embeddings (in line with Takens’/Whitney’s embedding theorems [106, 99]), we employ autoregressive ROMs to learn the dynamics of the latent embeddings $\boldsymbol{y}(t_{k})$.
+The simplest such model is MVAR, which provides an efficient, linear baseline with globally optimal, least squares solutions.
+For a nonlinear comparison, we employ LSTM networks [64], whose gating mechanisms effectively model complex temporal patterns; though numerous alternatives exist (Transformers [60], Neural Ordinary Differential Equations (ODEs) [98], etc.), LSTMs represent state-of-the-art in nonlinear sequential modeling.
+A single ROM is trained in both single- and two-population cases; in the latter, this supports coupled latent dynamics that capture temporal interactions between groups.
+For compactness in what follows, we denote the dimension of the concatenated two-population latent coordinates in Eq.
+(36) by $d=d_{1}+d_{2}+2(m+1)$, to match the single-population dimensions.
+Let an MVAR model of order $w$ (i.e., with $w$ time lags) without intercept be expressed as:
+$\boldsymbol{y}_{c}(t_{k})=\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j})+\boldsymbol{\varepsilon}(t_{k}),$ (38)
+where $\boldsymbol{y}_{c}(t_{k})$ is the latent variable vector, evaluated at discrete times $t_{k}$ for each of the $c=1,\ldots,C$ cases of different initial condition considered, and $\boldsymbol{A}_{j}\in\mathbb{R}^{d\times d}$ are the regression coefficient matrices for each lag $j$.
+The predictor vectors $\boldsymbol{y}_{c}(t_{k-j})\in\mathbb{R}^{d}$ include the $w$ past lag latent variables evaluated at discrete times $t_{k-1},\ldots,t_{k-w}$.
+Since the latent variables $\boldsymbol{y}_{c}(t_{k})$ are zero-mean (a consequence of the mean-centered POD projection) the intercept term is excluded from Eq.
+(38).
+The error term $\boldsymbol{\varepsilon}(t_{k})\in\mathbb{R}^{d}$ is assumed to follow a multivariate white noise process:
 $\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})]=\mathbf{0},\qquad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{k})^{\top}]=\Sigma\quad\text{and}\quad\mathbb{E}[\boldsymbol{\varepsilon}(t_{k})\boldsymbol{\varepsilon}(t_{r})^{\top}]=\mathbf{0}\quad\forall k\neq r,$ (39)
@@
 Given sequential (time series) data for $\{\boldsymbol{y}_{c}(t_{k})\}$, the unknown parameters of the MVAR in Eq.
-(22), i.e., the intercept vector $\boldsymbol{A}_{0}$ and the coefficient matrices $\boldsymbol{A}_{j}$, can be computed by the solution of the following least-squares problem:
-$\operatorname*{argmin}_{\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}}\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right),$ (24)
+(38), i.e., and the coefficient matrices $\boldsymbol{A}_{j}$, can be computed by the solution of the following least-squares problem:
+$\operatorname*{argmin}_{\{\boldsymbol{A}_{j}\}_{j=1}^{w}}\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right),$ (40)
 where the MVAR loss function is:
-$\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\sum_{c=1}^{C}\sum_{k=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert_{2}^{2},$ (25)
+$\mathcal{L}_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\sum_{c=1}^{C}\sum_{k=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert_{2}^{2},$ (41)
 with the predictions of the MVAR model given by:
-$\hat{\boldsymbol{y}}_{c}(t_{k})=\boldsymbol{A}_{0}+\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j}),$ (26)
+$\hat{\boldsymbol{y}}_{c}(t_{k})=\sum_{j=1}^{w}\boldsymbol{A}_{j}\boldsymbol{y}_{c}(t_{k-j}),$ (42)
 for $c=1,\ldots,C$ and $k=1,\ldots,K$.
 The optimization problem in Eq.
-(24) can be solved using methods such as regularized least squares; in particular, we used ridge regularization [86].
+(40) can be solved using regularized least squares methods [65].
 We note that MVARs require the trivial assumption of time series stationarity, which was the case for our illustrations.
@@
 We highlight that MVAR model training further requires the selection of an optimal lag window size $w$, which determines the length of the time series history used to model the temporal dependencies.
-Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in Appendix D.
-As a non-linear autoregressive model alternative to linear MVARs, we further consider LSTM networks, a type of Recurrent Neural Networks (RNNs) characterized by gating mechanisms, which allow the network to retain and propagate information over long sequences selectively, but also effectively handle vanishing gradients; a common challenge in traditional RNNs [87, 88, 89, 90].
+Here, we used two information-theoretic criteria for determining $w$, the details for which are provided in D.
+As a non-linear autoregressive model alternative to linear MVARs, we further consider LSTM networks [64], a type of Recurrent Neural Networks (RNNs) characterized by gating mechanisms, which allow the network to retain and propagate information over long sequences selectively, but also effectively handle vanishing gradients; a common challenge in traditional RNNs [56, 25, 113].
 Let an LSTM model with past $w$ latent states $\boldsymbol{y}_{c}(t_{k-j})$ for $j=1,\ldots,w$ be expressed as:
@@
 The gating mechanism of the LSTM is included in Eq.
-(27) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for conciseness of the presentation; a vanilla LSTM including 4 gates, and contains the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
+(43) at $F_{\mathrm{LSTM}}(\cdot)$, which we avoid here for conciseness of the presentation; a detailed presentation of the gating mechanism in $F_{\mathrm{LSTM}}(\cdot)$ and the related parameters $\boldsymbol{\theta}$ can be found in [64, 56].
+Here, we use a vanilla LSTM with 4 gates, containing the parameters $\boldsymbol{\theta}\in\mathbb{R}^{4N_{h}(d+N_{h}+1)}$.
 Similarly to the sequence-to-one learning of MVARs, given sequential data for $\{\boldsymbol{y}_{c}(t_{k})\}$ for a fixed lag window of width $w$, we train the LSTM model by solving the optimization problem:
@@
 where the LSTM loss function is the mean squared error (MSE) between the true value of the latent variable vector $\boldsymbol{y}_{c}(t_{k})$ and the predicted one $\hat{\boldsymbol{y}}_{c}(t_{k})$, as provided by Eqs.
-(27, 28), that is:
-$\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert^{2}.$ (30)
+(43, 44), that is:
+$\mathcal{L}_{\text{LSTM}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{\theta},\mathbf{W}_{y},\mathbf{b}_{y}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert^{2}_{2}.$ (46)
 To solve the optimization problem in Eq.
-(30), one typically uses gradient-based methods; in particular, we employed the Adam optimizer [91], which is a stochastic variant of gradient descent, known for its adaptive learning-rate capabilities.
+(46), one typically uses gradient-based methods.
+Here, we employ the Adam optimizer [74], which is a stochastic variant of gradient descent, known for its adaptive learning-rate capabilities.
 We highlight that while MVAR models are theoretically simple and efficient, their strict reliance on stationary data poses a significant limitation, requiring explicit pre‑processing that may distort dynamics or lose information [23].
@@
 However, this flexibility comes at the cost of needing much larger datasets, greater computational resources, and often lacking interpretability compared to MVAR models.
-Finally, the predicted latent coordinates $\hat{\boldsymbol{y}}_{c}(t_{k})$ obtained from the trained MVARs or LSTMs models are subsequently mapped back to the high-dimensional density space using the lifting operator defined in Eq.
-(21).
-3 Case study and methodology implementation
-3.1 Configuration
-The proposed methodology is employed on a crowd dynamics configuration where pedestrians are moving from left to right in a corridor of dimensions $48\,\text{m}\times 12\,\text{m}$ ($x$: length, $y$: width) in the presence of an obstacle, as shown in Fig.
+Finally, the predicted latent coordinates $\hat{\boldsymbol{y}}_{c}(t_{k})$ obtained from the trained MVARs or LSTMs models are subsequently mapped back to the high-dimensional density space using the lifting operator defined in Eqs.
+(35) or (37) for the single or two-population case, respectively.
+3 Case studies and methodology implementation
+We considered two representative benchmark problems in crowd dynamics: a unidirectional crowd flow moving through a corridor containing an obstacle, and a counterflow of two interacting populations in the same domain.
+The configurations of both case studies are described below, along with the implementation of the proposed methodology.
+3.1 Configurations
+The two crowd dynamics configurations simulate pedestrians moving in a corridor of dimensions $48\,\text{m}\times 12\,\text{m}$ ($x$: length, $y$: width) in the presence of an obstacle, as shown in Fig.
 2.
-Figure 2: The case study: The SFM simulating pedestrians moving in a corridor from the left to the right with an obstacle.
-The square obstacle of area $12.96\,\text{m}^{2}$ is centered at $(25.8\,\text{m},1.8\,\text{m})$, with the pedestrians navigating around it towards reaching the corridor’s right end.
-Hence, the computational domain is $\Omega=\left([0,48]\times[0,12]\right)\backslash\left([24,27.6]\times[0,3.6]\right)$.
-For modelling the pedestrians’ movement, we use the SFM [1] as a microscopic agent-based simulator.
+(a) Unidirectional flow
+(b) Counterflow
+Figure 2: Crowd dynamics configurations of SFM simulating pedestrians moving in a corridor past an obstacle.
+In the unidirectional flow configuration in Fig.
+2(a), pedestrians move from the left to the right of the corridor.
+In the counterflow configuration in Fig.
+2(b), two pedestrian populations move in opposite directions; group 1 (blue) moves from the left to the right, while group 2 (red) follows the opposite direction.
+The square obstacle of area $12.96\,\text{m}^{2}$ is centered at $(25.8\,\text{m},1.8\,\text{m})$, with the pedestrians navigating around it towards reaching their destination (right or left corridor’s boundary).
+Hence, the computational domain is $\Omega\backslash\Gamma$, with the corridor occupying $\Omega=[0,48]\times[0,12]$ and the obstacle $\Gamma=[24,27.6]\times[0,3.6]$.
+For modelling the pedestrians’ movement, we used the SFM [62, 61] as a microscopic agent-based simulator.
 SFM falls within the general form of Eqs.
@@
 Details on the SFM forces included in Eq.
-(31) and the related parameters that were considered in this work are provided in Appendix A.
-We simulate a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see in the Appendix B).
-Pedestrians are directed to move from left to right by setting their initial target at coordinates $(x,y)=(25\,\text{m},\,9.6\,\text{m})$, corresponding to the midpoint of the corridor in the horizontal direction and 80% of its vertical extent.
-Upon reaching this intermediate waypoint, their final destination is updated to $(x,y)=(48\,\text{m},\,4.2\,\text{m})$, which lies at the right end of the corridor and at 35% of its height.
-To maintain a constant population density, we implement periodic boundary conditions by re-spawning agents at the left boundary at $x=0\,\text{m}$, when agents reach the right boundary at $x=48\,\text{m}$.
-Pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator.
-Such a configuration ensures a controlled pedestrian flow through the simulated environment and also ensures the pedestrians avoid the obstacle each time they reenter the corridor.
+(47) and the related parameters that were considered in this work are provided in A.
+Remark 3.1.
+Here, we employed the Social Force model purely as a convenient baseline to demonstrate and validate the proposed methodology.
+Naturally, the same workflow is not tied to this specific choice: it can be coupled just as well with more sophisticated agent-based approaches—e.g., models that include explicit route planning, richer interaction rules, heterogeneous behaviors, or learned decision-making—depending on the fidelity required by the application.
+In the unidirectional flow configuration (see Fig.
+2(a), we simulated a crowd of $N=100$ pedestrians moving through a corridor, with initial spatial distributions ranging from homogeneous dispersions to clustered configurations (for more details, see C).
+Pedestrians are directed to move from the left to the right of the corridor.
+Since the SFM governs pedestrians’ motion through target-directed forces without intrinsic path planning, we employ a sequential two-target scheme to induce realistic obstacle-avoidance behavior.
+Pedestrians are initially guided towards an intermediate target zone below the obstacle at coordinates $x=25\,\text{m}$ and $y\in[6,7.8]\,\text{m}$.
+Upon reaching this zone, their final destination is updated to $x=48\,\text{m}$ with $y$-coordinates remaining at the values each agent had upon reaching the intermediate zone.
+To maintain a constant population density, we implemented periodic boundary conditions by re-spawning agents at the left boundary at $x=0\,\text{m}$, when agents reach the right boundary at $x=48\,\text{m}$.
+In the counterflow configuration (see Fig.
+2(b), we simulated a crowd of $N=100$ pedestrians divided equally into two groups with opposing flow directions: group 1 moves from left to right, and group 2 from right to left.
+We used the SFM in Eq.
+(47) with the same parameters as in the unidirectional case, and initial spatial distributions that are symmetric for each group along $x=24\,\text{m}$; group 1 follows the same initial distribution as in the unidirectional case.
+Employing a two-target scheme to induce realistic obstacle-avoidance behavior, we prescribed distinct preferred zones for each group.
+Group 1 is initially guided to the zone $x=25\,\text{m}$ and $y\in[5.4,7.8]\,\text{m}$, while group 2 to $x=25\,\text{m}$ and $y\in[9.6,11.4]\,\text{m}$.
+Upon reaching these zones, their final destination is updated to $x=48\,\mathrm{m}$ for group 1 and $x=0\,\mathrm{m}$ for group 2, while the $y$-coordinates remain fixed at the values each agent had upon reaching the intermediate zone.
+This separation promotes smooth obstacle avoidance and organized counterflow without introducing explicit route planning.
+As in the unidirectional case, periodic boundary conditions are applied at both ends of the corridor by re-spawning agents at the left/right boundary upon exiting at the right/left for groups 1/2, respectively.
+In both configurations pedestrians are prohibited from passing the corridor’s walls or entering the obstacle area by introducing strong wall forces in the SFM microscopic simulator.
+Such a configuration ensures a controlled pedestrian flow through the simulated environment and a realistic obstacle-avoidance behavior.
 3.2 Data generation of microscopic distributions
-We generate synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy.
+We generated synthetic data of pedestrian trajectories by numerically integrating the SFM with a first-order explicit Euler scheme using a small fixed time step of $\delta t_{micro}=0.025\,\text{s}$ to ensure stability and accuracy.
 Each simulation has a time span of $250\,\text{s}$.
@@
 As already discussed, we generate pedestrian trajectories from various initial spatial distributions modelling diverse crowd scenarios.
-In particular, we consider $C=20$ different initializations with zero velocities and positions distributed via: (i) uniform distributions to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) “double bell-shaped” Gaussian distributions to capture the formation of two distinct crowd groups and (iv) sinusoidal distributions to generate periodic macroscopic behavior.
-Details on the form and parameters of the above initializations are provided in Appendix B, distinguishing between those used for training and testing the proposed framework in Tables B.1 and B.2, respectively.
-The resulting datasets are comprised of 20 cases with different initializations, each case containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1000}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1000$ time steps correspond to the samples recorded from each trajectory.
+In particular, we consider $C=20$ different initializations with zero velocities and positions distributed via: (i) uniform distributions to provide a homogeneous baseline scenario, (ii) Gaussian clusters to model localized groups of pedestrians, (iii) “double bell-shaped” Gaussian distributions to capture the formation of two distinct crowd groups and (iv) cosinusoidal distributions to generate periodic macroscopic behavior.
+Details on the form and parameters of the above initializations are provided in C, distinguishing between those used for training and testing the proposed framework in Tables C.1 and C.2, respectively.
+For the unidirectional flow, the resulting datasets are comprised of 20 cases with different initializations, each containing a matrix $\mathcal{X}^{(c)}=\{\bm{x}_{i}(t_{k}),i=1,\ldots,N,k=1,\ldots,K\}\in\mathbb{R}^{200\times 1000}$, where $2N=200$ corresponds to the pedestrians’ positions and $K=1000$ time steps correspond to the samples recorded from each trajectory.
 Thus, the total datasets over all cases are $\mathcal{X}_{tr},\mathcal{X}_{ts}\in\mathbb{R}^{200\times(20\cdot 1000)}$ used for training and testing, respectively.
+For the counterflow, we generated data from the same set of $C=20$ initializations; see C.
+Here, the initialization for group 1 is identical to that of the unidirectional case (with half the agents), while group 2 is initialized symmetrically to group 1 along $x=24\,\text{m}$, preserving the $y$-positions.
+Each initialization case now produces two separate data matrices $\mathcal{X}^{(1,c)}$ for group 1 and $\mathcal{X}^{(2,c)}$ for group 2, both of dimension $\mathbb{R}^{100\times 1000}$.
+Consequently, the complete training and testing datasets for the counterflow are $\{\mathcal{X}_{tr}^{(1)},\mathcal{X}_{tr}^{(2)}\}$ and $\{\mathcal{X}_{ts}^{(1)},\mathcal{X}_{ts}^{(2)}\}$, respectively.
 3.3 Extraction of continuous density fields
 For the first step of our framework, we derive macroscopic density fields from pedestrian positions via KDE, as described in Section 2.2.1.
-To achieve this, we discretize the computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behavior, while minimizing artificial grid effects.
+To achieve this, we discretize the corridor computational domain $\Omega$ into $n_{x}\times n_{y}=80\times 20$ control volumes with uniform spacing $\delta x=\delta y=0.6\,\mathrm{m}$; this resolution reflects a typical personal space for each pedestrian, allowing for capturing emergent collective behavior, while minimizing artificial grid effects.
 For the KDE implementation on the cell centroids, we choose Gaussian kernels in Eq.
 (6) tuned with bandwidths $\mathbf{H}=diag(h_{x},h_{y})=(3,2)$ for suppressing noise and preserving features of the microscopic data, while ensuring that local structures remain clear without introducing spurious artifacts.
-To account for the physical obstacle in $\Omega$, a binary mask is applied to explicitly zero out density estimation on the cell centroids of the obstacle geometry.
+To account for the physical obstacle in $\Gamma\subset\Omega$, we applied a binary mask to the KDE-estimated density field, enforcing zero density in $\Gamma$.
+This is necessary because KDE can produce non-zero density estimates near obstacle boundaries when pedestrians are in close proximity.
 To account for the periodic boundary conditions, we further perform domain augmentation for the KDE support.
-All the details regarding KDE implementation are provided in Appendix C.
-Using the above procedure, we derive the density fields $\rho^{(c)}\big(x_{i},y_{j},t_{k}\big)$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing $\mathcal{X}_{ts}$ data sets.
-For conforming to the assumptions of Proposition 1, we finally normalize each density field following Eq.
+All the details regarding KDE implementation are provided in B.
+Using the above procedure, in the unidirectional flow case, we derived the density fields $\rho^{(c)}\big(x_{i},y_{j},t_{k}\big)$ ($c=1,\ldots,C$, $i=1,\ldots,n_{x}$, $j=1,\ldots,n_{y}$ and $k=1,\ldots,K$) from the pedestrian positions included in the training $\mathcal{X}_{tr}$ and testing $\mathcal{X}_{ts}$ data sets.
+For conforming to the assumptions of Proposition 1, we normalize each density field following Eq.
 (33), thus constructing the normalized density matrices $X_{tr},X_{ts}\in\mathbb{R}^{n_{c}\times n_{t}}$ in the form of Eq.
 (9), where $n_{c}=n_{x}\times n_{y}=1600$ and $n_{t}=C\cdot K=20\cdot 1000$.
+In the counterflow case, we followed the same procedure to extract and normalize the density fields of each group separately, as it is required by Proposition 2, yielding the matrices $X_{tr}^{(1)},X_{ts}^{(1)}\in\mathbb{R}^{n_{c}\times n_{t}}$ for group 1 and $X_{tr}^{(2)},X_{ts}^{(2)}\in\mathbb{R}^{n_{c}\times n_{t}}$ for group 2.
 3.4 POD for the restriction and lifting operators
-For constructing the restriction and lifting operators in Eqs.
-(20) and (21) respectively, we employ POD on the training dataset $X_{tr}$ to determine the latent dimension $d$ and the POD basis $U_{d}$.
-Using the economy-size SVD (which reduces computational complexity in thin matrices while maintaining exact reconstruction capability), we determine $d$ by the best low-rank approximation of $X_{tr}$ retaining $99\%$ of the energy; thus, computing the minimum number $d$ satisfying the criterion:
+We construct the restriction and lifting operators by employing POD on the training data, determining the latent dimension $d$ and the POD basis $U_{d}$ via the economy-size SVD.
+In all cases, we determine $d$ by the best low-rank approximation retaining $99\%$ of the energy; thus, computing the minimum number satisfying the criterion:
 $E_{d}=\dfrac{\sum_{i=1}^{d}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq 0.99$ (48)
-where $r=\min(n_{c},n_{t})=1600$ is the rank of $X_{tr}$, and $\sigma_{i}>0$ for $i=1,\ldots,r$ are the singular values in descending order.
-Next, using the first $d$ left singular vectors of $X_{tr}$, we form $U_{d}$ for completing the construction of the restriction and lifting operators in Eqs.
-(20, 21).
-To assess the accuracy of restriction and lifting operators, we evaluate the POD reconstruction error on the training and testing data sets by the relative $L_{2}$ error:
+where $r=\min(n_{c},n_{t})=1600$ is the rank of the data matrix, and $\sigma_{i}>0$ for $i=1,\ldots,r$ are the singular values in descending order.
+The basis $U_{d}$ is then formed by the first $d$ left singular vectors.
+For the unidirectional flow case, we applied this procedure to the training dataset $X_{tr}$ and the resulting basis $U_{d}$ is then used to construct the restriction and lifting operators in Eqs.
+(34) and (35), respectively.
+For the counterflow case, the above procedure is done per group, to compute the separate POD bases $U_{d_{1}}^{(1)}$ and $U_{d_{2}}^{(2)}$ from the group-specific training datasets $X_{tr}^{(1)}$ and $X_{tr}^{(2)}$, respectively.
+Following Proposition 2, we then computed the SVD of the cross-covariance matrix $C^{(12)}=\frac{1}{n_{t}}\bar{X}^{(1)}_{tr}\left(\bar{X}_{tr}^{(2)}\right)^{\top}$ of the training sets, which is required to form the augmented bases $\tilde{U}_{d_{l}}^{(l)}$ (for $l=1,2$) defined in Eq.
+(20).
+These augmented bases were subsequently used to construct the group-specific restriction and lifting operators in Eqs.
+(36) and (37).
+For both configurations, to assess the accuracy of restriction and lifting operators, we evaluate the POD reconstruction error on the training and testing data sets by the relative $L_{2}$ error:
 $e^{2,rec}_{k}=\dfrac{\lVert x(t_{k})-\mathcal{L}(\mathcal{R}(x(t_{k})))\rVert_{2}}{\lVert x(t_{k})\rVert_{2}},$ (49)
-where $x(t_{k})\in X_{tr},X_{ts}$ with $k=1,\ldots,K$ is the normalized density defined in Eq. (19).
+where $x(t_{k})\in X_{tr},X_{ts}$ denotes the normalized density in Eq.
+(33) at time $t_{k}$ for $k=1,\ldots,K$.
+For the counterflow, this error is computed separately for each group using the corresponding operators and data sets $X^{(l)}_{tr},X^{(l)}_{ts}$ for $l=1,2$.
 3.5 ROMs for latent space dynamics and their training
 Having obtained the latent representations $\boldsymbol{y}(t)\in\mathbb{R}^{d}$ from the restriction operator, the third step of the proposed approach is to construct ROMs for learning the dynamics in the latent space.
-For this purpose, we learn linear MVAR and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}(t_{k-1})\}$.
-To conform to this sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq.
-(20) on the training data $X_{tr}$.
-As a preliminary step, we first tune the lag window $w$, within a bound of $w_{\text{max}}=100$, using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [93, 94, 95]; details are provided in Appendix D.
+For this purpose, we learn linear MVAR and non-linear LSTM models, as described in Section 2.2.3, for predicting the latent variable $\hat{\boldsymbol{y}}_{c}(t_{k})$ of the $c$ case at timestep $t_{k}$ given the past $w$ lags $\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}$.Note that in the counterflow, a single ROM is trained on the concatenated latent vectors from both groups $\boldsymbol{y}_{c}(t_{k})=[\boldsymbol{y}^{(1)}_{c}(t_{k}),\boldsymbol{y}^{(2)}_{c}(t_{k})]^{\top}$, as detailed in Section 2.2.3.
+To conform to sequence-to-one learning, the features consist of the dataset $Y_{f}=\{\boldsymbol{y}_{c}(t_{k-w}),\ldots,\boldsymbol{y}_{c}(t_{k-1})\}\in\mathbb{R}^{d\times w\times C\cdot(K-w)}$ and the targets consist of the data set $Y_{t}=\{\boldsymbol{y}_{c}(t_{k})\}\in\mathbb{R}^{d\times C\cdot(K-w)}$, where each $\boldsymbol{y}_{c}(t_{k})$ is computed via the restriction operator in Eq.
+(34) on the training data $X_{tr}$ (for the unidirectional case) and via Eq.
+(36) on $X_{tr}^{(1)},X_{tr}^{(2)}$ (for the counterflow case).
+As a preliminary step, we first tune the lag window $w$, within a bound of $w_{\text{max}}=20$, using two information-theoretic criteria: the Bayesian Information Criterion (BIC) and the Akaike Information Criterion (AIC) [81, 2, 100]; details are provided in D.
 Both criteria are likelihood-based, with AIC tending to select larger windows than BIC.
-We note that BIC and AIC are well-suited for MVARs models since they assume stationarity of the data and white noise residuals.
+We note that BIC and AIC are well-suited for MVARs since they assume stationarity of the data and white noise residuals.
 For comparison, we adopt the same $w$ for learning LSTMs, as the one tuned for MVARs by these criteria.
-For the MVARs model training, we first verify that all input time series are stationary using the augmented Dickey–Fuller (ADF) test [96] with a threshold of significance level $p<0.01$.
-Then, we compute the coefficients $\boldsymbol{A}_{0}$ and $\boldsymbol{A}_{j}$ in Eq.
-(22) using ordinary least-squares (OLS) to solve the optimization problem in Eq.
+For the MVAR model training in both configurations, we first verified that all input time series are stationary using the augmented Dickey–Fuller (ADF) test [39] with a threshold of significance level $p<0.01$.
+Then, we computed the coefficients $\boldsymbol{A}_{j}$ in Eq.
+(38) using ordinary least squares to solve the optimization problem in Eq.
 (40).
-For the LSTM model, we consider a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq.
+For the counterflow case, where the related data-matrix was ill-conditioned, we employed ridge regularization with $\lambda=10^{-6}$.
+For the LSTM model in both configurations, we considered a single LSTM layer with $16$ units and a hyperbolic tangent activation function in the hidden layers; the output layer is linear, as denoted in Eq.
 (44).
@@
 The model’s parameters are initialized with a Glorot uniform initialization [52], the initial learning rate is set to $l_{r}=10^{-3}$, and the mini-batch size is set to $32$; the latter hyperparameters were tuned on a trial-and-error basis, providing a good compromise between computational efficiency and generalization without exceeding memory constraints.
-The maximum number of epochs was set to $200$ as no further improvement was achieved using more epochs.
-To further avoid overfitting of the LSTM, we split the training data into training and validation sets using a 80-20% ratio, and shuffle each data set without breaking the internal temporal structure of the sequences.
-We compute the loss function value on the validation set for every epoch and stop the training after 5 consecutive validation checks with no improvement.
+The maximum number of training epochs was set to $40$ for the unidirectional flow case and to $70$ for the counterflow case, based on the saturation of the training loss, which yielded no additional improvement beyond these epochs.
 To assess the convergence of both MVARs and LSTMs training, we report the values of the loss functions in Eqs.
@@
 Since the MVARs loss function is based on $L_{2}$ error, we further report the MSE at the latent space:
-$MSE_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\boldsymbol{A}_{0},\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\rVert^{2},$ (34)
+$MSE_{\text{MVAR}}\left(\{\boldsymbol{y}_{c}(t_{k})\},\{\hat{\boldsymbol{y}}_{c}(t_{k})\};\{\boldsymbol{A}_{j}\}_{j=1}^{w}\right)=\dfrac{1}{C(K-w)}\sum_{c=1}^{C}\sum_{t_{k}=w+1}^{K}\Big\lVert\boldsymbol{y}_{c}(t_{k})-\hat{\boldsymbol{y}}_{c}(t_{k})\Big\rVert^{2}_{2},$ (50)
 where $\hat{\boldsymbol{y}}_{c}(t_{k})$ is the prediction in Eq.
 (42), to enable direct comparison with the MSE-based loss function of LSTMs.
-Additionally, we report the computational times required for training both ROMs (for LSTMs, we report the mean error and 10–90% error percentiles statistics across the 50 runs); we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the OLS method employed for solving the related optimization problem.
+Additionally, we report the computational times required for training both ROMs (for LSTMs, we report the mean error and 10–90% error percentiles statistics across the 50 runs); we expect MVARs to be much faster than LSTMs due to the lower number of parameters and the method employed for solving the related optimization problem.
 3.6 Accuracy assessment of the proposed framework
-Having constructed/trained each step of the proposed framework–construction of the restriction/lifting operators and training of the ROMs–, we finally assess the forecasting performance of the whole framework on the $C=20$ different initializations of the testing set $\mathcal{X}_{ts}$; see Appendix B for details.
+Having constructed/trained each step of the proposed framework (construction of the restriction/lifting operators and training of the ROMs), we finally assess the forecasting performance of the whole framework on the $C=20$ different initializations of the testing set $\mathcal{X}_{ts}$; see C for details.
 In particular, we first obtain the “ground truth” normalized densities for each initialization, say $x^{(c)}(t_{0}+k\ \delta t)\equiv x^{(c)}(t_{k})\in X_{ts}$ for all time steps $k=1,\ldots,K$, using Eq.
 (33).
-Then, following the proposed restrict-evolve with ROM-lift framework in Eq.
-(5), we forecast the normalized density at the $k$ time step, using the $w$ past normalized densities in both an open-loop and closed-loop setting.
-For the open-loop setting evaluation, we compute the prediction:
-$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(x^{(c)}(t_{k-1})),\ldots,\mathcal{R}(x^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (35)
-while for the closed-loop setting evaluation, we compute the prediction:
-$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\Phi_{\text{ROM}}\left(\mathcal{R}(\hat{x}^{(c)}(t_{k-1})),\ldots,\mathcal{R}(\hat{x}^{(c)}(t_{k-w}));\mathbf{p}\right)\right),$ (36)
-where $\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators in Eqs.
-(20) and (21) and $\Phi_{\text{ROM}}$ is the latent dynamics ROM with parameters $\mathbf{p}$ that can be either the trained MVAR or LSTM model.
-In particular, for the LSTM model, we consider the parameter set resulting in the best training performance among the 50 runs.
-Notice that the open-loop prediction in Eq.
-(35) considers the restricted values of the observed normalized densities $\mathcal{R}(x^{(c)}(t_{k-j}))$ for $j=1,\ldots,w$ as inputs, while the closed-loop prediction in Eq.
-(36) instead considers the predicted–from previous steps–normalized densities $\mathcal{R}(\hat{x}^{(c)}(t_{k-j}))$ in a recursive fashion.
-Since the open-loop setting requires the “ground truth” densities, we use it here to evaluate the one-step predictions of our framework.
-For the desired, in practice, longer predictions, we utilize the closed-loop setting, which enables us to recursively feed subsequent predictions into Eq.
-(36), resulting in an autoregressive time-stepping framework also considering the effect of one-step forecast error accumulation.
-To quantify the end-to-end prediction accuracy, we compute the relative, reconstructed in the ambient-space, $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ (either for the open-loop or the closed-loop setting) per case $c$ in the testing set and time step $k$ as:
+While ROMs are trained in an open-loop setting using ground truth embeddings, we evaluate the proposed framework in a closed-loop setting (i.e., recursively feeding subsequent predictions to the autoregressive ROMs), as this reflects the intended use in practice.
+In particular, following the restrict-evolve with ROM-lift framework in Eq.
+(5), we compute the closed-loop prediction:
+$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\hat{\boldsymbol{y}}^{(c)}(t_{k})\right),\qquad\hat{\boldsymbol{y}}^{(c)}(t_{k})=\Phi_{\text{ROM}}\left(\hat{\boldsymbol{y}}^{(c)}(t_{k-1}),\ldots,\hat{\boldsymbol{y}}^{(c)}(t_{k-w});\mathbf{p}\right),$ (51)
+where $\hat{\boldsymbol{y}}^{(c)}(t_{k-j})$ are latent points predicted recursively by $\Phi_{\text{ROM}}$ in the previous $j=1,\ldots,w$ steps.
+Closed-loop forecasting requires the initial $j=1,\ldots,w$ embeddings of the observed densities $\mathcal{R}(x^{(c)}(t_{j}))$.
+Here,
+$\mathcal{R}$ and $\mathcal{L}$ are the restriction and lifting operators (defined in Eqs.
+(34,35) for the unidirectional flow or in Eqs.
+(36, 37) for the counterflow), and $\Phi_{\text{ROM}}$ is the latent dynamics ROM with parameters $\mathbf{p}$ that can be either the trained MVAR or LSTM model; for LSTMs, we select the best-performing model among the 50 training runs.
+For completeness, we provide the open-loop setting and corresponding one-step prediction results in E and F for the unidirectional flow and counterflow cases, respectively.
+In open-loop, the latent points are obtained from ground truth embeddings $\boldsymbol{y}^{(c)}(t_{k-j})=\mathcal{R}(x^{(c)}(t_{k-j}))$ rather than previous predictions $\hat{\boldsymbol{y}}^{(c)}(t_{k-j})$ in Eq.
+(51).
+For evaluation purposes, the prediction is obtained at every time step $k$.
+However, in practice, the lifting is applied on-demand only when a high-dimensional reconstruction is required.
+This renders the scheme particularly efficient, as the evolution is governed by the ROM in the latent space.
+To quantify the end-to-end prediction accuracy, we compute the relative, reconstructed in the ambient-space, $L_{1}$, $L_{2}$ and $L_{\infty}$ errors between the “ground truth” densities $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ per case $c$ in the testing set and time step $k$ as:
 $e^{1,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{1}}{\lVert x^{(c)}(t_{k})\rVert_{1}},\quad e^{2,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{2}}{\lVert x^{(c)}(t_{k})\rVert_{2}},\quad e^{\infty,(c)}_{k}=\dfrac{\lVert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rVert_{\infty}}{\lVert x^{(c)}(t_{k})\rVert_{\infty}}.$ (52)
@@
 To track down the prediction accuracy in time, we further depict the mean and 10–90% percentiles of the error evolution in time over all cases.
+For the counterflow, these errors are computed separately for each group $l=1,2$.
 All numerical computations were executed on a PC equipped with an 11th Gen Intel(R) Core(TM) i7-1165G7 2.80 GHz CPU and 16 GB of system memory.
+The source code is publicly available at github.com/patsatzisdim/NGEF_CD.
 4 Numerical results
-Here, we present the results of the proposed framework on predicting pedestrian flow following the configuration of a rectangular corridor with an obstacle described in Section 3.1.
+Here, we present the results of the proposed framework on predicting pedestrian flow in a rectangular corridor with an obstacle for the unidirectional flow and counterflow configurations described in Section 3.1.
+We end the section with a comparative discussion about the use of MVAR and LSTM models.
+4.1 Unidirectional flow case study
+We first considered the unidirectional flow of pedestrians moving from the left to the right of the corridor as shown in Fig.
+2(a).
 As already discussed in Section 3.2, we first collected pedestrian trajectory data using the SFM microscopic agent-based simulator under a wide range of initial conditions.
@@
 Following the criterion in Eq.
-(32), the minimum latent dimension required for retaining more than 99% of the energy in the training dataset is $d=13$ modes, as shown in Fig.
+(48), the minimum latent dimension required for retaining more than 99% of the energy in the training dataset is $d=6$ modes, as shown in Fig.
 3(a).
@@
 Figure 3: Accuracy of the restriction and lifting operators in Eqs.
-(20) and (21).
+(34) and (35) for the unidirectional flow case.
 Panel (a) shows the minimum number of POD modes (colored diamonds) required for retaining the desired percentage of energy in the training data set.
 Panel (b) depicts the average (over the $C=20$ cases with different initial conditions) reconstruction error $e^{2,rec}_{k}$ in Eq.
-(33) with $d=13$ POD modes, in time for the training and testing datasets.
+(49) with $d=6$ POD modes, in time for the training and testing datasets.
 Using the discovered POD basis $U_{d}$, we next constructed the restriction and lifting operators in Eqs.
@@
 3(b) displays the average relative error over all the $C=20$ cases with different initial conditions included in either the training or testing datasets.
-While initially the reconstruction error is $\sim$15%, it decreases to 10% at the first 30s30\ $\text${s} and further reduces to $\sim$7% after 50s50\ $\text${s} for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=13$.
+While initially the relative reconstruction error is $\sim$21%, it decreases to 4% at the first 20s20\ $\text${s} and further reduces to $\sim$3% after 50s50\ $\text${s} for both seen (training set) and unseen (testing set) data, indicating high accuracy of the restriction and lifting operators with a significantly reduced latent space dimension $d=6$.
 We note that this level of reconstruction error serves as an infimum approximation error baseline, i.e., the best possible error for the proposed framework at the ambient–density profiles–space.
 ROM forecasts in the long term are therefore expected to surpass this baseline error due to the additional error contributions from model-form error and training/generalization errors.
-We next train MVAR and LSTM models to learn ROMs for the latent dynamics (see Section 2.2.3), where the latent variables $\boldsymbol{y}(t)\in\mathbb{R}^{13}$ are obtained from the density profiles of the training set via the restriction operator.
-To determine the lag window, we employed the BIC and AIC criteria described in Appendix D; BIC indicated a lag window of $w=4$, whereas AIC favored a wider (as expected) $w=9$.
+We next trained MVAR and LSTM models to learn ROMs for the latent dynamics (see Section 2.2.3), where the latent variables $\boldsymbol{y}_{c}(t_{k})\in\mathbb{R}^{6}$ of the $c$ case at timestep $t_{k}$ are obtained from the density profiles of the training set via the restriction operator in Eq.
+(34).
+To determine the lag window, we employed the BIC and AIC criteria described in D; BIC indicated a lag window of $w=4$, whereas AIC favored a wider (as expected) $w=9$.
 Both windows were used for training both MVAR and LSTM models, denoted from now on as MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models.
@@
 For the LSTMs, as each model was trained over 50 runs with different parameter initializations, we report in Table 1 the mean and the 10–90% error percentiles across the 50 runs.
-Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models at the latent space.
+Table 1: Training results of MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models at the latent space for the unidirectional flow case.
 Loss functions $\mathcal{L}_{\text{MVAR}}\left(\cdot\right)$ in Eq.
 (41) for MVARs and $\mathcal{L}_{\text{LSTM}}\left(\cdot\right)$ in Eq.
-(30) for LSTMs at the end of training, along with computational times (in seconds) required for training.
+(46) for LSTMs are reported at the end of training, along with computational times (in seconds) required for training.
 For the LSTMs, the mean and 10–90% error percentiles (in parentheses) are reported across the 50 training processes with different parameter initializations.
@@
 Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s)
-MVAR(4) $1.91\times 10^{-4}$ $1.41\times 10^{-13}$ - $0.0021$
-MVAR(9) $1.34\times 10^{-4}$ $7.05\times 10^{-14}$ - $0.0045$
-LSTM(4) - - $2.35\ (2.22,2.50)\times 10^{-5}$ 252(236,302)252\ (236,302)
-LSTM(9) - - $1.81\ (1.68,1.93)\times 10^{-5}$ 310(302,326)310\ (302,326)
+MVAR(4) $4.65\times 10^{-8}$ $3.89\times 10^{-13}$ - $0.0013$
+MVAR(9) $2.04\times 10^{-8}$ $1.72\times 10^{-13}$ - $0.0017$
+LSTM(4) - - $7.12\ (6.27,8.16)\times 10^{-9}$ 62(55,66)62\ (55,66)
+LSTM(9) - - $1.51\ (1.30,1.81)\times 10^{-8}$ 86(77,89)86\ (77,89)
 Since the loss function of MVARs is $L_{2}$-based, we additionally compute the MSE in Eq.
 (50) to enable direct comparison with the MSE-based LSTM loss.
-MVARs are able to reach much smaller loss function values than LSTMs, providing a hint for their better accuracy.
+MVARs converge to much smaller loss function values in comparison to LSTMs.
 As expected, LSTMs require significantly higher computational time for training in comparison to MVARs.
-Table 1 further shows that the training of both types of ROMs converges to lower loss function values with increased lag window, albeit requiring significantly more computational time.
-One-Step/Open-loop Predictions
-Having constructed and trained all components of our framework, we now evaluate its end-to-end performance in one-step predictions in the ambient–density profile–space using the open-loop time-stepper in Eq.
-(35); note that for the LSTMs, we consider the model with the best training performance among the 50 model runs.
+Table 1 further shows that the training of both types of ROMs with increased lag window requires more computational time.
+Recursive/Closed-loop
+Predictions Having constructed and trained all components of our framework, we evaluated its end-to-end performance in the ambient–density profile–space.
+While the ROMs were trained in an open-loop setting, we assess them in the practical closed-loop setting (see Eq.
+(51)), where only the first $w$ observations are supplied and all subsequent predictions are generated autoregressively.
+For LSTM models, we consider the best-performing model from the 50 training runs.
 We quantify the prediction accuracy via the relative $L_{1}$, $L_{2}$, and $L_{\infty}$ errors, $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
 (52), computed between the “ground-truth” density profiles $x^{(c)}(t_{k})$ and the predicted/reconstructed ones $\hat{x}^{(c)}(t_{k})$ for the $C$, unseen in training, cases (of different initial conditions) included in the testing set $X_{ts}$ across time steps $k$.
-Table 2 provides a summary of the relative reconstructed errors in the ambient space, for one-step predictions, for all four ROMs, including the mean error and the 10–90% error percentiles, aggregated over all cases and time steps in the testing set.
-Table 2: Open-loop (one-step) prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models.
+Table 2 provides a summary of the closed-loop reconstructed errors, including the mean and the 10–90% error percentiles, aggregated over all cases and time steps.
+Table 2: Closed-loop/recursive prediction relative errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models for the unidirectional flow case.
 The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
@@
 Model $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
-MVAR(4) 0.127(0.072,0.201)0.127\ (0.072,0.201) 0.078(0.045,0.119)0.078\ (0.045,0.119) 0.073(0.039,0.117)0.073\ (0.039,0.117)
-MVAR(9) 0.127(0.072,0.200)0.127\ (0.072,0.200) 0.078(0.045,0.118)0.078\ (0.045,0.118) 0.072(0.039,0.116)0.072\ (0.039,0.116)
-LSTM(4) 0.280(0.265,0.325)0.280\ (0.265,0.325) 0.167(0.158,0.190)0.167\ (0.158,0.190) 0.138(0.121,0.167)0.138\ (0.121,0.167)
-LSTM(9) 0.309(0.297,0.351)0.309\ (0.297,0.351) 0.185(0.174,0.207)0.185\ (0.174,0.207) 0.135(0.115,0.167)0.135\ (0.115,0.167)
-As shown in Table 2, the open-loop (one-step) prediction accuracy with the use of MVAR models outperforms that of LSTMs; all LSTM relative errors are approximately twice as high as those of the MVAR models.
-However, LSTMs demonstrate narrower 10–90% error percentiles in comparison to MVARs, indicating less uncertainty in one-step predictions.
-Table 2 additionally shows that the accuracy of MVARs does not improve with a longer lag, while that of LSTMs slightly degrades.
-(a) Relative Errors with MVAR(4)
-(b) Relative Errors with MVAR(9)
-(c) Relative Errors with LSTM(4)
-(d) Relative Errors with LSTM(9)
-Figure 4: One-step (open-loop) prediction errors, in the ambient–density profile–space, for the testing set over time, using trained MVAR and LSTM models.
-The relative reconstructed $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq.
-(37) (using the one-step prediction in Eq.
-(35)) are shown for the MVAR(4) panel (a), MVAR(9) panel (b), LSTM(4) panel (c), and LSTM(9) panel (d) latent dynamics models.
-Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step.
-For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
-Fig.
-4 displays the average relative reconstructed errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10–90% percentile ranges, for one-step predictions, over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$.
-Fig.
-4(a,b) depict the results obtained with the MVAR(4) and MVAR(9) models, respectively, both demonstrating a remarkable approximation accuracy.
-The mean $L_{2}$ reconstructed relative errors are initially $\sim$15% and reduce to $\sim$7% after the first 50s50\ $\text${s} of the simulation.
-This level of errors matches the infimum approximation error baseline of POD reconstruction, shown in Fig.
-3(b).
-This very high approximation accuracy is not met by the LSTM(4) and LSTM(9) models, shown in Fig.
-4(c,d).
-In particular, the mean relative reconstructed $L_{2}$ errors of the LSTM(4) and LSTM(9) models are $\sim$17% and $\sim$18%, respectively, towards the end of the simulations.
-Fig.
-4 further confirms that the lag window does not affect the one-step prediction accuracy of MVARs, while it slightly reduces that of LSTMs.
-Recursive/Closed-loop Predictions
-For obtaining long-term (closed-loop) predictions with the proposed framework, we next adopt the recursive closed-loop time-stepper in Eq.
-(36), which is only supplied with the first $w$ observations and predicts the next time steps in an autoregressive fashion.
-To quantify the prediction accuracy, we again use the relative reconstructed $L_{1}$, $L_{2}$, and $L_{\infty}$ errors in the ambient space, computed for the cases included in the testing set $X_{ts}$ across time steps.
-Table 3 provides a summary of these reconstructed errors, including the mean and the 10–90% error percentiles, aggregated over all cases and time steps.
-Table 3: Closed-loop/recursive simulation errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models.
-The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
-(37) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models.
-Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
-Model $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
-MVAR(4) 0.231(0.139,0.376)0.231\ (0.139,0.376) 0.153(0.087,0.267)0.153\ (0.087,0.267) 0.141(0.074,0.247)0.141\ (0.074,0.247)
-MVAR(9) 0.184(0.114,0.292)0.184\ (0.114,0.292) 0.124(0.073,0.209)0.124\ (0.073,0.209) 0.119(0.062,0.210)0.119\ (0.062,0.210)
-LSTM(4) 0.419(0.221,0.648)0.419\ (0.221,0.648) 0.284(0.139,0.512)0.284\ (0.139,0.512) 0.281(0.115,0.501)0.281\ (0.115,0.501)
-LSTM(9) 0.433(0.230,0.590)0.433\ (0.230,0.590) 0.310(0.145,0.460)0.310\ (0.145,0.460) 0.285(0.121,0.441)0.285\ (0.121,0.441)
-As shown in Table 3, the long-term (closed-loop) prediction accuracy of the proposed framework with MVAR models outperforms significantly that of LSTM models.
-In addition, MVARs demonstrate narrower 10–90% error percentiles than LSTMs, indicating more reliable long-term predictions.
-Table 3 further shows that the consideration of a longer lag makes MVARs more accurate, while slightly degrading the accuracy of LSTMs.
-We next assessed the evolution/accumulation of the relative reconstructed errors $L_{2}$ and $L_{\infty}$ over time for the closed-loop setting.
-Fig.
-5 depicts the average reconstructed relative errors in the ambient–density profile–space, and their 10–90% error percentile ranges across the multiple cases of different, unseen initial conditions in the testing set for all four ROMs.
+MVAR(4) 0.180(0.085,0.246)0.180\ (0.085,0.246) 0.153(0.077,0.211)0.153\ (0.077,0.211) 0.172(0.095,0.242)0.172\ (0.095,0.242)
+MVAR(9) 0.163(0.087,0.228)0.163\ (0.087,0.228) 0.140(0.079,0.194)0.140\ (0.079,0.194) 0.160(0.092,0.230)0.160\ (0.092,0.230)
+LSTM(4) 0.177(0.128,0.235)0.177\ (0.128,0.235) 0.156(0.115,0.209)0.156\ (0.115,0.209) 0.185(0.133,0.243)0.185\ (0.133,0.243)
+LSTM(9) 0.165(0.112,0.223)0.165\ (0.112,0.223) 0.142(0.110,0.190)0.142\ (0.110,0.190) 0.166(0.116,0.227)0.166\ (0.116,0.227)
+As shown in Table 2, the proposed framework with the use of MVAR models slightly outperforms that of LSTMs in the closed-loop setting across error metrics.
+The MVAR(9) achieves the lowest errors, indicating accuracy in long-term predictions.
+The MVAR(4) also performs well, with broader percentiles across test cases.
+Increasing the lag window from $w=4$ to $w=9$ slightly improves MVAR accuracy, demonstrating the benefit of incorporating longer temporal dependencies within the linear autoregressive structure.
+The LSTM models exhibit slightly higher errors than MVARs, with tighter 10–90% percentiles, while the increase of lag window shows a similar to MVARs improvement.
+We next assessed the evolution/accumulation of the relative reconstructed errors $L_{2}$ and $L_{\infty}$ over time.
+Figure 4 depicts the average reconstructed relative errors in the ambient–density profile–space, and their 10–90% error percentile ranges across the multiple cases of different, unseen initial conditions in the testing set for all four ROMs.
 (a) Relative Errors with MVAR(4)
@@
 (f) Relative Errors with SFM, $1\%$ perturbation
-Figure 5: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the trained MVAR and LSTM models.
-Panels (a)–(d) display the relative $L_{2}$ (red) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq.
+Figure 4: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the trained MVAR and LSTM models for the unidirectional flow case.
+Panels (a)–(d) display the relative $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq.
 (52) for MVAR(4), MVAR(9), LSTM(4) and LSTM(9) models, respectively.
@@
 Indicatively, for our illustrations, to provide a comparative “baseline” of the reconstructed error for tiny perturbations of the density profiles, we quantified the sensitivity of the SFM simulations per se with respect to small perturbations of the initial conditions (see in Section 3.1).
-Specifically, we introduced $\pm 1\%$ perturbations in the initial conditions of the density profiles of the reference test cases and computed the relative reconstructed errors between these perturbed initial conditions and their unperturbed counterparts.
-The MVAR models Fig.
-5(a,b) demonstrate a remarkable approximation accuracy, with the mean relative $L_{2}$ error remaining bounded below $\sim$14%(17%) for the MVAR(9)(MVAR(4)) models, despite the long-term horizon.
-The respective 10–90% error percentiles are bounded in (10, 30)% relative errors for the MVAR(4) model, while for the MVAR(9) model, the 10–90% error percentiles reduce further to (7, 21)% relative errors.
-The mean values of errors, and especially of the MVAR(9) model, is well within the 10–90% error percentiles of the system’s deviation/“baseline” 10–90% $L_{2}$ error percentiles (2, 18)% relative error, as quantified by the SFM simulations of the testing set using the $\pm$ 1% perturbed initial conditions for the density profiles, shown in Fig.
-5(e,f).
-In contrast, the LSTM models Fig.
-5(c,d) provide less accurate predictions than the MVAR models, with mean relative $L_{2}$ error plateauing at $\sim$35% for both LSTM(4) and LSTM(9) models, towards the end of the simulation, well above the expected “baseline” error bounds.
-For completeness, we directly compare the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted fields $\hat{x}^{(c)}(t_{k})$ generated by the closed-loop/recursive simulations.
-This analysis uses a case from the test dataset featuring a Gaussian initialization corresponding to the 6th row of Table B.2.
-Figure 6 displays such comparison for the MVAR(9) model at four selected time steps (k=250, 500, 780, 930k=250,\ 500,\ 780,\ 930), during which the pedestrian crowd does not pass through the corridor boundary.
-Panels (a,c,e,g) show the ground truth at these time steps, while panels (b,d,f,h) present the corresponding MVAR(9) model predictions; analogous results for the MVAR(4), LSTM(4) and LSTM(9) models are provided in Appendix E, Figs.
-E.1, E.2 and E.3, respectively.
-(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
-(b) Predicted density at $t_{250}=62.5\,\text{s}$.
-(c) Ground truth density at $t_{500}=125\,\text{s}$.
-(d) Predicted density at $t_{500}=125\,\text{s}$.
-(e) Ground truth density at $t_{780}=195\,\text{s}$.
-(f) Predicted density at $t_{780}=195\,\text{s}$.
-(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
-(h) Predicted density at $t_{930}=232.5\,\text{s}$.
-Figure 6: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
-(36) using the MVAR(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
-Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
-The predictions from all four surrogate models successfully preserve the global structure and propagation direction of the ground truth density, demonstrating the framework’s capacity to capture dominant crowd dynamics and produce meaningful long-horizon predictions.
-The models also correctly respect the obstacle regions.
-The primary observed discrepancy is a minor underestimation of the peak density values coupled with the introduction of artificial density in low-density regions, a likely consequence of enforcing mass conservation via the POD restriction operator.
-This artifact is faintest in the MVAR(9) model, and remains only slightly more pronounced in the MVAR(4).
-By contrast, it becomes considerably stronger in the less accurate LSTM(4) and LSTM(9) models.
-MVAR vs. LSTM performance
-The comparison between model classes reveals that the use of MVARs provides a more accurate and reliable framework for long-term predictions than that of LSTMs.
-This is clearly demonstrated by their significantly lower errors and narrower error percentiles in the closed-loop setting (Table 3, Fig.
+Specifically, we introduced $\pm 1\%$ perturbations in the initial conditions of the density profiles of the reference test cases and computed the relative reconstructed errors between these perturbed initial conditions and their unperturbed counterparts, as shown in Fig.
+4(e,f).
+The time evolution in Fig.
+4 reveals a characteristic pattern across all models: errors are higher initially, decrease drastically during the first $20\,\text{s}$ as the POD reconstruction error baseline diminishes, and then slowly increase towards the end of the simulation due to error accumulation in the recursive forecasting.
+The MVAR models in Fig.
+4(a,b) demonstrate remarkable approximation accuracy with the mean relative $L_{2}$ error remaining bounded below $\sim$14%(15%) for the MVAR(9)(MVAR(4)) models by the end of the simulation horizon.
+The respective 10–90% error percentiles are bounded in (7, 25)% errors for the MVAR(4) model, while for the MVAR(9) model they tighten further to (10, 20)%.
+Both MVAR models achieve errors that compare favorably with the intrinsic system baseline obtained from $\pm 1\%$ perturbed SFM simulations, which shows mean $L_{2}$ error of $\sim 8\%$ and 10–90% percentiles $(3,32)\%$ (Fig.
+4(e,f)).
+The LSTM models in Fig.
+4(c,d) provide slightly less accurate predictions than the MVAR models.
+By the end of the simulation, the LSTM(4) mean relative $L_{2}$ error reaches approximately $16\%$, with 10–90% percentiles widening to $(11,25)\%$.
+The LSTM(9) improves marginally, with mean error around $13\%$ and tighter percentiles $(10,18)\%$.
+The MVAR and LSTM models demonstrate similar trends, with MVARs providing slightly higher accuracy; a result confirming the averaged in time results reported in Table 2.
+For completeness, we directly compared the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted fields $\hat{x}^{(c)}(t_{k})$ generated by the closed-loop/recursive simulations.
+The predictions of the proposed framework for all ROMs (see Figs E.2E.5) successfully preserve the global structure and propagation direction of the ground truth density, demonstrating the framework’s capacity to capture dominant crowd dynamics and produce meaningful long-horizon predictions.
+The spatial distribution of the absolute errors between ground truth and closed-loop predictions (see Fig.
+E.6) further shows that the all models correctly respect the obstacle regions.
+The primary observed discrepancy is a minor underestimation of the peak density values, and of density in regions immediately downstream and upstream of these peaks coupled with the introduction of artificial density in low-density regions; the latter is a likely consequence of enforcing mass conservation via the POD restriction operator.
+These artifacts are fainter in the MVAR(9) and LSTM(9) models, and become slightly stronger in the less accurate MVAR(4) and LSTM(4) models, showing a lessening effect as time passes (see Fig.
+E.6).
+Finally, the open-loop (one-step) prediction results are provided in E; see Table E.1 and Fig.
+E.1.
+Similarly to the closed-loop findings above, all four ROMs achieve comparable accuracy in the open-loop setting, with mean $L_{2}$ errors of approximately $0.038$ across models and narrow 10–90% percentiles.
+This indicates that while both models capture the local dynamics very accurately in one-step predictions ($L_{2}$ errors just above the POD baseline), their long-term forecasting is, in practice, affected by error accumulation that can degrade their performance.
+4.2 Counterflow case study
+For the second case study, we applied the proposed framework in a counterflow configuration in which two interacting populations are divided equally into two groups with opposing flow directions; see Fig.
+2(b).
+Following the same procedure as in the unidirectional flow case (for implementation details see Sections 3.23.4), we constructed the restriction and lifting operators in Eqs.
+(36) and (37) by retaining $d_{1}=6$ and $d_{2}=8$ POD modes per group $l=1,2$ and $m=4$ cross-covariance modes.
+The accuracy of the resulting operators is very high, with relative $L_{2}$ errors reducing to $3-4\%$ after the first $50\,\text{s}$ for both groups (see Fig.
+F.1 in F).
+With latent variables $\boldsymbol{y}_{c}(t_{k})=[\boldsymbol{y}^{(1)}_{c}(t_{k}),\boldsymbol{y}^{(2)}_{c}(t_{k})]\in\mathbb{R}^{24}$ determined, we trained MVAR and LSTM models following Section 3.5.
+The AIC-identified lag was $w=10$, and stationarity was verified (ADF test).
+The MVAR(10) again converged to lower loss function values with substantially less computational time than the LSTM(10) one (see Table F.1 in F), though both models show deteriorated performance compared to the unidirectional flow case.
+To evaluate the end-to-end performance of the proposed framework in the ambient–density profile–space in long-term (closed-loop) predictions, we followed the same quantification procedure as in Section 4.1.
+We again observe that the closed-loop prediction accuracy with the use of MVAR is superior to that of LSTM (see Table F.2 in F).
+In particular, the mean, overall testing cases and time steps, reconstructed $L_{2}$ error of the MVAR(10) model is $8\%$ with tight 10–90% percentiles, while that of the LSTM(10) model is $\sim 10\%$ with wider percentiles for both population groups.
+The detailed error evolution over time is displayed in Fig.
+5.
+The MVAR(10) model shows mean $L_{2}$ errors dropping to $\sim 9-10\%$ within the first $20\,\text{s}$ for both population groups.
+These errors remain bounded over the long-time horizon at values slightly over $10\%$, exhibiting high accuracy in comparison to the baseline accuracy of pure restriction and lifting operators ($3-4\%$ after the first $50\,\text{s}$ for both groups; see Fig.
+F.1).
+In contrast, LSTM(10) exhibits slightly larger errors that stabilize at $13-14\%$ by the simulation end.
+Across both models, group 2 shows slightly higher errors and wider uncertainty bands.
+(a) Relative Errors with MVAR(10) for group 1
+(b) Relative Errors with MVAR(10) for group 2
+(c) Relative Errors with LSTM(10) for group 1
+(d) Relative Errors with LSTM(10) for group 2
+Figure 5: Closed-loop/recursive simulation errors in the ambient–density profile–space, across the testing set over time, using the MVAR and LSTM models for the counterflow case.
+Panels (a,b) and (c,d) display the relative reconstructed $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ in Eq.
+(52) for MVAR(10) and LSTM(10) models, respectively.
+The left and right column panels show errors for group 1 and 2, respectively.
+Mean relative errors (solid) and 10–90% error percentiles (dashed) are reported over $C=20$ cases per time step.
+For the LSTM model, results correspond to the run achieving the best training performance out of the 50 independent initializations.
+Additionally, we provide a direct comparison of the ground truth densities with their closed-loop predictions for both groups (see Figs.
+F.3 and F.4 in F).
+Both MVAR and LSTM models accurately reproduce the two dominant density lobes and capture migration and alignment patterns of the two interacting groups.
+These results, together with the spatial distribution of the absolute errors between ground truth and closed-loop predictions (see Fig.
+F.5), further show that both models respect obstacle regions, without significant errors accumulating in the immediate obstacle proximity, while accurately reproducing localized congestion zones formed upstream and downstream of the obstacle.
+Minor discrepancies include peak attenuation and spatial dispersion of the Gaussian-like lobe structure of each group, which are slightly more pronounced with the LSTM(10) model, especially as time progresses; a finding that is consistent with LSTM’s higher error metrics compared to MVAR in Fig.
+5.
+For completeness, we also provide the open-loop/one-step prediction results in Table F.2 and Fig.
+F.2 of F.
+As in the unidirectional flow case, both MVAR and LSTM models achieve comparable accuracy with mean $L_{2}$ errors $\sim 4-6\%$ and very narrow 10–90% percentiles.
+This very accurate performance in one-step predictions demonstrates once again the effect of error accumulation introduced in the recursive predictions for long-term horizons.
+Finally, the proposed framework achieves higher accuracy in the counterflow case than in the unidirectional flow case, and this difference is consistent across ROM model selections.
+This may be attributed to the larger latent dimension ($24$ vs $6$), which provides a richer representation space.
+However, this comes at increased computational cost, highlighting the accuracy-efficiency trade-off inherent in dimensionality selection.
+4.3 MVAR vs. LSTM performance
+The comparison between model classes in both unidirectional and counterflow case studies reveals that the use of MVARs provides a slightly more accurate and reliable framework for long-term predictions than that of LSTMs.
+This is clearly demonstrated by their lower errors and narrower error percentiles in the closed-loop setting (Table 2, Fig.
+4 and Table F.2, Fig.
 5).
-Crucially, this performance gap widens substantially when compared to the open-loop results (Table 2, Fig.
-4), where LSTMs performed more competitively.
-For instance, the MVAR(9) model’s mean $L_{2}$ error across time increased from $\sim$7% (open-loop) to $\sim$14% (closed-loop) within the “baseline” reconstruction error from POD and perturbed initial conditions, respectively, while the LSTM(9) model’s error changing respectively from $\sim$19% to $\sim$35% is well above these “baselines”.
-This superior performance of MVARs is attributed to the challenge of error accumulation in closed-loop forecasting; the linear structure of the MVAR model proves to be much less susceptible to this instability than complex, nonlinear LSTMs.
-While LSTMs can theoretically capture richer nonlinear dynamics, their performance severely degrades due to the “curse of dimensionality” in their training, i.e., finding a global optimum, accompanied by a distribution shift [99, 100] (i.e., the error accumulation of the recursive scheme during inference, that shifts the distribution upon which the models were trained), despite their ability to learn the one-step dynamics effectively.
+While all models perform remarkably well in an open-loop setting (see Table E.1, Fig.
+E.1 and Table F.2, Fig.
+F.2), the effect of error accumulation in recursive predictions degrades their predictive accuracy.
+For instance, in the unidirectional flow case the MVAR(4) model’s mean $L_{2}$ error across time increased from $\sim$4% (open-loop) to $\sim$15% (closed-loop), and the LSTM(4) model’s error increased from $\sim$4% to $\sim$16%.
+This consistent pattern in both case studies with different latent dimensions ($6$ vs $24$) and interaction complexities is attributed to the challenge of error accumulation in closed-loop forecasting; the linear structure of the MVAR model proves to be less susceptible to this instability than complex, nonlinear LSTMs.
+While LSTMs can theoretically capture richer nonlinear dynamics, their performance degrades due to the “curse of dimensionality” in their training, i.e., finding a global optimum, accompanied by a distribution shift [20, 54] (i.e., the error accumulation of the recursive scheme during inference, that shifts the distribution upon which the models were trained), despite their ability to learn the one-step dynamics effectively.
 In addition, there are several challenges regarding the “correct” initialization of LSTMs [70], especially in dynamical systems and multiscale modeling tasks.
-In contrast, the training of the linear MVAR models is based only on OLS, thus giving a global optimum.
-For the particular applications, the results suggest that MVARs are a much better choice for long-term horizon predictions in the latent space.
+In contrast, the training of the linear MVAR models is based on least squares regression, thus giving a global optimum.
+For the particular applications, the results suggest that MVARs are a better choice for long-term horizon predictions in the latent space.
 This is a result in line also with other studies (see e.g., [91]).
-Beyond accuracy, the proposed framework offers a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics.
-In particular, for the benchmark case of $100$ pedestrians moving in the corridor for $250\text{s}$, the conventional numerical integration requires a computational time of $20\text{s}$ for the microscopic SFM simulator simulation plus $30\text{s}$ for the extraction of the density fields (in a coarser time discretization), summing up to a total of $50\text{s}$.
-In contrast, our framework requires just $0063\text{s}, 0.0$$0161\text{s}$ for the MVARs models and $6.75\text{s}$, $6.77\text{s}$ for the LSTM ones to complete the predictions across the long-time horizon integration–representing a speed-up of 3,100x and 7.4x for the MVARs and LSTMs, respectively.
-Even when accounting for the offline costs (POD basis construction and ROM training), MVARs retain a significant advantage, with training times under $0.005\text{s}$ (as shown in Table 1).
-The MVARs variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts.
+The proposed framework offers a balanced trade-off between prediction accuracy and computational efficiency for coarse-grained crowd dynamics.
+For the unidirectional flow case ($250\,\text{s}$ horizon, $K=1000$ steps), one conventional simulation requires $80\,\text{s}$ for SFM integration plus approximately $18\,\text{s}$ for density field extraction via KDE, totaling $98\,\text{s}$ per complete prediction.
+Our framework’s online closed-loop prediction requires $w\times(t_{ID}+t_{R})\,\text{s}$ for density extraction and restriction of the first $w$ lagged snapshots, followed by $(1000-w)\times(t_{F}+t_{L})$ for per-step ROM forecasting and lifting operations; see Eq.
+(51).
+This results in total online execution times of $0.397$, $0.887$, $1.413$, and $1.871\,\text{s}$ for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models, respectively, corresponding to speed-ups of $247\times$, $110\times$, $69\times$, and $52\times$ relative to the conventional approach.
+For the counterflow case, one conventional simulation requires a total of $77.35\,\text{s}$, while our framework yields total online execution times of $0.781$ and $2.141\,\text{s}$ for the MVAR(10) and LSTM(10) models, respectively, corresponding to speed-ups of $99\times$ and $36\times$.
+These online times include lifting at every time step for a fair comparison with the conventional simulation output.
+In practice, with on-demand lifting (i.e., applying $\mathcal{L}$ only when density fields are needed), the speed-ups would be even greater, as the framework evolves in the latent space (Eq.
+(51)).
+Offline costs (one-time) for the unidirectional flow case include $(80+18)\times 20=1960\,\text{s}$ for generating $20$ training trajectories via SFM with KDE extraction, $4.24\,\text{s}$ for POD basis computation, and model training times of $0.0013$, $0.0017$, $62$, and $86\,\text{s}$ for MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models respectively (the LSTMs times are averaged over 50 runs; see Table 1).
+For the counterflow case, the offline costs are $1547\,\text{s}$ for generating $20$ training trajectories, $15.51\,\text{s}$ for the augmented POD basis, and model training times of $0.031$ and $104\,\text{s}$ for MVAR(10) and LSTM(10) models respectively (see Table F.1).
+While offline costs appear significant, they become negligible when amortized over multiple predictions.
+For example, the total offline cost for LSTM(9) in the unidirectional flow case ($2050.24\,\text{s}$) equals just $21$ conventional simulations; a cost quickly recouped when dozens or hundreds of predictions are required, as in parameter studies or real-time applications.
+Overall, MVAR variants prove particularly efficient for real-time applications, while maintaining better accuracy than their LSTM counterparts.
 These gains are especially valuable for real-world scenarios, where the number of pedestrians renders the conventional approaches infeasible.
@@
 By first encoding the full high‐dimensional state into a compact latent representation, we learn the system’s evolution operator using surrogate ROMs, bypassing the prohibitive computational cost of learning PDEs via, for example, CNNs or other high‐dimensional ML models such as DNNs.
-Working in latent space with delayed POD coordinates simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics.
-In fact, under the usual generic assumptions of Takens’ theorem, applying time-delay embeddings to POD modes produces a delay-coordinate latent model that is diffeomorphic to the original system dynamics (see also results and discussion in [67, 62, 102, 68, 69]).
-Thus, our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution.
+Working in latent spaces with delayed POD coordinates simplifies the enforcement of physical constraints (such as mass conservation) over long time horizons, since, as we demonstrate, the POD algorithm used to find an appropriate set of coordinates of the latent space explicitly conserves the mass of the reconstructed dynamics.
+In fact, under the usual generic assumptions of Takens’/ Whitneys’ embeddings theorems, applying time-delay embeddings to POD modes produces a delay-coordinate latent model that is diffeomorphic to the original system dynamics (see also results and discussion in [91, 102, 71, 9, 44]).
+For our illustrations, we assessed our approach using data generated from the SFM in a corridor with an obstacle under both unidirectional and counterflow scenarios, though more advanced crowd models incorporating behavioral aspects or optimal path planning [15, 34] can be readily employed.
+Our approach does not attempt to learn any closed‑form PDE, but instead learns the effective solution operator of such an unavailable PDE, mapping initial crowd density to its future evolution.
 The resulting ROMs trained in the latent space, in the form of linear (MVARs) or nonlinear multivariate autoregressive models (such as LSTMs), encode and reconstruct this operator numerically via the POD reconstruction.
 As a result, our framework allows for rapid and accurate “what‐if” scenarios exploration and real‐time forecasting of crowd dynamics.
-Critically, we demonstrate that linear MVAR models within our framework not only outperform more complex LSTMs in long-term predictive accuracy and reliability, but also achieve a computational speed-up of over three orders of magnitude compared to conventional simulation.
+Critically, we demonstrate that linear MVAR models within our framework not only outperform more complex LSTMs in long-term predictive accuracy and reliability, but also achieve a computational speed-up of over two orders of magnitude compared to conventional simulation.
 This combination of interpretability, speed, and accuracy makes our approach, particularly with MVARs, an efficient approach for large‐scale, real-world crowd‐dynamics analysis and distributed control [36, 5, 8, 53, 27, 6, 82].
-Finally, our finding that simple linear MVAR models outperform nonlinear LSTM networks in the latent space in terms of complexity, training time, and interpretability is a significant insight.
-It aligns with a growing body of literature [62, 102, 69, 63] that demonstrates the power of linear ROMs constructed in judiciously chosen latent spaces (e.g., POD-based delay-coordinate embeddings, Koopman operator) for approximating complex nonlinear or chaotic dynamics, offering a promising path toward interpretable and efficient data-driven modeling.
+The predictive performance of linear MVAR models in our framework (consistently matching or exceeding that of nonlinear LSTMs) aligns with a growing body of literature [91, 49, 9, 44] demonstrating that linear ROMs constructed in judiciously chosen latent spaces (e.g., POD-based delay-coordinate embeddings, Koopman operator) can effectively approximate complex nonlinear or chaotic dynamics.
+For example, nonlinear manifold learning techniques such as Diffusion Maps have been shown to provide more accurate embeddings of complex dynamics compared to POD, particularly during transients [37].
+As formulated in its vanilla form, the proposed method inherits the well-known limitations of POD-based reduced-order models, and in particular their strong dependence on the geometries and boundary conditions represented in the training data, and the present algorithm is no exception: if the test configuration departs substantially away from the snapshot ensemble, the linear subspace may fail to represent the solution manifold accurately.
+Handling boundary-aware cases would require a different modeling approach, such as NOs.
+Incorporating NOs in latent spaces within the proposed framework is a promising direction which will be investigated in future studies, including a broader comparison with alternative surrogate models as for example via Transformer-based architectures [60, 101] that have shown a good performance for long-horizon modeling, and latent Neural ODEs [98], which offer a continuous-time perspective for learning reduced dynamics.
+Moreover, approaches based on Learning Effective Dynamics (LED) [71, 24, 109, 42] and its extensions, such as G-LED [50] and interpretable LED [84], provide complementary frameworks that emphasize structure preservation and interpretability of the learned operators.
+A vital extension of this work for real-world deployment is the transition to probabilistic forecasting to quantify predictive uncertainty.
+While techniques such as Monte Carlo dropout or quantile regression could be applied, the linear Gaussian structure of the MVAR model offers a particularly direct and interpretable pathway to achieve this by propagating initial condition uncertainty through its stochastic latent dynamics.
+Finally, the proposed mass-conserving framework provides a foundation for applications to real-world trajectory data, systematic benchmarking against alternative surrogates, and extension to other collective motion systems such as vehicular traffic.
 Acknowledgments
@@
 where $C_{i}\in\mathbb{R}^{+}$ is the strength of the repulsive interaction against a wall or an obstacle, $D_{i}$ is its corresponding decay parameter, and $d_{iB}$ is the euclidean distance between the pedestrian and the closest point of the object $B$; $\bm{n}_{iB}$ is the respective the unit vector.
-Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exerft dynamic resistance.
+Unlike pedestrian–pedestrian interactions, no tangential (sliding) friction is typically included for static obstacles, as they do not exert dynamic resistance.
 Table A.1: Social Force Model simulation parameters.
@@
 Although individual variations exist in reality, using a common set of parameters is a standard approach in the literature [61] to maintain a manageable calibration process while enabling the reliable simulation of diverse crowd dynamics, including the present case study.
-Appendix B Initial Conditions of Microscopic Distributions
-Here, we describe the initial conditions considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested.
-We considered zero initial velocities for all pedestrians, while we used the following distributions for initializing the position $\bm{x}_{i}=(x_{i},y_{i})\in\Omega$ of the $i$-th pedestrian.
-Positions are sampled from:
-  • Uniform distributions over the domain $x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\subset\Omega$:
-    $p(x_{i},y_{i})=\mathcal{U}\left([x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\right).$ (B.1)
-  • Two dimensional normal distributions with means $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
-    $p(x_{i},y_{i})=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\left(-\frac{(x_{i}-\mu_{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y_{i}-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right),$ (B.2)
-  • “Double bell-shaped” independent joint distributions, which are comprised by normal distributions along the yy–axis with mean $\mu_{y}$ and standard deviation $\sigma_{y}$, and a mixture of normal distributions along the xx–axis with means $(\mu_{x_{1}},\mu_{x_{2}})$ and standard deviations $(\sigma_{x_{1}},\sigma_{x_{2}})$:
-    $p(x_{i},y_{i})$ $=\frac{1}{\sqrt{2\pi}\,\sigma_{y}}\,\exp\!\left(-\frac{(y-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right)$ (B.3)
-    $\quad\times\frac{1}{2}\left[\frac{1}{\sqrt{2\pi}\,\sigma_{x_{1}}}\exp\!\left(-\frac{(x-\mu_{x_{1}})^{2}}{2\sigma_{x_{1}}^{2}}\right)+\frac{1}{\sqrt{2\pi}\,\sigma_{x_{2}}}\exp\!\left(-\frac{(x-\mu_{x_{2}})^{2}}{2\sigma_{x_{2}}^{2}}\right)\right].$
-  • Cosine distributions centered at $(c_{x},c_{y})$ with scale parameters $(s_{x},s_{y})$:
-    $p(x_{i},y_{i})=\frac{1}{\pi s_{x}s_{y}}\cos\left(\frac{x_{i}-c_{x}}{s_{x}}\right)\cos\left(\frac{y_{i}-c_{y}}{s_{y}}\right).$ (B.4)
-The training and testing data sets were generated from $C=20$ initial conditions, with 5 samples drawn from each of the four distributions described above.
-To ensure plausible and varied crowd configurations, the parameters for these distributions were selected heuristically to provide a broad spatial coverage while mitigating boundary artifacts.
-Furthermore, during the sampling process, pedestrian positions were resampled if they were within a critical distance equal to $2r_{i}$ of an existing pedestrian or a wall/obstacle to prevent initial overlaps and collisions.
-The specific parameter configurations used for the training and testing data sets are detailed Table B.1 and Table B.2, respectively.
-Table B.1: Microscopic initial distribution configurations and parameters for constructing the training dataset.
-Case c Distribution Type Parameters
-1 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=15,y_{\text{min}}=3,y_{\text{max}}=9$
-2 Uniform, Eq. (B.1) $x_{\text{min}}=33,x_{\text{max}}=46,y_{\text{min}}=3,y_{\text{max}}=9$
-3 Uniform, Eq. (B.1) $x_{\text{min}}=2,x_{\text{max}}=20,y_{\text{min}}=4,y_{\text{max}}=8$
-4 Uniform, Eq. (B.1) $x_{\text{min}}=28,x_{\text{max}}=46,y_{\text{max}}=4,y_{\text{max}}=8$
-5 Uniform, Eq. (B.1) $x_{\text{min}}=3,x_{\text{max}}=17,y_{\text{max}}=4,y_{\text{max}}=8$
-6 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
-7 Gaussian, Eq. (B.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
-8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
-9 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
-10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=5,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
-11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=8,\;\mu_{y}=7,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=0.8$
-12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
-13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=28,\;\mu_{y}=6,\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=32,\;\sigma_{x_{2}}=1.1$
-14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=36,\;\mu_{y}=6,\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=40,\;\sigma_{x_{2}}=1.2$
-15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=9,\;\mu_{y}=6,\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=1$
-16 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
-17 Cosine, Eq. (B.4) $c_{x}=20,\;c_{y}=8,s_{x}=10,\;s_{y}=3$
-18 Cosine, Eq. (B.4) $c_{x}=30,\;c_{y}=7,s_{x}=10,\;s_{y}=3$
-19 Cosine, Eq. (B.4) $c_{x}=40,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
-20 Cosine, Eq. (B.4) $c_{x}=10,\;c_{y}=6,s_{x}=9,\;s_{y}=3$
-Table B.2: Microscopic initial distribution configurations and parameters for constructing the testing dataset.
-Case c Distribution Type Parameters
-1 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=16,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
-2 Uniform, Eq. (B.1) $x_{\text{min}}=32,\;x_{\text{max}}=46,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
-3 Uniform, Eq. (B.1) $x_{\text{min}}=2,\;x_{\text{max}}=19,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
-4 Uniform, Eq. (B.1) $x_{\text{min}}=29,\;x_{\text{max}}=46,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
-5 Uniform, Eq. (B.1) $x_{\text{min}}=4,\;x_{\text{max}}=17,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
-6 Gaussian, Eq. (B.2) $\mu_{x}=12,\;\mu_{y}=7,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
-7 Gaussian, Eq. (B.2) $\mu_{x}=11,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=2$
-8 Gaussian, Eq. (B.2) $\mu_{x}=38,\;\mu_{y}=6,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
-9 Gaussian, Eq. (B.2) $\mu_{x}=39,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
-10 Gaussian, Eq. (B.2) $\mu_{x}=37,\;\mu_{y}=7,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
-11 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=7,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.1,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=0.9$
-12 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
-13 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=29,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=33,\;\sigma_{x_{2}}=1.1$
-14 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=37,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=41,\;\sigma_{x_{2}}=1.2$
-15 Double Gaussian, Eq. (B.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\;\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=1$
-16 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=4,\;s_{x}=8,\;s_{y}=3$
-17 Cosine, Eq. (B.4) $c_{x}=25,\;c_{y}=6,\;s_{x}=10,\;s_{y}=3$
-18 Cosine, Eq. (B.4) $c_{x}=35,\;c_{y}=5,\;s_{x}=10,\;s_{y}=3$
-19 Cosine, Eq. (B.4) $c_{x}=45,\;c_{y}=7,\;s_{x}=5,\;s_{y}=3$
-20 Cosine, Eq. (B.4) $c_{x}=15,\;c_{y}=6,\;s_{x}=7,\;s_{y}=3$
-Appendix C Density fields extraction via Kernel Density Estimation
+Appendix B Density fields extraction via Kernel Density Estimation
 Here, we provide additional details for the first step of the proposed framework, regarding the extraction of continuous density profiles from discrete pedestrians’ positions.
@@
 Its multivariate form is given by:
-$K_{h}(\mathbf{x}-\mathbf{x}_{i}(t))=\frac{1}{2\pi\det(\mathbf{H})^{1/2}}\exp\left(-\tfrac{1}{2}(\mathbf{x}-\mathbf{x}_{i}(t))^{\top}\mathbf{H}^{-1}(\mathbf{x}-\mathbf{x}_{i}(t))\right),$ (C.1)
-where $\mathbf{H}$ is the symmetric positive-definite bandwidth matrix, $\mathbf{x}\in\mathbb{R}^{2}$ is the spatial domain and $\mathbf{x}_{i}(t)\in\mathbb{R}^{2}$ is the position of the $i$-th pedestrian.
+$K_{h}(\mathbf{x}-\mathbf{x}_{i}(t))=\frac{1}{2\pi\det(\mathbf{H})^{1/2}}\exp\left(-\tfrac{1}{2}(\mathbf{x}-\mathbf{x}_{i}(t))^{\top}\mathbf{H}^{-1}(\mathbf{x}-\bm{x}_{i}(t))\right),$ (B.1)
+where $\mathbf{H}$ is the symmetric positive-definite bandwidth matrix, $\mathbf{x}\in\mathbb{R}^{2}$ is the spatial domain and $\bm{x}_{i}(t)\in\mathbb{R}^{2}$ is the position of the $i$-th pedestrian.
 The choice of the bandwidth matrix, which in this case takes the form $\mathbf{H}=\operatorname{diag}(h_{x},h_{y})$, plays a central role in the KDE estimate.
@@
 Silverman’s rule of thumb [103], originally formulated for univariate distributions, extends naturally to multivariate settings, indicating an optimal bandwidth:
-$h_{i}=\sigma_{i}\left(\frac{4}{N(d+2)}\right)^{1/(d+4)},$ (C.2)
+$h_{i}=\sigma_{i}\left(\frac{4}{N(d+2)}\right)^{1/(d+4)},$ (B.2)
 where $N$ is the number of data points, $d$ is the number of spatial dimensions, in this case $d=2$, and $\sigma_{i}$ is the standard deviation of the pedestrian positions along $i=x,y$ spatial directions.
@@
 We then define the extended domain as:
-$\Omega^{\text{ext}}=[a-\Delta x,b+\Delta x]\times[c,d],$ (C.3)
+$\Omega^{\text{ext}}=[a-\Delta x,b+\Delta x]\times[c,d],$ (B.3)
 while the support in the $y$-direction remains unbounded.
@@
 Let $\epsilon=\Delta x$ define a threshold for proximity to the boundary.
-For each position $\mathbf{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at
-$\mathbf{x}_{i}^{\text{aug1}}(t)=(x_{i}(t)-(b-a),y_{i}(t)).$
-Similarly, for every $\mathbf{x}_{i}(t)$ such that $x_{i}(t)\geq b-\epsilon$, we introduce a “mirrored” pedestrians at
-$\mathbf{x}_{i}^{\text{aug2}}(t)=(x_{i}(t)+(b-a),y_{i}(t)).$
+For each position $\bm{x}_{i}(t)=(x_{i}(t),y_{i}(t))$ such that $x_{i}(t)\leq a+\epsilon$, we introduce a “mirrored” pedestrians at
+$\bm{x}_{i}^{\text{aug1}}(t)=(x_{i}(t)-(b-a),y_{i}(t)).$
+Similarly, for every $\bm{x}_{i}(t)$ such that $x_{i}(t)\geq b-\epsilon$, we introduce a “mirrored” pedestrians at
+$\bm{x}_{i}^{\text{aug2}}(t)=(x_{i}(t)+(b-a),y_{i}(t)).$
 The resulting augmented set of pedestrian positions used for KDE is then given by
-$\{\mathbf{x}_{i}^{\prime}(t)\}=\{\mathbf{x}_{i}(t)\}\cup\{\mathbf{x}_{i}^{\text{aug1}}(t)\}\cup\{\mathbf{x}_{i}^{\text{aug2}}(t)\}.$ (C.4)
-A pseudo-code for the estimation of the density in $\Omega\subset\mathbb{R}^{2}$ is given in Algorithm C.1.
-Algorithm C.1 Kernel Density Estimation (KDE) for extracting density fields from pedestrian positions.
+$\{\bm{x}_{i}^{\prime}(t)\}=\{\bm{x}_{i}(t)\}\cup\{\bm{x}_{i}^{\text{aug1}}(t)\}\cup\{\bm{x}_{i}^{\text{aug2}}(t)\}.$ (B.4)
+A pseudo-code for the estimation of the density in $\Omega\subset\mathbb{R}^{2}$ is given in Algorithm B.1.
+Algorithm B.1 Kernel Density Estimation (KDE) for extracting density fields from pedestrian positions.
 1:Number of pedestrians NN, computational domain $\Omega=[a,b]\times[c,d]$, spatial discretization $n_{x}$$n_{y}$, obstacle subdomain $\Gamma=[a_{o},b_{o}]\times[c_{o},d_{o}]\subset\Omega$ and pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\in\Omega\backslash\Gamma$
@@
 3:Set augmentation on $x$-direction $\Delta x=(b-a)/n$
-4:Define extended support $\Omega^{\text{ext}}$ to handle periodic boundary conditions $\triangleright$ Use Eq. (C.3)
+4:Define extended support $\Omega^{\text{ext}}$ to handle periodic boundary conditions $\triangleright$ Use Eq. (B.3)
 5:Discretize extended domain $\Omega^{\text{ext}}$ in $(n_{x}+2n_{x}/n)\times n_{y}$ cells
-6:Introduce “mirrored” pedestrians to form $\{\mathbf{x}_{i}^{\prime}(t)\}$ $\triangleright$ Use Eq. (C.4)
+6:Introduce “mirrored” pedestrians to form $\{\mathbf{x}_{i}^{\prime}(t)\}$ $\triangleright$ Use Eq. (B.4)
 7:Initialize zero density field $\rho((x_{i},y_{j}),t)$ across the extended domain $\Omega^{\text{ext}}$
 8:for each pedestrian $p=1,2,\ldots,N$ do $\triangleright$ Account for “mirrored” pedestrians
-9:  Compute kernel contribution $K_{h}(\mathbf{x}-\mathbf{x}_{p}(t))$ over the extended support $\Omega^{\text{ext}}$ $\triangleright$ Use Eq. (C.1)
+9:  Compute kernel contribution $K_{h}(\mathbf{x}-\mathbf{x}_{p}(t))$ over the extended support $\Omega^{\text{ext}}$ $\triangleright$ Use Eq. (B.1)
 10:  Add contribution to density field $\rho((x_{i},y_{j}),t)$
@@
 13:Retrieve density $\rho((x_{i},y_{j}),t)$ in computational domain $\Omega$
-14:Employ binary mask function to ensure zero density in obstacle subdomain $\Gamma$, by setting $\triangleright$ See Eq. (C.5)
+14:Employ binary mask function to ensure zero density in obstacle subdomain $\Gamma$, by setting $\triangleright$ See Eq. (B.5)
 […]
@@
 This is typically achieved by introducing a binary mask function $\mathcal{M}(x,y)$:
-$M(x,y)=\begin{cases}0,&\text{if }(x,y)\in\Gamma\subset\Omega,\\ 1,&\text{otherwise}.\end{cases}$ (C.5)
+$M(x,y)=\begin{cases}0,&\text{if }(x,y)\in\Gamma\subset\Omega,\\ 1,&\text{otherwise}.\end{cases}$ (B.5)
 which is then applied to the initially estimated densities to enforce zero density in $\Gamma$, implying:
-$\rho((x_{i},y_{j}),t)=\mathcal{M}(x_{i},y_{j})\,\rho((x_{i},y_{j}),t),\quad\forall(x_{i},y_{j})\in\Omega.$ (C.6)
+$\rho((x_{i},y_{j}),t)=\mathcal{M}(x_{i},y_{j})\,\rho((x_{i},y_{j}),t),\quad\forall(x_{i},y_{j})\in\Omega.$ (B.6)
+Appendix C Initial Conditions of Microscopic Distributions
+Here, we describe the initial conditions considered in the SFM to construct the datasets, upon which the proposed framework is trained and tested.
+We considered zero initial velocities for all pedestrians, while we used the following distributions for initializing the position $\bm{x}_{i}=(x_{i},y_{i})\in\Omega$ of the $i$-th pedestrian.
+Positions are sampled from:
+  • Uniform distributions over the domain $x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\subset\Omega$:
+    $p(x_{i},y_{i})=\mathcal{U}\left([x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\right).$ (C.1)
+  • Two dimensional normal distributions with means $(\mu_{x},\mu_{y})$ and standard deviations $(\sigma_{x},\sigma_{y})$:
+    $p(x_{i},y_{i})=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\left(-\frac{(x_{i}-\mu_{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y_{i}-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right),$ (C.2)
+  • “Double bell-shaped” independent joint distributions, which are comprised by normal distributions along the yy–axis with mean $\mu_{y}$ and standard deviation $\sigma_{y}$, and a mixture of normal distributions along the xx–axis with means $(\mu_{x_{1}},\mu_{x_{2}})$ and standard deviations $(\sigma_{x_{1}},\sigma_{x_{2}})$:
+    $p(x_{i},y_{i})$ $=\frac{1}{\sqrt{2\pi}\,\sigma_{y}}\,\exp\!\left(-\frac{(y-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right)$ (C.3)
+    $\quad\times\frac{1}{2}\left[\frac{1}{\sqrt{2\pi}\,\sigma_{x_{1}}}\exp\!\left(-\frac{(x-\mu_{x_{1}})^{2}}{2\sigma_{x_{1}}^{2}}\right)+\frac{1}{\sqrt{2\pi}\,\sigma_{x_{2}}}\exp\!\left(-\frac{(x-\mu_{x_{2}})^{2}}{2\sigma_{x_{2}}^{2}}\right)\right].$
+  • Cosine distributions centered at $(c_{x},c_{y})$ with scale parameters $(s_{x},s_{y})$:
+    $p(x_{i},y_{i})=\frac{1}{\pi s_{x}s_{y}}\cos\left(\frac{x_{i}-c_{x}}{s_{x}}\right)\cos\left(\frac{y_{i}-c_{y}}{s_{y}}\right).$ (C.4)
+The training and testing data sets were generated from $C=20$ initial conditions, with 5 samples drawn from each of the four distributions described above.
+To ensure plausible and varied crowd configurations, the parameters for these distributions were selected heuristically to provide a broad spatial coverage while mitigating boundary artifacts.
+Furthermore, during the sampling process, pedestrian positions were resampled if they were within a critical distance equal to $2r_{i}$ of an existing pedestrian or a wall/obstacle to prevent initial overlaps and collisions.
+The specific parameter configurations used for the training and testing data sets are detailed Table C.1 and Table C.2, respectively.
+Table C.1: Microscopic initial distribution configurations and parameters for constructing the training dataset.
+Case c Distribution Type Parameters
+1 Uniform, Eq. (C.1) $x_{\text{min}}=2,x_{\text{max}}=15,y_{\text{min}}=3,y_{\text{max}}=9$
+2 Uniform, Eq. (C.1) $x_{\text{min}}=33,x_{\text{max}}=46,y_{\text{min}}=3,y_{\text{max}}=9$
+3 Uniform, Eq. (C.1) $x_{\text{min}}=2,x_{\text{max}}=20,y_{\text{min}}=4,y_{\text{max}}=8$
+4 Uniform, Eq. (C.1) $x_{\text{min}}=28,x_{\text{max}}=46,y_{\text{max}}=4,y_{\text{max}}=8$
+5 Uniform, Eq. (C.1) $x_{\text{min}}=3,x_{\text{max}}=17,y_{\text{max}}=4,y_{\text{max}}=8$
+6 Gaussian, Eq. (C.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+7 Gaussian, Eq. (C.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
+8 Gaussian, Eq. (C.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
+9 Gaussian, Eq. (C.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+10 Gaussian, Eq. (C.2) $\mu_{x}=37,\;\mu_{y}=5,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+11 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=8,\;\mu_{y}=7,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=0.8$
+12 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
+13 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=28,\;\mu_{y}=6,\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=32,\;\sigma_{x_{2}}=1.1$
+14 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=36,\;\mu_{y}=6,\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=40,\;\sigma_{x_{2}}=1.2$
+15 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=9,\;\mu_{y}=6,\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=1$
+16 Cosine, Eq. (C.4) $c_{x}=10,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
+17 Cosine, Eq. (C.4) $c_{x}=20,\;c_{y}=8,s_{x}=10,\;s_{y}=3$
+18 Cosine, Eq. (C.4) $c_{x}=30,\;c_{y}=7,s_{x}=10,\;s_{y}=3$
+19 Cosine, Eq. (C.4) $c_{x}=40,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
+20 Cosine, Eq. (C.4) $c_{x}=10,\;c_{y}=6,s_{x}=9,\;s_{y}=3$
+Table C.2: Microscopic initial distribution configurations and parameters for constructing the testing dataset.
+Case c Distribution Type Parameters
+1 Uniform, Eq. (C.1) $x_{\text{min}}=2,\;x_{\text{max}}=16,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
+2 Uniform, Eq. (C.1) $x_{\text{min}}=32,\;x_{\text{max}}=46,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
+3 Uniform, Eq. (C.1) $x_{\text{min}}=2,\;x_{\text{max}}=19,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
+4 Uniform, Eq. (C.1) $x_{\text{min}}=29,\;x_{\text{max}}=46,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
+5 Uniform, Eq. (C.1) $x_{\text{min}}=4,\;x_{\text{max}}=17,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
+6 Gaussian, Eq. (C.2) $\mu_{x}=12,\;\mu_{y}=7,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
+7 Gaussian, Eq. (C.2) $\mu_{x}=11,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=2$
+8 Gaussian, Eq. (C.2) $\mu_{x}=38,\;\mu_{y}=6,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
+9 Gaussian, Eq. (C.2) $\mu_{x}=39,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+10 Gaussian, Eq. (C.2) $\mu_{x}=37,\;\mu_{y}=7,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
+11 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=7,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.1,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=0.9$
+12 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=10,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
+13 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=29,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=33,\;\sigma_{x_{2}}=1.1$
+14 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=37,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=41,\;\sigma_{x_{2}}=1.2$
+15 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\;\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=1$
+16 Cosine, Eq. (C.4) $c_{x}=15,\;c_{y}=4,\;s_{x}=8,\;s_{y}=3$
+17 Cosine, Eq. (C.4) $c_{x}=25,\;c_{y}=6,\;s_{x}=10,\;s_{y}=3$
+18 Cosine, Eq. (C.4) $c_{x}=35,\;c_{y}=5,\;s_{x}=10,\;s_{y}=3$
+19 Cosine, Eq. (C.4) $c_{x}=45,\;c_{y}=7,\;s_{x}=5,\;s_{y}=3$
+20 Cosine, Eq. (C.4) $c_{x}=15,\;c_{y}=6,\;s_{x}=7,\;s_{y}=3$
 Appendix D Lag window size selection via the Bayesian and the Akaike Information Criteria
@@
 $w^{*}_{\text{BIC}}=\underset{w\in\{1,\dots,w_{\text{max}}\}}{\arg\min}\,\text{BIC}(w),\quad w^{*}_{\text{AIC}}=\underset{w\in\{1,\dots,w_{\text{max}}\}}{\arg\min}\,\text{AIC}(w).$ (D.3)
-Appendix E Comparison of the “ground truth” density fields with the predicted ones, using MVAR(4), LSTM(4) and LSTM(9) models
-(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
-(b) Predicted density at $t_{250}=62.5\,\text{s}$.
-(c) Ground truth density at $t_{500}=125\,\text{s}$.
-(d) Predicted density at $t_{500}=125\,\text{s}$.
-(e) Ground truth density at $t_{780}=195\,\text{s}$.
-(f) Predicted density at $t_{780}=195\,\text{s}$.
-(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
-(h) Predicted density at $t_{930}=232.5\,\text{s}$.
+Appendix E Open-loop prediction errors and closed-loop reconstructions for the unidirectional flow case
+Here, we provide the detailed open-loop prediction errors and the closed-loop reconstructions of the proposed framework for the unidirectional flow case discussed in Section 4.1.
+The open-loop setting prediction of the proposed restrict-evolve with ROM-lift framework in Eq. (5) reads:
+$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\hat{\boldsymbol{y}}^{(c)}(t_{k})\right),\qquad\hat{\boldsymbol{y}}^{(c)}(t_{k})=\Phi_{\text{ROM}}\left(\boldsymbol{y}^{(c)}(t_{k-1}),\ldots,\boldsymbol{y}^{(c)}(t_{k-w});\mathbf{p}\right),$ (E.1)
+where $\boldsymbol{y}^{(c)}(t_{k-j})=\mathcal{R}(x^{(c)}(t_{k-j}))$ are restricted embeddings of observed densities for $j=1,\ldots,w$.
+Similarly to the close-loop predictions in Eq.
+(51), the open-loop prediction expression is used in Eq.
+(52) to quantify the one-step prediction accuracy of the proposed framework.
+Table E.1 provides a summary of the relative reconstructed errors in the ambient space, for one-step predictions, for all four ROMs, including the mean error and the 10–90% error percentiles, aggregated over all cases and time steps in the testing set.
+Table E.1: Open-loop (one-step) prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models for the unidirectional flow case.
+The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq.
+(52) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models.
+Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
+Model $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
+MVAR(4) 0.048(0.033, 0.065)0.048\ (0.033,\,0.065) 0.038(0.027, 0.051)0.038\ (0.027,\,0.051) 0.042(0.030, 0.057)0.042\ (0.030,\,0.057)
+MVAR(9) 0.047(0.032, 0.063)0.047\ (0.032,\,0.063) 0.037(0.026, 0.049)0.037\ (0.026,\,0.049) 0.041(0.030, 0.055)0.041\ (0.030,\,0.055)
+LSTM(4) 0.048(0.034, 0.065)0.048\ (0.034,\,0.065) 0.038(0.027, 0.051)0.038\ (0.027,\,0.051) 0.043(0.032, 0.058)0.043\ (0.032,\,0.058)
+LSTM(9) 0.047(0.033, 0.063)0.047\ (0.033,\,0.063) 0.037(0.026, 0.049)0.037\ (0.026,\,0.049) 0.042(0.031, 0.056)0.042\ (0.031,\,0.056)
+Figure E.1 displays the relative reconstruction errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10–90% percentile ranges for one-step predictions via the open-loop time-stepper (Eq.
+(E.1)).
+Errors are averaged over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$.
+(a) Relative Errors with MVAR(4)
+(b) Relative Errors with MVAR(9)
+(c) Relative Errors with LSTM(4)
+(d) Relative Errors with LSTM(9)
+Figure E.1: One-step (open-loop) prediction errors, in the ambient–density profile–space, for the testing set over time, using trained MVAR and LSTM models for the unidirectional flow case.
+The relative reconstructed $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq.
+(52) (using the one-step prediction in Eq.
+(E.1)) are shown for the MVAR(4) panel (a), MVAR(9) panel (b), LSTM(4) panel (c), and LSTM(9) panel (d) latent dynamics models.
+Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step.
+For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
+Figures E.2, E.3 E.4 and E.5 compare the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ for closed-loop predictions (Eq.
+(51)) using MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models, respectively.
+The test case uses a Gaussian initialization (6th row of Table C.2) at four time steps ($k=250,500,780,930$), during which the pedestrian crowd does not pass through the corridor boundary.
+Panels (a,c,e,g) show the ground truth, while panels (b,d,f,h) present the corresponding closed-loop predictions.
+The spatial distribution of the resulting absolute errors is shown in Figure E.6 at time steps $k=250,930$.
+(a) Ground truth density at $t_{250}=62.5\,\text{s}$
+(b) Predicted density at $t_{250}=62.5\,\text{s}$
+(c) Ground truth density at $t_{500}=125\,\text{s}$
+(d) Predicted density at $t_{500}=125\,\text{s}$
+(e) Ground truth density at $t_{780}=195\,\text{s}$
+(f) Predicted density at $t_{780}=195\,\text{s}$
+(g) Ground truth density at $t_{930}=232.5\,\text{s}$
+(h) Predicted density at $t_{930}=232.5\,\text{s}$
 Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
-(36) using the MVAR(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+(51) for the unidirectional flow case, using the MVAR(4) model for an unseen case of the testing set; initialization at the 6th row of Table C.2.
 Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
 (a) Ground truth density at $t_{250}=62.5\,\text{s}$.
-(b) Predicted density at $t_{250}=62.5\,\text{s}$.
-(c) Ground truth density at $t_{500}=125\,\text{s}$.
-(d) Predicted density at $t_{500}=125\,\text{s}$.
-(e) Ground truth density at $t_{780}=195\,\text{s}$.
-(f) Predicted density at $t_{780}=195\,\text{s}$.
-(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
-(h) Predicted density at $t_{930}=232.5\,\text{s}$.
+(b) Predicted density at $t_{250}=62.5\,\text{s}$
+(c) Ground truth density at $t_{500}=125\,\text{s}$
+(d) Predicted density at $t_{500}=125\,\text{s}$
+(e) Ground truth density at $t_{780}=195\,\text{s}$
+(f) Predicted density at $t_{780}=195\,\text{s}$
+(g) Ground truth density at $t_{930}=232.5\,\text{s}$
+(h) Predicted density at $t_{930}=232.5\,\text{s}$
 Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
-(36) using the LSTM(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+(51) for the unidirectional flow case, using the MVAR(9) model for an unseen case of the testing set; initialization at the 6th row of Table C.2.
 Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
-(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
-(b) Predicted density at $t_{250}=62.5\,\text{s}$.
-(c) Ground truth density at $t_{500}=125\,\text{s}$.
-(d) Predicted density at $t_{500}=125\,\text{s}$.
-(e) Ground truth density at $t_{780}=195\,\text{s}$.
-(f) Predicted density at $t_{780}=195\,\text{s}$.
-(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
-(h) Predicted density at $t_{930}=232.5\,\text{s}$.
+(a) Ground truth density at $t_{250}=62.5\,\text{s}$
+(b) Predicted density at $t_{250}=62.5\,\text{s}$
+(c) Ground truth density at $t_{500}=125\,\text{s}$
+(d) Predicted density at $t_{500}=125\,\text{s}$
+(e) Ground truth density at $t_{780}=195\,\text{s}$
+(f) Predicted density at $t_{780}=195\,\text{s}$
+(g) Ground truth density at $t_{930}=232.5\,\text{s}$
+(h) Predicted density at $t_{930}=232.5\,\text{s}$
 Figure E.4: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
-(36) using the LSTM(9) model for an unseen case of the testing set; initialization at the 6th row of Table B.2.
+(51) for the unidirectional flow case, using the LSTM(4) model for an unseen case of the testing set; initialization at the 6th row of Table C.2.
+Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
+(a) Ground truth density at $t_{250}=62.5\,\text{s}$
+(b) Predicted density at $t_{250}=62.5\,\text{s}$
+(c) Ground truth density at $t_{500}=125\,\text{s}$
+(d) Predicted density at $t_{500}=125\,\text{s}$
+(e) Ground truth density at $t_{780}=195\,\text{s}$
+(f) Predicted density at $t_{780}=195\,\text{s}$
+(g) Ground truth density at $t_{930}=232.5\,\text{s}$
+(h) Predicted density at $t_{930}=232.5\,\text{s}$
+Figure E.5: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(51) for the unidirectional flow case, using the LSTM(9) model for an unseen case of the testing set; initialization at the 6th row of Table C.2.
 Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
+(a) MVAR(4) absolute error at time $t_{250}=62.5\,\mathrm{s}$
+(b) MVAR(4) absolute error at time $t_{930}=232.5\,\mathrm{s}$
+(c) MVAR(9) absolute error at time $t_{250}=62.5\,\mathrm{s}$
+(d) MVAR(9) absolute error at time $t_{930}=232.5\,\mathrm{s}$
+(e) LSTM(4) absolute error at time $t_{250}=62.5\,\mathrm{s}$
+(f) LSTM(4) absolute error at time $t_{930}=232.5\,\mathrm{s}$
+(g) LSTM(9) absolute error at time $t_{250}=62.5\,\mathrm{s}$
+(h) LSTM(9) absolute error at time $t_{930}=232.5\,\mathrm{s}$
+Figure E.6: Absolute error between the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities for the unidirectional flow case, corresponding to the predictions in Figs.
+E.2, E.3, E.4, and E.5.
+Panels (a)–(h) show the spatial distribution of the absolute error $\lvert x^{(c)}(t_{k})-\hat{x}^{(c)}(t_{k})\rvert$ in $\Omega$ at two indicative time steps $t_{250}=62.5\,\mathrm{s}$ (left column) and $t_{930}=232.5\,\mathrm{s}$ (right column).
+Panels (a,b) correspond to MVAR(4), (c,d) to MVAR(9), (e,f) to LSTM(4), and (g,h) to LSTM(9) models.
+Appendix F Detailed results for the counterflow case
+Here, we provide the detailed results of the proposed framework for the counterflow case discussed in Section 4.2, including baseline errors of the restriction and lifting operators, training of the MVAR and LSTM model, open-loop and closed-loop reconstruction performance.
+Figure F.1 shows the accuracy of the restriction and lifting operators in Eqs.
+(36) and (37) constructed via the augmented basis in Eq.
+(20), displaying the evolution of the relative $L_{2}$ reconstruction error in time per group.
+Table F.1 reports the training results of the MVAR(10) and LSTM(10) models, while Table F.2 summarizes the open- and closed-loop end-to-end performance of the proposed framework.
+Figure F.2 displays the relative reconstruction errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ per group and their 10–90% percentile ranges for one-step predictions via the open-loop time-stepper (Eq.
+(E.1)).
+Errors are averaged over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$.
+Figures F.3 and F.4 compare the ground truth normalized density fields $x^{(l,c)}(t_{k})$ and the predicted ones $\hat{x}^{(,c)}(t_{k})$ for closed-loop simulations using MVAR(10) and LSTM(10) models, respectively.
+The test case uses a double gaussian initialization (15th row of Table C.2) at four time steps ($k=250,560,780,950$), during which the pedestrian groups do not pass through the corridor boundary.
+The two interacting populations are shown in red (group 1) and blue (group 2).
+Panels (a,c,e,g) show the ground truth, while panels (b,d,f,h) present the corresponding closed-loop predictions.
+The spatial distribution of the resulting absolute errors is shown in Figure F.5 at time steps $k=250,950$.
+(a) Augmented POD reconstruction error (group 1)
+(b) Augmented POD reconstruction error (group 2)
+Figure F.1: Accuracy of the restriction and lifting operators in Eqs.
+(36) and (37) for the counterflow case.
+Panels (a) and (b) depict the average (over $C=20$ cases with different initial conditions) reconstruction errors $e_{k}^{2,rec}$ in Eq.
+(13) for groups 1 and 2, respectively, in time for the training and testing datasets.
+The augmented basis in Eq.
+(20) is constructed with $d_{1}=6$, $d_{2}=8$ POD modes retained for the group-specific matrices $X^{(1)}_{tr},X^{(2)}_{tr}$ and $m=4$ modes retained for cross-covariance matrix $C^{(12)}$.
+Table F.1: Training results of MVAR(10) and LSTM(10) models at the latent space for the counterflow case.
+Loss functions $\mathcal{L}_{\text{MVAR}}(\cdot)$ in Eq.
+(41) for the MVAR model and $\mathcal{L}_{\text{LSTM}}(\cdot)$ in Eq.
+(46) for the LSTM model are reported at the end of training, along with the computational times (in seconds) required for training.
+For the LSTM, the mean and 10–90% error percentiles (in parentheses) are shown across the 50 independent training runs with different random initializations.
+Additionally, the MSE loss function of the MVAR model in Eq.
+(50) is reported for direct comparison to the MSE-based loss $\mathcal{L}_{\text{LSTM}}(\cdot)$.
+Model $\mathcal{L}_{\text{MVAR}}(\cdot)$ $MSE_{\text{MVAR}}(\cdot)$ $\mathcal{L}_{\text{LSTM}}(\cdot)$ Comput. Time (s)
+MVAR(10) $5.096\times 10^{-6}$ $1.07\times 10^{-11}$ - $0.031$
+LSTM(10) - - $1.44\ (1.38,1.71)\times 10^{-7}$ 104(84,226)104\ (84,226)
+Table F.2: Open-loop/one-step and closed-loop/recursive prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models for the counterflow case.
+The relative reconstructed $L_{1}$, $L_{2}$, and $L_{\infty}$ errors $e^{1,(c)}_{k}$, $e^{2,(c)}_{k}$, and $e^{\infty,(c)}_{k}$ in Eq.
+(52) are reported separately for group 1 and group 2, using the MVAR(10) and LSTM(10) latent dynamics models.
+Mean values and 10–90% error percentiles (in parentheses) are shown over all the $C=20$ cases and $K=990$ time steps.
+group 1 group 2
+$L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$ $L_{1}$ error, $e^{1,(c)}_{k}$ $L_{2}$ error, $e^{2,(c)}_{k}$ $L_{\infty}$ error, $e^{\infty,(c)}_{k}$
+Open-Loop
+MVAR(10) $0.063$ (0.041, 0.100)(0.041,\,0.100) $0.046$ (0.027, 0.076)(0.027,\,0.076) $0.046$ (0.024, 0.078)(0.024,\,0.078) $0.070$ (0.049, 0.089)(0.049,\,0.089) $0.048$ (0.034, 0.063)(0.034,\,0.063) $0.051$ (0.034, 0.073)(0.034,\,0.073)
+LSTM(10) $0.075$ (0.049, 0.109)(0.049,\,0.109) $0.057$ (0.034, 0.086)(0.034,\,0.086) $0.060$ (0.032, 0.095)(0.032,\,0.095) $0.082$ (0.061, 0.101)(0.061,\,0.101) $0.056$ (0.041, 0.071)(0.041,\,0.071) $0.064$ (0.043, 0.092)(0.043,\,0.092)
+Closed-Loop
+MVAR(10) $0.108$ (0.064, 0.164)(0.064,\,0.164) $0.083$ (0.046, 0.123)(0.046,\,0.123) $0.091$ (0.046, 0.151)(0.046,\,0.151) $0.118$ (0.086, 0.152)(0.086,\,0.152) $0.088$ (0.058, 0.115)(0.058,\,0.115) $0.099$ (0.060, 0.144)(0.060,\,0.144)
+LSTM(10) $0.132$ (0.090, 0.201)(0.090,\,0.201) $0.104$ (0.067, 0.168)(0.067,\,0.168) $0.113$ (0.061, 0.197)(0.061,\,0.197) $0.123$ (0.095, 0.157)(0.095,\,0.157) $0.095$ (0.068, 0.123)(0.068,\,0.123) $0.115$ (0.070, 0.152)(0.070,\,0.152)
+(a) Relative Errors with MVAR(10) (group 1)
+(b) Relative Errors with MVAR(10) (group 2)
+(c) Relative Errors with LSTM(10) (group 1)
+(d) Relative Errors with LSTM(10) (group 2)
+Figure F.2: One-step (open-loop) prediction errors, in the ambient–-density profile–-space, for the testing set over time, using trained MVAR and LSTM models for the counterflow case.
+The relative reconstructed $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq.
+(52) (using the one-step prediction in Eq.
+(E.1)) are shown for the MVAR(10) and LSTM(10) latent dynamics models in panels (a,b) (c,d) per group, respectively.
+Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step.
+For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
+(a) Ground truth density at $t_{250}=62.5\,\mathrm{s}$
+(b) Predicted density at $t_{250}=62.5\,\mathrm{s}$
+(c) Ground truth density at $t_{560}=140\,\mathrm{s}$
+(d) Predicted density at $t_{560}=140\,\mathrm{s}$
+(e) Ground truth density at $t_{780}=195\,\mathrm{s}$
+(f) Predicted density at $t_{780}=195\,\mathrm{s}$
+(g) Ground truth density at $t_{950}=237.5\,\mathrm{s}$
+(h) Predicted density at $t_{950}=237.5\,\mathrm{s}$
+Figure F.3: Comparison of the “ground truth” $x^{(l,c)}(t_{k})$ and the predicted $\hat{x}^{(l,c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(51) for groups $l=1,2$ in the counterflow case, using the MVAR(10) model for an unseen case of the testing set; initialization at the 15th row of Table C.2.
+Panels (a), (c), (e) and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f) and (h) show the predicted one, for the timesteps $t_{250}=62.5\,\mathrm{s}$, $t_{560}=140\,\mathrm{s}$, $t_{780}=195\,\mathrm{s}$, and $t_{950}=237.5\,\mathrm{s}$, respectively.
+The group-specific densities are superimposed with the blue and red lobes corresponding to groups 1 and 2, respectively.
+(a) Ground truth density at $t_{250}=62.5\,\mathrm{s}$
+(b) Predicted density at $t_{250}=62.5\,\mathrm{s}$
+(c) Ground truth density at $t_{560}=140\,\mathrm{s}$
+(d) Predicted density at $t_{560}=140\,\mathrm{s}$
+(e) Ground truth density at $t_{780}=195\,\mathrm{s}$
+(f) Predicted density at $t_{780}=195\,\mathrm{s}$
+(g) Ground truth density at $t_{950}=237.5\,\mathrm{s}$
+(h) Predicted density at $t_{950}=237.5\,\mathrm{s}$
+Figure F.4: Comparison of the “ground truth” $x^{(l,c)}(t_{k})$ and the predicted $\hat{x}^{(l,c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq.
+(51) for groups $l=1,2$ in the counterflow case, using the LSTM(10) model for an unseen case of the testing set; initialization at the 15th row of Table C.2.
+Panels (a), (c), (e) and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f) and (h) show the predicted one, for the timesteps $t_{250}=62.5\,\mathrm{s}$, $t_{560}=140\,\mathrm{s}$, $t_{780}=195\,\mathrm{s}$, and $t_{950}=237.5\,\mathrm{s}$, respectively.
+The group-specific densities are superimposed with the blue and red lobes corresponding to groups 1 and 2, respectively.
+(a) MVAR(10) absolute error at time $t_{250}=62.5\,\mathrm{s}$
+(b) MVAR(10) absolute error at time $t_{950}=237.5\,\mathrm{s}$
+(c) LSTM(10) absolute error at time $t_{250}=62.5\,\mathrm{s}$
+(d) LSTM(10) absolute error at time $t_{950}=237.5\,\mathrm{s}$
+Figure F.5: Absolute error between the “ground truth” $x^{(l,c)}(t_{k})$ and the predicted $\hat{x}^{(l,c)}(t_{k})$ normalized densities in the counterflow case, corresponding to the predictions for groups $l=1,2$ in Figs.
+F.3 and F.4.
+Panels (a)–(d) show the spatial distribution of the absolute error $\lvert x^{(l,c)}(t_{k})-\hat{x}^{(l,c)}(t_{k})\rvert$ in $\Omega$ at two indicative time steps $t_{250}=62.5\,\mathrm{s}$ (left column) and $t_{950}=237.5\,\mathrm{s}$ (right column).
+Panels (a,b) correspond to MVAR(10), and panels (c,f) to LSTM(10) models.